### v2

# A Weighted Loss Approach to Robust Federated Learning under Data Heterogeneity

Abstract
Federated learning (FL) is a machine learning paradigm that enables multiple data holders to collaboratively train a machine learning model without sharing their training data with external parties. In this paradigm, workers locally update a model and share with a central server their updated gradients (or model parameters). While FL seems appealing from a privacy perspective, it opens a number of threats from a security perspective as (Byzantine) participants can contribute poisonous gradients (or model parameters) harming model convergence. Byzantine-resilient FL addresses this issue by ensuring that the training proceeds as if Byzantine participants were absent. Towards this purpose, common strategies ignore outlier gradients during model aggregation, assuming that Byzantine gradients deviate more from honest gradients than honest gradients do from each other. However, in heterogeneous settings, honest gradients may differ significantly, making it difficult to distinguish honest outliers from Byzantine ones. In this paper, we introduce the Worker Label Alignement Loss (WoLA), a weighted loss that aligns honest worker gradients despite data heterogeneity, which facilitates the identification of Byzantines’ gradients. This approach significantly outperforms state-of-the-art methods in heterogeneous settings. In this paper, we provide both theoretical insights and empirical evidence of its effectiveness.
1 Introduction
Federated learning (FL) [19, 18, 41] enables a group of machines (workers) to collaboratively train a model without sharing their private data. In this paradigm, a central server coordinates the process by sending the current model to all workers. Each worker then performs a training step on its local data and sends an update back to the server. Typically, this update is the gradient computed from a mini-batch [19, 3]. The server aggregates the gradients, updates the model, and sends it back to the workers, repeating the cycle.
In an ideal setup, the server aggregates gradients by simply averaging them. However, some workers may fail or behave maliciously by sending corrupted updates (also referred to as Byzantine behaviour). Robust training seeks to achieve model performance comparable to that achieved in the absence of Byzantine workers. To this end, a common strategy is to replace simple averaging with robust aggregation methods on the server side [19, 18]. These methods are designed to reduce the impact of outliers, which are expected to include Byzantine updates. However, their effectiveness strongly depends on the assumption that honest workers hold homogeneous data distributions [18, 3, 33, 4]. In practice, this assumption may not hold [3, 24, 29, 15], as each worker typically has access to only a small, potentially unrepresentative portion of the dataset.
With heterogeneous data, the divergence among honest gradients increases, making it easier for Byzantine machines to mask their updates [3, 19, 24]. In this setting, a robust aggregator alone is not enough to resist Byzantine attacks. A common approach is to add a pre-aggregation step, such as Bucketing [24] or Nearest Neighbor Mean (NNM) [24, 3, 19, 18]. While effective under weak adversarial conditions (i.e., when the proportion of Byzantine updates is low), we show that the performance of these solutions drops significantly in stronger adversarial scenarios.
We present a novel solution for robust FL in heterogeneous and highly adversarial settings. Specifically, we address label distribution skew [29, 35, 46]—a form of data heterogeneity where label distributions vary across workers, while the feature distribution for each class remains consistent across them. This scenario is common in practice [29, 35, 46]. For instance, in a mobile image recognition app, users of different ages may capture different subjects, resulting in variations in label distributions across devices. In another case, hospitals may specialize in certain types of diseases, leading to an overrepresentation of specific labels in their local patient records.
Label skewness leads to inconsistent training across workers, as each worker optimizes the model based on its own label distribution [24, 46, 4]. Local training tends to overemphasize majority classes while neglecting minority ones [46, 4]. Since these class imbalances vary across workers, local objectives become misaligned, leading to highly inconsistent updates.
Rather than attempting to reduce variance of updates through pre-aggregation, we address the root cause: misaligned worker objectives. We introduce the Worker Label Alignment (WoLA) loss, a weighted loss function that aligns local training objectives. WoLA reweights point-wise errors for each worker to generate gradients as if all local datasets were drawn from the same label distribution. By steering honest gradients in a common direction, WoLA reduces variance and improves Byzantine resilience.
WoLA is model-agnostic, requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine machines, and adds no computational cost. It is also straightforward to implement and can be combined with other robustness strategies.
The only requirement is that a common training objective—represented by a probability distribution—is fixed and shared by the server once before training begins. For instance, this objective could be the global label distribution (i.e., the distribution obtained by merging all local datasets), derived either from prior knowledge or from local label distributions shared by the workers with the server. Using this information, the server estimates and broadcasts the global distribution. Despite the fact that Byzantine nodes can lie in their data distributions ahead of training, we show in this paper that their influence remains limited and that this approach still brings a significant gain in robustness compared to state-of-the-art approaches.
Our contributions can be summarized as follows:
  • We introduce WoLA, a weighted loss function designed to address label skewness in highly adversarial settings (Section 5).
  • We provide theoretical insights for the effectiveness of WoLA, with all proofs included in the appendix (Section 6)
  • We empirically show that WoLA significantly outperforms state-of-the-art methods through extensive experiments involving various models, datasets, aggregators, and Byzantine attacks. Our code is publicly available [^1] https://anonymous.4open.science/r/WoLA-E4CB to ensure reproducibility (Section 8).
The paper is organized as follows. Section 2 reviews the background and related work. Section 4 introduces the motivation behind the WoLA. Section 5 formally defines the loss and explores its properties. Section 6 presents theoretical justifications for its design. Section 7 describes the experimental setup, and Section 8 presents the experimental results. Finally, Section 9 concludes the paper and outlines directions for future work.
2 Background
This section outlines key FL concepts and defines our threat model. All notations are summarized in Table 1.
Table 1: Notation—In (c) and (d), superscript indices denote classes, while subscripts denote workers.
Notation Meaning
$[k]$ Set of integers from $1$ to $k$, with $k\in\mathbb{N}$
##\## Cardinality of a set
$\mathbbm{1}$ Indicator function. For example, $\mathbbm{1}(x\leq 0)=1$ if $x\leq 0$, and 00 otherwise.
$\nabla$ Gradient
[…]
$\mathbb{E}[X]$ Expected value of the random variable $X$
$\mathbb{E}[X|A]$ Conditional expectation of the random variable $X$ given the event $A$
i.i.d. Independent and identically distributed
a.s. Almost surely is a probabilistic concept meaning that an event happens with probability $1$.
[…]
Notation Meaning
$\Phi$ Model
$n$ Number of workers
$\mathcal{H}$ Set of honest workers, with $\mathcal{H}\subset[n]$
$f$ Number of Byzantine workers, with $f=n-\#\mathcal{H}$
Notation Meaning
$C$ Number of classes
$\mathcal{D}$ Global training set, composed of pairs $(x,y)\in\mathcal{X}\times[C]$, where $x$ is an observation, and $y$ its corresponding label.
$\mathcal{D}_{i}$ Local dataset of worker $i$, with $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$
$N$ Size of the global dataset: $N=\#\mathcal{D}$
$N_{i}$ Size of worker $i$’s local dataset: $N_{i}=\#\mathcal{D}_{i}$
$N^{c}$ Number of class-$c$ instances in the global dataset, with $N=\sum_{c=1}^{C}N^{c}$
$N_{i}^{c}$ Number of class-$c$ instances in worker $i$’s dataset, with $N_{i}=\sum_{c=1}^{C}N_{i}^{c}$
$p$ Label distribution of $\mathcal{D}$: $p=(p^{1},p^{2},\ldots,p^{C})$ with $p^{c}=N^{c}/N$
$p_{i}$ Label distribution of $\mathcal{D}_{i}$: $p_{i}=(p_{i}^{1},p_{i}^{2},\ldots,p_{i}^{C})$ with $p^{c}_{i}=N_{i}^{c}/N_{i}$
Notation Meaning
$\mathcal{L}$ Loss
$\ell(y,\Phi(x))$ Pointwise loss of prediction $\Phi(x)$ and label $y$
$\mathcal{L}_{i}$ Loss of worker $i$: $\mathcal{L}_{i}=\sum_{(x,y)\in\mathcal{D}_{i}}\ell\Big{(}y,\Phi(x)\Big{)}/N_{i}$, with $\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\ N_{i}/N=\mathcal{L}$
$\mu^{c}$ Class-wise gradient of class $c$ in $\mathcal{D}$, i.e., $\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}$
$\mu^{c}_{i}$ Class-wise gradient of class $c$ in $\mathcal{D}_{i}$, i.e., $\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}_{i}$
$\mathcal{WL}$ Worker Label Alignement Loss (WoLA)
$\mathcal{WL}_{i}$ WoLA of worker $i$, with $\sum_{i\in\mathcal{H}}\mathcal{WL}_{i}\ N_{i}/N=\mathcal{WL}$
$q$ Shared training objective across workers: a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$.
2.1 Byzantine-free FL
We consider a distributed architecture with $n$ workers and a central server, aimed at solving a classification task with $C$ classes. The workers hold local datasets $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$. Each worker $i$ holds $N_{i}$ data points $(x,y)\in\mathcal{X}\times[C]$, where $\mathcal{X}$ is the input space, $x$ is an observation, and $y$ is its label. Given a model $\Phi$, the standard local loss function (not WoLA) for worker $i$ is:
$\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\ell\big{(}y,\Phi(x)\big{)},$
where $\ell$ is a point-wise loss function, assumed to be differentiable with respect to the model parameters.
In each communication round, the server broadcasts the model parameters to all workers. Each worker $i$ is expected to compute the gradient using its own data $\mathcal{D}_{i}$ and send back the resulting gradient $\nabla\mathcal{L}_{i}$.
2.2 Threat Model
In this paper, we focus on handling Byzantine workers under the common assumption of a trusted FL server [3, 24, 4]. We consider an adversarial setting where fewer than half the workers are Byzantine, i.e., $f<n/2$. These malicious workers can send arbitrary updates to the server and have full knowledge of all local updates, following the threat model in [3, 24, 4]. The server’s objective is to identify a stationary point of the average loss over the honest workers. That is, to find a point where $\nabla\mathcal{L}=0$, with
$\mathcal{L}:=\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\frac{N_{i}}{N},$
and $N=\sum_{i\in\mathcal{H}}N_{i}$. Assuming such a point exists, it can be approximated via gradient descent and robustness strategies.
2.3 Global Dataset, Local Datasets & Label Distribution Skew
To model data heterogeneity, we consider the following label skew scenario: all data points are drawn from independent and identically distributed (i.i.d) random variables, forming the global dataset $\mathcal{D}$. These points are then randomly assigned to honest workers, resulting in local datasets $\mathcal{D}_{i}$ for $i\in\mathcal{H}$, such that $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$. Each worker receives data with a different label distribution, introducing heterogeneity. Although Byzantine workers may also have their own data, we exclude them from the global dataset as they play no role in our analysis.
The global label distribution refers to the distribution of labels in the full dataset $\mathcal{D}$, while the local label distributions refer to label distributions in a worker’s datasets $\mathcal{D}_{i}$, for $i\in\mathcal{H}$.
2.4 Gradient Dissimilarity
Karimireddy et al. [24] introduce Gradient Dissimilarity as a measure of system robustness, lower gradient dissimilarity implies stronger robustness. It quantifies how much individual gradients deviate from the average of the honest gradients:
Gradient $\text{Gradient Dissimilarity}=\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\|\nabla\mathcal{L}_{i}-\nabla\mathcal{L}\|^{2},$
Data heterogeneity tends to increase gradient dissimilarity [19, 18, 24].
3 Related Work
This section reviews state-of-the-art approaches relevant to our context.
3.1 Robustness Strategies with Homogeneous Data
To mitigate Byzantine behavior during aggregation, a common approach is to replace simple updates averaging with robust aggregation methods [34, 19]. Examples include the coordinate-wise median (CWMed) [45], geometric median (GM) [38, 34], coordinate-wise trimmed mean (CwTM) [45], Krum [7], Multi-Krum (MKrum) [20], Minimum Diameter Averaging [20], and Comparative Gradient Elimination [22]. These methods heavily rely on the assumption that honest workers have homogeneous data. Their robustness under data heterogeneity remains unclear [18, 3, 33].
Besides robust aggregators, additional strategies can enhance robustness in homogeneous settings. These approaches rely on a clean dataset held by the server to evaluate and filter worker updates [44, 9, 14, 36].
3.2 Robustness Strategies Under Data Heterogeneity
We begin by defining Gradient Dissimilarity, a robustness measure commonly used in studies addressing Byzantine resilience under data heterogeneity [19, 18, 24]. Next, we review state-of-the-art approaches in Byzantine FL, followed by methods developed for the Byzantine-free setting. Finally, we highlight the advantages of our proposed method.
3.2.1 Byzantine FL
A common strategy to reduce gradient variance is to apply a pre-aggregation step, such as Bucketing [24], Nearest-Neighbor Mixing (NNM) [3], or FoundationFL [15]. These methods act before the main aggregation rule, resulting in a two-step aggregation process—for example, applying NNM before Krum. Other methods like BOBA [4] and RAGE [12] use different strategies and do not rely on pre-aggregation.
Bucketing [24] randomly partitions inputs into groups (buckets), averages the inputs in each bucket, and feeds these averages into the aggregation rule. This random partitioning reduces the empirical variance of honest updates in expectation (i.e., on average). However, in some iterations, it may fail to reduce heterogeneity, giving Byzantine workers a better chance to disrupt learning [18].
In contrast, NNM [3] is a deterministic pre-aggregation method that replaces each gradient with the average of its $n-f$ nearest neighbors (including itself). NNM introduces additional computational cost, which prevents its use in large-scale distributed learning tasks [18].
FoundationFL [15] synthesizes additional updates at each training step. The server assigns a score to each update and replicates the highest-scoring gradient $m$ times, where $m$ is a hyperparameter set by the authors to $m=n/2$. A critical limitation of FoundationFL is its underlying assumption [^2] The score is based on the distance to the coordinate-wise minimum and maximum. Given updates $g_{1},\dots,g_{n}$, define $g_{\text{max}}=\max(g_{1},\dots,g_{n})$ and $g_{\text{min}}=\min(g_{1},\dots,g_{n})$ (coordinate-wise). The score $s_{i}$ for worker $i$ is then $s_{i}=\min(\|g_{\text{max}}-g_{i}\|,\|g_{\text{min}}-g_{i}\|)$. A high score means the update is far from the extremes.: extreme values correspond to Byzantine updates. However, this assumption is not reliable under data heterogeneity, particularly against the A Little Is Enough (ALIE) attack [5].
BOBA [4] differs from standard robustness strategies. It projects client gradients onto a $C-1$ dimensional subspace using truncated Singular Value Decomposition (SVD), estimates the $C-1$ simplex from a server-side dataset, reconstructs each client’s label distribution based on their position relative to the simplex, and discards those with abnormal values. In addition to requiring server-side data, at each round, BOBA incurs extra cost from computing the truncated SVD of the gradient matrix, which scales with both the number of clients and the gradient dimension [^3] The computational complexity of BOBA is $\mathcal{O}(kCnd)$, where $k$ is the number of truncated SVDs computed per round (typically between 3 and 5 in thei experiments), $C$ is the number of classes, $n$ the number of workers, and $d$ the gradient dimension..
RAGE [12] performs a weighted average using confidence scores assigned to gradients. These scores are computed iteratively, downweighting gradients with large projections on the principal eigenvector of the empirical covariance matrix of gradients. The covariance is recomputed with updated scores until the largest eigenvalue is sufficiently small. The computational complexity of RAGE makes it impractical for large-scale optimization [24].
3.2.2 Byzantine-free FL
In a Byzantine-free setting, non-IID data remains a major obstacle to model convergence. Several techniques have been proposed to address this challenge, including knowledge distillation [^4] Knowledge distillation encourages a student model to approximate the teacher model’s output logits, enabling it to mimic the teacher’s behavior. and logit calibration.
Some methods rely on knowledge distillation and require access to an additional test set [28, 11, 10, 30]. In contrast, [47] and [39] avoid relying on proxy datasets. They use a generator trained on the server or client to enable data-free knowledge distillation. However, these approaches require significant communication and coordination, and introduce substantial computational overhead [40, 21].
Other approaches introduce personalized loss functions based on local data distributions, including DMFL [35], FedLC [46], and FedDisk [32]. DMFL and FedLC, inspired by [8], adjust logits [^5] Logits are the outputs of the final classification layer and inputs to the softmax. before applying softmax cross-entropy, increasing margins for minority classes. These methods require the tuning of a hyperparameter with no bounded range. FedDisk uses loss with sample-level weights, learned from a global and multiple local density estimators, independent of class labels.
All these propositions either require an additional tes set, introduce significant overhead [40, 21, 32], or rely on hyperparameter tuning—a particularly difficult task in our setting: (i) without a test set, local tuning is possible, but data heterogeneity may cause inconsistent values; and (ii) since these methods assume a Byzantine-free setting, no robust protocol is provided to handle malicious behavior.
3.2.3 Our proposition
To the best of our knowledge, using a loss function to improve Byzantine resilience under data heterogeneity is a novel approach. WoLA is a model-agnostic method that requires no hyperparameters, no server-side dataset, and no knowledge of the number of Byzantine machines. It adds no computational cost, is easy to implement, and can be combined with other robustness strategies. Its only requirement is a shared training objective.
While many approaches focus on handling update heterogeneity [3, 24, 4, 15, 12], WoLA tackles its root cause—label skewness—by aligning gradients from the start. This enables WoLA to significantly outperform state-of-the-art methods. In the next section, we explain the motivation behind its design.
4 Rationale for Worker Label Alignement Loss
This section first shows that local gradients approximate weighted averages of class-wise gradients. Next, it illustrates that class-wise gradients can differ significantly. Finally, it presents the main conclusion drawn from these observations.
4.1 Gradients as Weighted Averages
The key rationale behind WoLA is that the gradient over a sufficiently large local dataset can be approximated as a weighted average of class-wise gradients, with weights given by the local label distribution.
More precisely, the gradient of worker $i$ for a standard loss is:
$\nabla\mathcal{L}_{i}$ $=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}$
$=\sum_{c=1}^{C}\frac{N^{c}_{i}}{N_{i}}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}},$
where:
$\frac{N^{c}_{i}}{N_{i}}=\Bigg{(}\text{Proportion of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=p^{c}_{i},$
and
$\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}}$ (1)
$\quad\quad=\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=\mu_{i}^{c}$
Given a sufficient number of instances of class $c\in[C]$, the following approximation is reasonable (see Appendix Section A.2, Lemma 2 for a formal statement):
(Mean gradient of class $\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}$ (2)
$\quad\quad\approx\Bigg{(}\text{Mean gradient of class }c\text{ in}\mathcal{D}\Bigg{)}$
$\quad\quad=\frac{\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}}\vcentcolon=\mu_{c}$
We call $\mu_{i}^{c}$ the class-$c$ gradient of worker $i$, and $\mu^{c}$ the global class-$c$ gradient. When clear from context, we omit the term global.
Using approximation equation (2), we can express the local gradient as a weighted average of global class gradients:
$\nabla\mathcal{L}_{i}\approx\sum_{c=1}^{C}p_{i}^{c}\mu^{c}$ (3)
Assuming that the vectors $\mu^{c}$ for $c\in[C]$ differ—either in norm or direction—the differences in local gradients arise from the local label distributions, through the weights $p_{i}^{c}$.
In the next subsection, we show that the vectors $\mu^{c}$ for $c\in[C]$ are indeed different.
4.2 Difference in Class Gradients
Figure 1 shows standard (non-FL) training of a shallow neural network on the Iris dataset [17]. The loss decreases and accuracy improves over epochs, indicating successful model convergence. Independently of the optimization, at each step we compute the class gradients for Setosa, Versicolour, and Virginica, and track the cosine similarity between these gradients over training steps.
The figure reveals that the angles between class-wise gradients often exceed $\pi/2$, indicating that the gradients point in widely different directions. In some cases, such as between Versicolour and Virginica, the gradients are nearly opposite.
This highlights that different classes can push the model in distinct—and sometimes conflicting—directions during training. Similar observations have been reported in previous studies [46, 4].
[…]
4.3 Key Takeaways
Local gradients approximate weighted averages of global class gradients, with weights given by the local label distribution. Since class gradients may differ, aligning local gradients requires aligning their weightings—as if local datasets shared the same label distribution. WoLA is designed to achieve this goal.
5 Worker Label Alignement Loss
This section introduces WoLA. It first defines the training objective required for WoLA, then presents its formula. The final subsections provide an intuitive understanding of WoLA’s properties. These properties are rigorously stated in Section 6 and proved in the Appendix A.
5.1 Shared Training Objective
WoLA requires a shared training objective across workers. This objective is formalised by a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$. A larger $q^{i}$ means that point-wise errors for samples with label $i$ are given more weight during training, and vice versa. To define this distribution, we considered three situations:
Prior knowledge
The global label distribution is known in advance, based on expert knowledge or historical data. For example, in tumor detection on MRI scans or crop type recognition from satellite imagery, national statistics can serve as a reliable proxy for label distributions.
Local label distributions are shared
Each worker shares its local label distribution with the server before training. The server then computes the global distribution. In this case, it is important to note that (i) the process reduces data privacy, and (ii) it introduces a new opportunity for Byzantine workers to attack the system. We carefully evaluate this attack scenario in Section 7.2. Regarding privacy concerns, we assume in this paper that the server is trusted as we consider privacy as an orthogonal problem that is actively being investigated in the literature. Nevertheless, off-the-shelf SecureAggregation protocols such as [6] could be easily used to release the trusted server assumption.
Consensus among workers on the training objective
All workers agree that the global model should favor certain classes during training, regardless of the actual global label distribution. For example, in a task to detect hateful or illegal content on social networks—where hate content is rare—it may be agreed to assign higher importance to these cases. In medical diagnosis, rare diseases may also be given more weight. In contrast, for automated recruitment systems, all classes may be weighted equally to avoid algorithmic bias.
The training objective $q$ is shared by the server with all workers once before training begins, allowing them to adjust their local loss functions accordingly.
5.2 WoLA definition
WoLA locally adjusts the weights of point-wise errors to simulate an identical label distribution across workers. More precisely, WoLA for worker $i$, denoted $\mathcal{WL}_{i}$, is defined as:
$\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\ell\Big{(}y,\Phi(x)\Big{)},$
While the formula is undefined for $p^{y}_{i}=0$, it is not problematic, as it indicates that label $y$ is absent from the local dataset $\mathcal{D}_{i}$, meaning the corresponding weight never appears in the sum.
A sample from class $c$ is overweighted when class $c$ is underrepresented locally compared to the training objective, i.e., when $p_{i}^{c}<q^{c}\Leftrightarrow 1<q^{c}/p_{i}^{c}$; and underweighted otherwise.
5.3 Local Updates Alignement
WoLA makes local datasets appear as if they had identical label distributions matching the training objective $q$—that is, as if $p_{i}^{c}=q^{c}$ for all $c\in[C]$ and all $i\in\mathcal{H}$. Specifically, for worker $i$, the gradient under WoLA is:
$\nabla\mathcal{WL}_{i}\vcentcolon=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\frac{q^{c}}{p^{c}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\sum_{c=1}^{C}q^{c}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}y,\Phi(x)\Big{)}}{N^{c}_{i}},$
using approximation in equation (2), it follows:
$\nabla\mathcal{WL}_{i}\approx\sum_{c=1}^{C}q^{c}\mu^{c}$ (4)
With WoLA, the approximation of the worker $i$’s gradient no longer depends on local label proportions. It becomes a weighted average of class gradients, using the shared training objective $q$ as weights. Since equation (4) holds for all $i\in[C]$, it follows that $\nabla\mathcal{WL}_{i}\approx\nabla\mathcal{WL}_{j}$ for all $j\in[C]$. This suggests that local updates are more similar under WoLA, potentially reducing gradient dissimilarity and improving robustness.
Figure 2 illustrates the impact of WoLA on updates using a fictitious example with three classes: Dog, Flower, and Frog, under the training objective $q=(1/3,1/3,1/3)$. It shows three local datasets with label distributions $(3/6,2/6,1/6)$, $(2/6,2/6,2/6)$, and $(1/6,1/6,4/6)$. Two of them differ from the training objective $q$. WoLA addresses this mismatch by upweighting samples from local minority classes and downweighting those from local majority classes, thereby aligning updates.
(a) Point-wise standard loss gradients on $\mathcal{D}_{1}$
(b) Point-wise standard loss gradients on $\mathcal{D}_{2}$
(c) Point-wise standard loss gradients on $\mathcal{D}_{3}$
(d) Point-wise WoLA gradients on $\mathcal{D}_{1}$
(e) Point-wise WoLA gradients on $\mathcal{D}_{2}$
(f) Point-wise WoLA gradients on $\mathcal{D}_{3}$
Figure 2: Fictitious FL classification task with three classes (Dog, Flower, Frog) under the training objective $q=(1/3,1/3,1/3)$, comparing standard loss with WoLA. Each image represents a data point from a local dataset. Arrows indicate 2D point-wise loss gradients; in black squares, arrows show the resulting updates. Subfigures (a–c) display gradients under the standard loss, while (d–f) show gradients with WoLA. In (d) and (f), dataset sizes are scaled to reflect the effect of WoLA weights, although datasets are actually unchanged. For example, $\mathcal{D}_{3}$ has a local label distribution of $(1/6,1/6,4/6)$, while $q$ is uniform. This results in weights: $1/3\div 1/6=4/2$ for Dog and Flower, and $1/3\div 4/6=1/2$ for Frog. To avoid fractional images, the denominator 2 is omitted, giving 4 duplicates of Dog and Flower gradients, and 1 of Frog. The same logic applies to $\mathcal{D}_{1}$. Subfigures (b) and (e) are identical, as $\mathcal{D}_{2}$ label distribution matches the training objective. Local updates with WoLA are more aligned than with the standard loss.
5.4 Targeted Model
Assuming that the training objective $q$ matches the global label distribution $p$, i.e., $p=q$, the average gradients of the standard loss and the WoLA loss—denoted $\nabla\mathcal{L}$ and $\nabla\mathcal{WL}$ respectively—are similar. Using the approximations from equations (3) and (4), we have:
$\nabla\mathcal{L}$ $=\sum_{i\in\mathcal{H}}\nabla\mathcal{L}i\frac{N_{i}}{N}$
$\approx\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p_{i}^{c}\mu^{c}\frac{N_{i}}{N}=\sum_{c=1}^{C}p^{c}\mu^{c}=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p^{c}\mu^{c}\frac{N_{i}}{N}$
$=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}q^{c}\mu^{c}\frac{N_{i}}{N}\approx\sum_{i\in\mathcal{H}}\nabla\mathcal{WL}_{i}\frac{N_{i}}{N}=\nabla\mathcal{WL}$
Therefore, when $p=q$, the training targets the same model regardless of whether the standard or WoLA loss is used.
5.5 Key Takeaways
WoLA requires a shared training objective. Previous analyses suggest that: (i) using WoLA reduces gradient dissimilarity, and (ii) when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
6 Theoretical Arguments
This section uses probabilistic tools to rigorously state the properties of WoLA. We first describe the data modeling framework, then present asymptotic results that offer insights into WoLA’s behavior.
6.1 Data Modeling
Let $(X_{1},Y_{1},Z_{1}),(X_{2},Y_{2},Z_{2}),\ldots,(X_{N},Y_{N},Z_{N})$ be independent and identically distributed (i.i.d.) random triplets drawn from a distribution $\mathfrak{D}$ over the sample space $\mathcal{X}\times[C]\times\mathcal{H}$. Each pair $(X_{i},Y_{i})$ represents a datapoint in the dataset $\mathcal{D}$. For each $k\in[N]$, the random variable $Z_{k}$ takes values in $\mathcal{H}$ and indicates the worker to which the $k$-th datapoint is assigned. We also denote by $(X,Y,Z)$ a generic triplet with the same distribution $\mathfrak{D}$.
For any triplet $(X_{k},Y_{k},Z_{k})$, we assume that $X_{k}$ and $Z_{k}$ are conditionally independent given $Y_{k}$. That is, conditioned on the label, the features provide no information about the assignment, and vice versa. Formally, $P(X_{k},Z_{k}|Y_{k})=P(X_{k}|Y_{k})P(Z_{k}|Y_{k}).$
6.2 Theoretical Results
The results are asymptotic, considering the case where the size of the global dataset tends to infinity, i.e., $N\to\infty$, and are based on the Law of Large Numbers. Full proofs and details are given in Appendix A.
Let $L_{i}$ and $W_{i}$ be the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$, respectively. Let $L$ and $W$ denote the average gradients of honest workers for the standard loss and WoLA, respectively.
With the standard loss, local gradients remain different even with an infinite amount of data:
Proposition 1
Under the standard loss, the gradient of worker $i$ converges almost surely as $N\to\infty$:
[…]
Moreover, the average gradient satisfies:
[…]
In contrast, with WoLA, local gradients become asymptotically identical, regardless of the training objective shared by the server:
Proposition 2
Let $q$ be the training objective distribution shared by the server. Under WoLA, the gradient of worker $i$ converges almost surely as $N\to\infty$:
[…]
Moreover, the average gradient also satisfies:
[…]
This directly implies that, under WoLA, the gradient dissimilarity vanishes asymptotically as $N\to\infty$ (see Appendix A.5).
Finally, if the training objective matches the global label distribution, i.e., $q=p$, then the standard loss gradient and the WoLA gradient become asymptotically identical:
Proposition 3
Let $q$ be the training objective and $p$ the global label distribution. If $q=p$, then:
[…]
This equality implies that, in the asymptotic regime, when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
(a) Results under ALIE attack.
(b) Results under LF attack.
[…]
(a) Boxplot of accuracy gains based on Table 2.
(b) Boxplot of accuracy gains under the strongest attack, based on Table 3.
[…]
Table 2: Global Performance: Test Accuracy (s.d.) %, averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 93.6 (0.3) 92.9 (0.4) 86.1 (1.6) 66.3 (4.5) 93.6 (0.3) 92.2 (0.6) 82.3 (3.3) 43.5 (11.7) 91.9 (1.8) 84.6 (4.0) 58.8 (12.1) 17.3 (7.3)
BKT 93.5 (0.3) 92.8 (0.5) 86.1 (1.6) 66.3 (4.5) 93.6 (0.4) 92.5 (0.7) 82.3 (3.3) 43.5 (11.7) 93.3 (0.7) 89.6 (1.9) 58.8 (12.1) 17.3 (7.3)
FoundFL 89.6 (1.8) 74.4 (4.2) 56.8 (4.6) 47.5 (5.0) 85.4 (5.2) 69.6 (7.3) 57.7 (6.7) 50.4 (6.8) 61.6 (15.9) 50.4 (12.8) 43.0 (12.1) 38.0 (11.8)
NNM 93.7 (0.2) 93.3 (0.3) 92.7 (0.4) 75.5 (6.9) 93.9 (0.3) 93.0 (0.6) 90.1 (2.5) 63.2 (9.0) 93.9 (0.6) 91.5 (1.4) 67.5 (14.5) 35.2 (8.4)
WoLA 93.7 (0.3) 93.4 (0.3) 92.7 (0.3) 81.9 (0.6) 93.9 (0.3) 93.5 (0.4) 92.4 (0.4) 80.5 (2.2) 93.6 (0.5) 93.0 (0.7) 88.9 (2.6) 54.5 (8.0)
WoLA+NNM 93.7 (0.3) 93.5 (0.3) 93.2 (0.3) 84.3 (0.6) 94.0 (0.3) 93.6 (0.4) 93.2 (0.3) 83.1 (2.8) 93.7 (0.6) 93.4 (0.7) 90.0 (3.0) 64.6 (8.7)
WoLA 93.7 (0.3) 93.3 (0.3) 92.5 (0.3) 81.3 (0.8) 93.8 (0.3) 93.3 (0.3) 92.1 (0.6) 76.8 (5.4) 93.6 (0.6) 92.9 (0.7) 88.1 (2.8) 47.3 (17.0)
WoLA+NNM 93.7 (0.3) 93.4 (0.3) 93.0 (0.3) 83.8 (0.7) 93.9 (0.4) 93.5 (0.3) 92.8 (0.4) 73.5 (13.0) 93.7 (0.6) 93.3 (0.7) 89.8 (2.8) 48.7 (15.8)
Accuracy Gain Over Top Non-WoLA Methods
WoLA 0.0 0.0 0.0 +6.4 0.0 +0.4 +2.4 +17.3 -0.4 +1.5 +21.4 +16.5
WoLA+NNM 0.0 +0.2 +0.6 +8.8 +0.1 +0.6 +3.1 +19.9 -0.2 +1.8 +22.5 +26.6
WoLA 0.0 0.0 -0.2 +5.9 -0.1 +0.3 +2.0 +13.6 -0.4 +1.4 +20.5 +9.3
WoLA+NNM 0.0 +0.1 +0.3 +8.3 -0.1 +0.5 +2.8 +10.3 -0.2 +1.8 +22.3 +10.7
(b) Fashion MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 77.1 (0.7) 74.0 (1.2) 64.7 (3.1) 45.3 (6.3) 76.6 (1.9) 71.6 (2.9) 59.2 (5.5) 32.3 (7.7) 72.6 (2.6) 60.3 (4.9) 41.2 (7.4) 16.7 (6.9)
BKT 76.8 (0.7) 74.5 (1.1) 64.7 (3.1) 45.3 (6.3) 77.2 (1.7) 73.6 (2.5) 59.2 (5.5) 32.3 (7.7) 75.7 (2.8) 67.4 (4.7) 41.2 (7.4) 16.7 (6.9)
FoundFL 71.2 (3.1) 62.2 (5.8) 52.3 (3.9) 44.7 (7.7) 63.6 (6.9) 54.9 (7.0) 47.7 (8.9) 41.6 (9.5) 47.8 (10.6) 41.7 (9.6) 38.1 (9.1) 31.2 (9.3)
NNM 78.4 (0.6) 76.5 (0.9) 72.3 (1.8) 54.8 (5.9) 79.4 (1.3) 76.7 (1.7) 69.9 (3.3) 48.8 (7.4) 79.4 (1.9) 74.0 (3.5) 54.2 (9.2) 30.5 (6.7)
WoLA 78.6 (0.6) 77.6 (0.6) 75.8 (0.8) 63.9 (1.4) 79.5 (1.3) 78.1 (1.6) 76.0 (1.7) 62.3 (4.0) 78.0 (1.8) 76.3 (2.3) 70.8 (4.8) 42.3 (7.3)
WoLA+NNM 78.6 (0.6) 78.0 (0.6) 77.1 (0.7) 68.5 (1.0) 79.6 (1.3) 78.6 (1.5) 77.5 (1.4) 67.5 (4.0) 78.5 (1.8) 77.4 (2.3) 73.9 (5.0) 48.8 (8.6)
WoLA 78.5 (0.5) 77.3 (0.5) 74.5 (0.9) 59.1 (2.3) 79.3 (1.5) 76.9 (1.6) 72.8 (2.0) 55.0 (6.0) 77.7 (1.7) 73.7 (2.4) 64.9 (4.3) 36.4 (11.2)
WoLA+NNM 78.5 (0.6) 77.6 (0.4) 75.9 (0.8) 64.2 (2.1) 79.6 (1.5) 77.3 (1.6) 74.9 (1.5) 55.7 (9.3) 78.3 (1.6) 75.0 (2.3) 69.3 (4.8) 38.9 (8.3)
Accuracy Gain Over Top Non-WoLA Methods
WoLA +0.2 +1.1 +3.6 +9.1 +0.1 +1.4 +6.1 +13.4 -1.4 +2.2 +16.6 +11.1
WoLA+NNM +0.2 +1.5 +4.8 +13.7 +0.2 +1.9 +7.6 +18.7 -0.9 +3.4 +19.8 +17.6
WoLA +0.1 +0.7 +2.2 +4.3 -0.1 +0.2 +2.9 +6.2 -1.7 -0.4 +10.7 +5.3
WoLA+NNM +0.1 +1.1 +3.6 +9.4 +0.2 +0.6 +5.0 +6.9 -1.1 +0.9 +15.1 +7.7
(c) Purchase100
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 70.4 (0.7) 65.3 (1.1) 50.5 (1.9) 24.8 (2.8) 67.9 (1.0) 58.0 (2.0) 37.1 (3.1) 8.0 (2.1) 62.2 (1.8) 44.4 (3.3) 15.9 (2.3) 2.6 (1.1)
BKT 70.1 (0.9) 64.6 (1.2) 50.5 (1.9) 24.8 (2.8) 68.2 (1.1) 59.3 (1.9) 37.1 (3.1) 8.0 (2.1) 64.5 (1.9) 49.3 (3.5) 15.9 (2.3) 2.6 (1.1)
FoundFL 52.2 (3.6) 36.5 (3.1) 28.4 (2.6) 22.4 (2.4) 38.8 (5.9) 28.9 (4.1) 20.7 (4.5) 13.3 (4.9) 24.6 (8.2) 15.6 (6.6) 10.3 (4.3) 6.4 (2.8)
NNM 71.6 (0.7) 67.9 (0.9) 61.3 (1.8) 38.1 (2.3) 70.6 (0.9) 63.9 (1.6) 52.0 (2.8) 28.3 (4.0) 69.6 (1.2) 58.8 (3.2) 28.9 (5.6) 15.1 (3.2)
WoLA 70.6 (0.6) 69.2 (0.6) 66.7 (0.7) 52.7 (1.0) 69.2 (0.8) 67.0 (0.8) 61.3 (1.2) 43.1 (1.7) 63.2 (1.1) 57.1 (2.1) 43.6 (2.6) 21.8 (2.5)
WoLA+NNM 70.9 (0.6) 69.8 (0.6) 68.3 (0.7) 57.6 (0.9) 69.3 (0.8) 67.8 (0.7) 65.3 (1.1) 53.5 (1.7) 64.0 (1.1) 60.6 (1.7) 52.6 (2.5) 33.0 (3.7)
WoLA 70.5 (0.6) 68.8 (0.5) 65.2 (0.6) 48.2 (1.7) 69.0 (0.8) 66.3 (0.9) 59.3 (1.6) 33.5 (6.1) 62.9 (1.0) 55.9 (1.9) 40.9 (3.4) 11.2 (5.5)
WoLA+NNM 70.7 (0.6) 69.3 (0.5) 67.0 (0.6) 53.3 (2.6) 69.2 (0.7) 67.2 (0.8) 63.1 (1.5) 39.9 (9.9) 63.7 (1.0) 59.3 (1.7) 45.7 (5.2) 19.5 (7.6)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -1.0 +1.3 +5.4 +14.6 -1.4 +3.1 +9.3 +14.8 -6.4 -1.7 +14.6 +6.8
WoLA+NNM -0.8 +1.8 +7.0 +19.5 -1.3 +3.9 +13.3 +25.2 -5.6 +1.8 +23.7 +17.9
WoLA -1.2 +0.8 +4.0 +10.2 -1.6 +2.4 +7.2 +5.2 -6.7 -2.9 +11.9 -3.9
WoLA+NNM -0.9 +1.4 +5.7 +15.3 -1.5 +3.2 +11.1 +11.6 -5.9 +0.5 +16.7 +4.5
(d) CIFAR10
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 69.4 (1.1) 60.4 (2.1) 41.4 (4.2) 21.5 (4.2) 63.9 (1.7) 48.1 (3.3) 32.2 (3.6) 16.1 (3.5) 51.9 (3.5) 36.7 (4.5) 24.5 (4.8) 12.0 (3.5)
BKT 67.9 (1.0) 61.6 (2.0) 41.4 (4.2) 21.5 (4.2) 62.3 (1.8) 51.4 (3.3) 32.2 (3.6) 16.1 (3.5) 52.8 (2.9) 37.5 (3.7) 24.5 (4.8) 12.0 (3.5)
FoundFL 45.1 (12.2) 35.1 (9.9) 32.6 (9.6) 22.5 (8.6) 27.2 (6.8) 20.5 (6.1) 25.2 (5.8) 19.3 (5.7) 20.0 (4.4) 18.2 (4.6) 18.3 (3.9) 15.2 (4.6)
NNM 71.7 (0.8) 66.9 (1.5) 56.7 (3.9) 38.8 (7.0) 68.7 (1.2) 59.1 (3.4) 40.6 (5.2) 29.2 (6.5) 62.7 (2.4) 48.7 (3.7) 32.5 (5.4) 22.5 (3.9)
WoLA 71.6 (0.6) 67.2 (0.8) 58.8 (1.2) 40.0 (2.2) 68.7 (0.7) 63.2 (1.1) 51.9 (2.9) 32.3 (3.9) 62.5 (1.8) 54.0 (3.1) 38.9 (4.7) 20.3 (5.0)
WoLA+NNM 72.7 (0.5) 70.0 (0.8) 64.2 (1.1) 51.6 (2.1) 70.3 (0.6) 66.9 (0.9) 59.1 (2.7) 45.0 (3.8) 65.4 (1.5) 60.1 (3.1) 49.0 (3.9) 34.2 (7.2)
WoLA 71.5 (0.6) 66.4 (0.9) 56.3 (1.6) 34.4 (3.3) 68.4 (0.7) 62.1 (1.5) 46.9 (3.9) 25.6 (6.4) 62.2 (1.9) 52.4 (3.7) 33.8 (5.5) 16.3 (4.8)
WoLA+NNM 72.7 (0.5) 69.4 (0.8) 62.5 (1.2) 45.4 (3.2) 69.9 (0.7) 66.2 (1.1) 55.4 (2.8) 37.5 (8.5) 65.0 (1.6) 58.2 (3.2) 42.7 (4.9) 21.3 (7.5)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.1 +0.3 +2.2 +1.2 -0.1 +4.1 +11.3 +3.1 -0.3 +5.4 +6.4 -2.1
WoLA+NNM +1.0 +3.2 +7.5 +12.9 +1.5 +7.9 +18.5 +15.8 +2.6 +11.4 +16.4 +11.8
WoLA -0.2 -0.5 -0.4 -4.4 -0.3 +3.0 +6.3 -3.6 -0.5 +3.7 +1.3 -6.2
WoLA+NNM +0.9 +2.5 +5.9 +6.7 +1.2 +7.1 +14.7 +8.3 +2.2 +9.5 +10.2 -1.2
Table 3: Worst-case Performance: Test Accuracy (s.d.) % under the worst attack, averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 93.4 (0.3) 91.9 (0.5) 73.6 (5.1) 27.5 (5.8) 93.3 (0.4) 90.1 (1.3) 62.8 (8.1) 12.6 (5.7) 89.3 (3.2) 70.7 (8.6) 29.5 (11.5) 7.1 (3.4)
BKT 93.2 (0.5) 91.5 (0.9) 73.6 (5.1) 27.5 (5.8) 93.2 (0.5) 90.8 (1.5) 62.8 (8.1) 12.6 (5.7) 92.6 (1.1) 84.6 (4.0) 29.5 (11.5) 7.1 (3.4)
FoundFL 80.4 (4.4) 36.5 (8.2) 10.5 (1.2) 10.9 (1.7) 76.4 (15.1) 27.1 (8.7) 10.9 (1.2) 13.5 (3.3) 48.5 (22.4) 20.5 (6.3) 14.7 (4.0) 16.9 (7.6)
NNM 93.6 (0.2) 92.8 (0.3) 91.3 (0.6) 47.0 (27.5) 93.8 (0.3) 91.6 (1.7) 83.9 (8.9) 22.5 (11.7) 93.7 (0.6) 85.6 (3.7) 49.6 (24.0) 18.1 (5.9)
WoLA 93.5 (0.3) 92.9 (0.3) 90.6 (0.4) 66.7 (0.7) 93.8 (0.3) 92.9 (0.3) 89.9 (0.7) 64.3 (3.0) 93.3 (0.5) 92.1 (0.8) 77.6 (8.1) 28.8 (6.2)
WoLA+NNM 93.6 (0.3) 93.2 (0.3) 92.6 (0.2) 70.5 (1.6) 93.9 (0.3) 93.2 (0.3) 92.3 (0.4) 66.7 (4.7) 93.6 (0.5) 92.6 (0.7) 78.9 (11.2) 24.1 (11.4)
WoLA 93.5 (0.3) 92.8 (0.3) 90.6 (0.4) 67.3 (1.0) 93.7 (0.3) 92.8 (0.3) 89.6 (0.8) 63.7 (8.9) 93.3 (0.5) 92.0 (0.8) 75.9 (8.1) 24.5 (7.3)
WoLA+NNM 93.6 (0.3) 93.1 (0.3) 92.3 (0.4) 70.3 (0.7) 93.7 (0.3) 93.1 (0.3) 91.9 (0.5) 51.4 (27.9) 93.6 (0.6) 92.6 (0.8) 80.2 (9.5) 14.9 (6.6)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.1 +0.1 -0.8 +19.7 -0.1 +1.3 +6.0 +41.9 -0.4 +6.5 +28.1 +10.7
WoLA+NNM 0.0 +0.3 +1.3 +23.5 0.0 +1.6 +8.3 +44.3 -0.1 +7.0 +29.3 +6.0
WoLA -0.1 0.0 -0.8 +20.3 -0.1 +1.3 +5.7 +41.3 -0.4 +6.4 +26.3 +6.5
WoLA+NNM +0.1 +0.2 +0.9 +23.2 -0.1 +1.5 +7.9 +28.9 -0.1 +7.0 +30.6 -3.2
(b) Fashion MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 75.6 (0.8) 69.6 (1.3) 50.6 (7.7) 18.3 (8.8) 73.7 (2.6) 64.6 (4.0) 40.8 (8.5) 10.1 (4.7) 66.7 (2.6) 45.2 (5.1) 22.5 (6.1) 4.6 (3.3)
BKT 75.9 (0.8) 71.4 (1.5) 50.6 (7.7) 18.3 (8.8) 76.2 (2.0) 70.2 (2.8) 40.8 (8.5) 10.1 (4.7) 74.5 (3.3) 59.2 (6.6) 22.5 (6.1) 4.6 (3.3)
FoundFL 65.5 (4.5) 35.9 (8.8) 11.7 (1.6) 12.2 (3.9) 53.6 (15.5) 23.6 (10.3) 20.6 (11.9) 15.0 (5.2) 31.4 (12.0) 25.7 (10.1) 15.8 (5.9) 14.8 (7.2)
NNM 77.5 (0.5) 73.3 (1.2) 62.9 (4.1) 25.3 (13.4) 77.9 (1.6) 73.4 (2.3) 54.5 (7.2) 18.8 (8.0) 77.2 (2.2) 69.2 (4.7) 36.6 (12.8) 14.1 (2.7)
WoLA 77.9 (0.6) 75.8 (0.5) 71.8 (1.0) 51.5 (1.5) 78.9 (1.3) 76.2 (1.6) 71.6 (2.2) 46.5 (7.0) 77.0 (1.9) 73.4 (2.5) 58.9 (6.2) 24.7 (7.5)
WoLA+NNM 78.0 (0.7) 76.6 (0.6) 74.8 (0.6) 54.8 (1.7) 79.1 (1.4) 76.9 (1.5) 75.0 (1.6) 49.5 (8.3) 77.9 (1.8) 74.9 (2.6) 61.6 (11.0) 18.9 (7.9)
WoLA 77.9 (0.5) 75.4 (0.4) 70.1 (1.0) 48.1 (2.5) 78.7 (1.3) 75.1 (1.6) 67.2 (3.4) 44.5 (5.1) 76.7 (1.5) 70.1 (2.8) 47.7 (5.9) 18.0 (7.4)
WoLA+NNM 77.8 (0.7) 75.9 (0.4) 73.2 (1.1) 50.8 (2.4) 79.1 (1.6) 75.7 (1.8) 71.1 (2.0) 35.4 (18.6) 77.8 (1.6) 71.8 (2.2) 57.9 (8.2) 11.3 (5.1)
Accuracy Gain Over Top Non-WoLA Methods
WoLA +0.5 +2.5 +8.9 +26.1 +1.0 +2.8 +17.1 +27.7 -0.3 +4.2 +22.3 +9.9
WoLA+NNM +0.5 +3.3 +11.9 +29.4 +1.2 +3.5 +20.5 +30.7 +0.7 +5.6 +25.0 +4.1
WoLA +0.4 +2.2 +7.2 +22.8 +0.9 +1.7 +12.7 +25.7 -0.5 +0.9 +11.1 +3.2
WoLA+NNM +0.3 +2.6 +10.3 +25.4 +1.2 +2.3 +16.5 +16.7 +0.6 +2.5 +21.4 -3.5
(c) Purchase100
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 68.2 (0.6) 59.2 (0.8) 29.0 (1.2) 8.7 (2.3) 64.0 (1.5) 48.5 (2.9) 18.5 (3.1) 0.9 (0.3) 54.9 (2.5) 28.5 (3.8) 2.1 (0.8) 0.9 (0.3)
BKT 68.6 (0.8) 58.1 (0.9) 29.0 (1.2) 8.7 (2.3) 65.8 (1.6) 50.3 (3.1) 18.5 (3.1) 0.9 (0.3) 58.2 (3.4) 30.0 (7.1) 2.1 (0.8) 0.9 (0.3)
FoundFL 33.6 (1.9) 4.2 (1.7) 0.6 (0.3) 0.4 (0.1) 21.5 (6.8) 2.6 (1.3) 1.6 (0.9) 1.6 (0.9) 13.2 (4.7) 3.2 (1.1) 2.9 (0.3) 2.5 (1.3)
NNM 70.4 (0.5) 63.3 (1.5) 52.2 (1.4) 6.4 (1.0) 67.1 (1.7) 53.8 (4.3) 33.6 (4.4) 3.2 (1.8) 61.8 (3.1) 36.9 (8.1) 3.2 (2.3) 1.0 (0.4)
WoLA 69.7 (0.6) 66.7 (0.6) 61.4 (0.7) 37.7 (1.0) 68.0 (0.8) 63.8 (0.8) 49.0 (1.4) 19.3 (1.2) 60.3 (1.0) 47.3 (2.3) 18.2 (2.6) 6.1 (2.8)
WoLA+NNM 70.3 (0.6) 68.0 (0.6) 66.3 (0.6) 44.2 (0.8) 68.7 (0.8) 65.8 (0.6) 62.0 (1.5) 37.6 (1.9) 61.7 (1.0) 53.5 (1.5) 39.7 (3.3) 3.0 (1.8)
WoLA 69.5 (0.6) 66.2 (0.5) 59.8 (0.6) 31.9 (1.9) 67.7 (0.8) 62.9 (0.8) 45.7 (2.5) 21.6 (4.6) 59.9 (0.9) 45.8 (2.0) 23.4 (4.5) 4.2 (1.0)
WoLA+NNM 70.2 (0.6) 67.5 (0.5) 64.8 (0.4) 36.3 (6.4) 68.5 (0.7) 65.0 (0.7) 58.4 (1.6) 20.0 (13.5) 61.4 (0.9) 51.6 (1.6) 17.4 (15.2) 1.9 (0.7)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.7 +3.4 +9.2 +29.0 +0.8 +10.0 +15.4 +16.1 -1.5 +10.4 +15.1 +3.6
WoLA+NNM -0.1 +4.7 +14.1 +35.5 +1.5 +12.0 +28.4 +34.4 -0.1 +16.6 +36.5 +0.5
WoLA -0.9 +2.8 +7.6 +23.2 +0.5 +9.1 +12.1 +18.4 -1.8 +8.9 +20.3 +1.7
WoLA+NNM -0.2 +4.2 +12.6 +27.7 +1.3 +11.2 +24.8 +16.8 -0.4 +14.8 +14.2 -0.6
(d) CIFAR10
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 63.9 (2.9) 50.1 (4.7) 25.4 (8.0) 14.4 (3.5) 55.3 (5.3) 22.2 (5.0) 16.3 (4.1) 10.9 (3.1) 36.3 (5.4) 17.7 (5.5) 15.4 (5.2) 8.2 (2.7)
BKT 65.2 (1.6) 55.6 (3.7) 25.4 (8.0) 14.4 (3.5) 58.3 (3.8) 34.0 (7.3) 16.3 (4.1) 10.9 (3.1) 42.2 (6.3) 18.0 (5.0) 15.4 (5.2) 8.2 (2.7)
FoundFL 36.7 (13.2) 14.0 (4.2) 10.6 (1.8) 11.0 (2.8) 21.6 (6.2) 12.0 (2.5) 11.0 (1.9) 11.3 (3.3) 16.9 (4.0) 12.6 (3.1) 12.7 (2.8) 9.6 (2.1)
NNM 65.0 (2.3) 53.3 (4.6) 42.0 (8.6) 19.2 (8.3) 58.8 (4.1) 31.3 (8.9) 15.2 (4.2) 14.0 (3.0) 45.7 (5.8) 17.9 (6.4) 15.6 (3.7) 11.8 (3.1)
WoLA 69.7 (0.9) 64.0 (1.1) 49.0 (1.5) 23.5 (3.0) 66.0 (1.1) 59.2 (2.1) 39.7 (5.1) 21.7 (2.1) 57.0 (2.2) 46.9 (4.3) 28.4 (3.8) 11.9 (2.5)
WoLA+NNM 70.5 (1.0) 65.2 (1.6) 57.7 (1.4) 34.6 (3.5) 67.1 (1.6) 60.3 (2.3) 49.2 (3.5) 27.3 (2.9) 59.2 (1.6) 49.7 (6.2) 37.1 (7.3) 22.0 (3.2)
WoLA 69.5 (0.9) 62.9 (1.5) 44.1 (2.5) 18.9 (2.5) 65.6 (1.3) 57.4 (2.1) 33.5 (5.4) 13.3 (2.6) 56.8 (2.2) 45.1 (6.2) 21.5 (5.5) 8.4 (4.7)
WoLA+NNM 70.6 (0.9) 64.3 (1.6) 55.0 (1.2) 34.0 (2.7) 66.3 (2.0) 59.5 (2.4) 44.2 (4.1) 26.5 (10.4) 58.7 (1.8) 45.6 (6.0) 28.8 (7.2) 13.8 (5.7)
Accuracy Gain Over Top Non-WoLA Methods
WoLA +4.5 +8.4 +6.9 +4.3 +7.2 +25.2 +23.4 +7.7 +11.3 +28.8 +12.8 +0.1
WoLA+NNM +5.4 +9.5 +15.7 +15.5 +8.3 +26.3 +32.9 +13.4 +13.5 +31.7 +21.5 +10.2
WoLA +4.4 +7.3 +2.1 -0.3 +6.8 +23.4 +17.2 -0.7 +11.1 +27.0 +5.8 -3.4
WoLA+NNM +5.4 +8.6 +13.0 +14.8 +7.5 +25.4 +28.0 +12.5 +13.0 +27.5 +13.1 +2.0
Table 4: Gradient Dissimilarity (s.d.), averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd).
(a) MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 0.1 (0.03) 0.33 (0.09) 1.09 (0.3) 2.07 (0.35) 0.69 (0.15) 1.8 (0.36) 3.27 (0.85) 4.28 (1.18) 4.46 (0.94) 7.04 (1.35) 8.91 (1.55) 4.56 (0.98)
BKT 0.06 (0.01) 0.12 (0.03) 1.09 (0.3) 2.07 (0.35) 0.13 (0.02) 0.36 (0.14) 3.27 (0.85) 4.28 (1.18) 0.55 (0.18) 2.79 (0.67) 8.91 (1.55) 4.56 (0.98)
FoundFL 0.37 (0.13) 0.84 (0.32) 0.59 (0.19) 0.52 (0.27) 0.84 (0.45) 1.28 (0.58) 0.88 (0.46) 0.89 (0.47) 4.54 (1.83) 3.86 (1.83) 3.61 (1.78) 3.45 (1.86)
NNM 0.06 (0.01) 0.08 (0.02) 0.13 (0.03) 0.83 (0.19) 0.12 (0.02) 0.36 (0.24) 0.71 (0.31) 1.92 (0.81) 0.38 (0.08) 1.22 (0.35) 4.88 (1.38) 4.74 (1.02)
WoLA 0.03 (0.0) 0.03 (0.0) 0.06 (0.02) 0.11 (0.02) 0.04 (0.0) 0.06 (0.01) 0.11 (0.03) 0.33 (0.18) 0.07 (0.01) 0.16 (0.05) 0.69 (0.35) 1.68 (0.42)
WoLA+NNM 0.03 (0.0) 0.03 (0.0) 0.04 (0.01) 0.05 (0.01) 0.04 (0.0) 0.05 (0.01) 0.06 (0.01) 0.19 (0.12) 0.05 (0.01) 0.07 (0.01) 0.46 (0.33) 0.81 (0.18)
WoLA 0.03 (0.0) 0.04 (0.0) 0.06 (0.01) 0.1 (0.03) 0.04 (0.01) 0.05 (0.01) 0.13 (0.06) 0.68 (0.68) 0.07 (0.01) 0.15 (0.03) 0.55 (0.26) 2.24 (0.8)
WoLA+NNM 0.03 (0.0) 0.03 (0.0) 0.03 (0.01) 0.06 (0.01) 0.04 (0.0) 0.05 (0.01) 0.06 (0.02) 0.36 (0.29) 0.05 (0.01) 0.07 (0.01) 0.37 (0.25) 1.5 (0.87)
(b) Fashion MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 3.84 (0.56) 4.2 (0.83) 4.22 (0.76) 2.84 (0.62) 8.15 (1.76) 7.67 (1.56) 7.32 (1.5) 4.15 (1.39) 14.01 (1.24) 13.47 (1.37) 11.15 (1.69) 4.55 (1.5)
BKT 1.26 (0.11) 1.82 (0.23) 4.22 (0.76) 2.84 (0.62) 2.45 (0.56) 3.8 (1.17) 7.32 (1.5) 4.15 (1.39) 5.5 (0.84) 8.82 (1.23) 11.15 (1.69) 4.55 (1.5)
FoundFL 1.28 (0.69) 1.31 (0.68) 1.08 (0.45) 0.76 (0.32) 3.01 (1.32) 2.18 (1.18) 2.24 (1.28) 1.92 (1.01) 6.18 (2.86) 5.93 (2.26) 5.7 (2.17) 4.02 (1.96)
NNM 2.54 (0.26) 2.54 (0.34) 2.92 (0.48) 1.92 (0.67) 5.11 (1.32) 5.0 (1.02) 5.26 (1.52) 3.23 (1.12) 9.94 (1.62) 10.1 (1.87) 9.48 (2.09) 6.0 (1.29)
WoLA 0.18 (0.02) 0.17 (0.02) 0.17 (0.03) 0.14 (0.02) 0.41 (0.17) 0.38 (0.15) 0.43 (0.17) 0.29 (0.12) 0.89 (0.15) 1.16 (0.24) 1.57 (0.48) 1.51 (0.7)
WoLA+NNM 0.17 (0.02) 0.15 (0.02) 0.13 (0.02) 0.11 (0.02) 0.38 (0.16) 0.29 (0.11) 0.28 (0.11) 0.28 (0.14) 0.78 (0.12) 0.72 (0.12) 0.97 (0.38) 0.99 (0.33)
WoLA 0.19 (0.02) 0.19 (0.03) 0.18 (0.05) 0.15 (0.04) 0.51 (0.2) 0.51 (0.21) 0.51 (0.25) 0.73 (0.7) 0.95 (0.17) 1.35 (0.31) 1.53 (0.27) 2.17 (0.88)
WoLA+NNM 0.18 (0.02) 0.15 (0.02) 0.14 (0.04) 0.12 (0.03) 0.45 (0.18) 0.4 (0.16) 0.35 (0.16) 0.4 (0.29) 0.82 (0.1) 0.79 (0.13) 0.91 (0.23) 1.46 (0.75)
(c) Purchase100
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 0.14 (0.01) 0.2 (0.02) 0.27 (0.02) 0.42 (0.05) 0.55 (0.08) 0.67 (0.1) 0.99 (0.11) 1.55 (0.23) 2.14 (0.29) 2.84 (0.33) 3.59 (0.43) 3.74 (0.42)
BKT 0.13 (0.01) 0.17 (0.01) 0.27 (0.02) 0.42 (0.05) 0.39 (0.04) 0.53 (0.05) 0.99 (0.11) 1.55 (0.23) 1.29 (0.16) 2.12 (0.28) 3.59 (0.43) 3.74 (0.42)
FoundFL 0.24 (0.02) 0.32 (0.03) 0.28 (0.02) 0.32 (0.03) 0.98 (0.17) 1.01 (0.12) 0.98 (0.15) 1.04 (0.17) 4.29 (0.73) 3.79 (0.68) 3.4 (0.46) 2.94 (0.54)
NNM 0.12 (0.01) 0.15 (0.01) 0.16 (0.02) 0.24 (0.02) 0.34 (0.04) 0.45 (0.05) 0.6 (0.1) 0.99 (0.19) 1.03 (0.14) 1.76 (0.36) 3.03 (0.52) 3.45 (0.49)
WoLA 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.03 (0.0) 0.04 (0.0) 0.04 (0.0) 0.05 (0.01) 0.09 (0.01) 0.12 (0.01) 0.15 (0.02) 0.24 (0.03)
WoLA+NNM 0.02 (0.0) 0.01 (0.0) 0.02 (0.0) 0.01 (0.0) 0.03 (0.0) 0.03 (0.0) 0.04 (0.0) 0.03 (0.0) 0.08 (0.0) 0.09 (0.01) 0.11 (0.01) 0.16 (0.02)
WoLA 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.05 (0.01) 0.03 (0.0) 0.04 (0.0) 0.06 (0.01) 0.13 (0.05) 0.09 (0.0) 0.14 (0.03) 0.26 (0.09) 0.44 (0.19)
WoLA+NNM 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.03 (0.01) 0.03 (0.0) 0.03 (0.0) 0.04 (0.01) 0.11 (0.06) 0.08 (0.0) 0.11 (0.02) 0.23 (0.09) 0.41 (0.19)
(d) CIFAR10
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 1.81 (0.27) 2.49 (0.26) 2.45 (0.33) 1.67 (0.48) 5.06 (0.8) 5.73 (0.88) 4.68 (0.83) 2.12 (0.96) 8.56 (0.7) 9.14 (0.86) 7.11 (1.04) 2.93 (1.25)
BKT 0.8 (0.06) 1.09 (0.08) 2.45 (0.33) 1.67 (0.48) 1.57 (0.2) 2.26 (0.49) 4.68 (0.83) 2.12 (0.96) 2.84 (0.26) 4.12 (0.64) 7.11 (1.04) 2.93 (1.25)
FoundFL 1.36 (0.56) 1.24 (0.48) 1.46 (0.56) 1.08 (0.63) 2.45 (1.0) 1.91 (1.22) 3.39 (1.24) 3.08 (1.34) 3.84 (1.68) 3.6 (1.37) 4.11 (2.21) 2.93 (1.91)
NNM 1.13 (0.11) 1.49 (0.23) 1.36 (0.34) 1.28 (0.4) 2.24 (0.29) 3.11 (0.74) 2.29 (0.76) 1.96 (0.78) 3.39 (0.24) 5.01 (0.92) 3.43 (0.73) 3.07 (0.94)
WoLA 0.47 (0.03) 0.44 (0.03) 0.38 (0.04) 0.36 (0.11) 0.72 (0.05) 0.78 (0.1) 0.68 (0.11) 0.54 (0.14) 1.01 (0.08) 1.23 (0.24) 1.16 (0.38) 0.87 (0.35)
WoLA+NNM 0.44 (0.02) 0.39 (0.02) 0.37 (0.03) 0.43 (0.11) 0.63 (0.04) 0.63 (0.06) 0.65 (0.1) 0.61 (0.13) 0.84 (0.06) 0.99 (0.14) 1.01 (0.23) 0.94 (0.28)
WoLA 0.47 (0.03) 0.44 (0.03) 0.38 (0.04) 0.32 (0.05) 0.72 (0.05) 0.76 (0.1) 0.69 (0.15) 0.54 (0.25) 0.99 (0.09) 1.18 (0.19) 1.18 (0.25) 1.09 (0.47)
WoLA+NNM 0.44 (0.02) 0.39 (0.02) 0.37 (0.04) 0.35 (0.05) 0.63 (0.04) 0.62 (0.05) 0.6 (0.08) 0.53 (0.2) 0.82 (0.07) 0.96 (0.15) 0.92 (0.22) 1.07 (0.47)
7 Experimental Setup
This section describes our experimental setup, including the baseline methods, the WoLA approach under training objective attack, and the datasets used. We detail the configuration of the heterogeneity regimes, workers, and the types of Byzantine attacks we consider. We also present the robust aggregation methods, evaluation metrics, optimization algorithms, models, and hyperparameters we used. Finally, we explain how we ensure the reproducibility of our experiments.
7.1 Baseline Approaches
We compare WoLA against three baselines: Bucketing (BKT) [24], NNM [3], and FoundationFL (FoundFL) [15]. BKT requires setting a bucket size $s$. Following the authors’ recommendation, we use $s=\left\lfloor{n}/({2f})\right\rfloor$. When $f>n/4$, the bucket size drops to $1$, making BKT ineffective. For instance, with $17$ workers, BKT has no effect when $f=6$ or $8$. FoundFL requires a hyperparameter $m$. As suggested in the original paper, we set $m=n/2$, i.e., $m=8$ in all our experiments. We also include a baseline using only a robust aggregator without any additional defense, denoted as -.
We evaluate different variants of WoLA. In particular, to assess robustness against training objective attacks, we implement the worst-case scenario described in Subsection 7.2, denoting the resulting loss as WoLA. Additionally, we examine the combination of WoLA and WoLA with NNM.
We do not compare WoLA with methods that (i) require a server-side dataset [4, 28, 11, 10, 30], (ii) involve high computational costs [4, 12, 32, 47, 39], or (iii) rely on hyperparameter tuning [46, 35].
7.2 WoLA under Training Objective Attack
To define the training objective, we consider a scenario where workers share their local label distributions with the server. In this case, Byzantine machines may try to manipulate the objective by sending malicious distributions. We analyze the impact of such attacks.
Just as Byzantine workers can send arbitrary updates with full knowledge of honest gradients, we assume they can also send arbitrary distributions with full knowledge of local label distributions. Rather than listing all possible attacks, we focus on one that causes the maximum distortion: Byzantine machines send a zero vector except for a single entry set to $1$, corresponding to the minority class in the global dataset (proof are provided in Appendix A.7).
To defend against attacks, we propose using the geometric median as the aggregator. This approach offers two key advantages: (i) Unlike methods like Trimmed Mean or Krum, which exclude outliers, the geometric median treats all updates equally. This is important because outlier distributions are not necessarily malicious. Excluding them may discard honest distribution and degrade performance. (ii) When aggregating probability distributions, the geometric median produces a probability distribution—unlike the coordinate-wise median, for instance—ensuring the training objective retains its expected form.
Our goal is to show that a strong attack, even when countered by a simple defense, has limited impact on model performance. As the chosen attack represents a worst-case scenario, it provides a meaningful benchmark for evaluating training objective attacks.
We refer to our approach in this scenario as WoLA.
7.3 Datasets, Heterogeneity & Workers
We conduct the experiments on four datasets: MNIST [27], Fashion-MNIST [42], CIFAR10 [26], and Purchase100 [37]. Purchase100 contains shopping records for thousands of customers. We use a simplified version [37], where each customer is represented by $600$ binary features indicating item purchases. The data is clustered into $100$ classes, and the task is to predict each customer’s class.
To simulate data heterogeneity across workers, we sample from the global dataset using a Dirichlet distribution with concentration parameter $\alpha$, following the method in [3, 23]. We consider three heterogeneity levels—low, medium, and high—corresponding to $\alpha\in\{3,1,0.3\}$.
We use $n=17$ workers in all experiments. The number of Byzantine workers varies as $f\in\{2,4,6,8\}$. Each worker holds a local dataset of fixed size, equal to the total dataset size divided by the number of honest workers.
7.4 Byzantine Attacks & Robust Aggregators
Byzantine workers carry out one of five state-of-the-art gradient attacks: Fall of Empires (FOE) [43], A Little is Enough (ALIE) [5], Sign Flipping (SF) [2], Label Flipping (LF) [2], and Mimic [24].
In the Mimic attack, all Byzantine workers replicate a selected honest update and send it to the server. Inspired by the ALIE attack—which places Byzantine values between the honest mean and outliers to overcome robustness defenses—we instead select the honest outlier that is most surrounded by other honest updates (see Appendix for details). With this strategy, the “extreme” honest outliers hide Byzantine updates, which makes the attack harder to detect while remaining effective in heterogeneous settings. For instance, based on the extensive results in Appendix B.4, Tables 14, 15, and 16, Mimic often appears as the worst attack for the strategies –, BKT, and NNM.
All state-of-the-art strategies are evaluated using four robust aggregators: CWMed, CwTM, GM, and MKrum.
7.5 Metrics, Tables & Boxplots
We use two metrics: (i) Test Accuracy, averaged over training with a test set that matches the global training label distribution, and (ii) Gradient Dissimilarity, also averaged over training.
We ran experiments on four datasets, three levels of heterogeneity, four aggregation methods, and five attacks, averaging results over three random seeds for each scenario. Full results are in Appendix B.4. For clarity, we summarize key findings in three tables. Table 2 shows Test Accuracy, averaged over all aggregation methods (CWMed, CwTM, GM, MKrum) and all attacks (ALIE, FOE, LF, Mimic, SF), to reflect overall performance. Table 4 presents the same summary for Gradient Dissimilarity. Table 3 shows Test Accuracy under the strongest attack (i.e., the one causing the lowest performance), averaged over all aggregators, to highlight worst-case performance.
In Tables 2 and 3, we report the Accuracy Gain Over Top Non-WoLA Methods—the difference between each WoLA variant and the best-performing non-WoLA method (among –, BKT, FoundFL, and NNM). These values highlight performance gains over state-of-the-art strategies.
We also include boxplots of the gains from Tables 2 and 3 in Figures 4(a) and 4(b), to provide an overall view. In the following, when describing a given situation, we refer to the typical performance as the interquartile range of the corresponding boxplot. The average gain is also taken from these boxplots.
7.6 Optimization Algorithm
WoLA relies on approximating the global class gradients using local datasets (see Subsection 5.3). Approximation accuracy improves with increased data per class at each optimization step. As a result, the best performance is likely achieved with non-stochastic gradient descent, where honest workers compute gradients over the entire dataset. In practice, however, stochastic methods are more common—workers typically compute gradients on random mini-batches. To match this setting, we follow [3] and adopt the distributed stochastic heavy ball method (see Algorithm in Appendix B.1). This method has demonstrated strong robustness against Byzantine behavior in both heterogeneous [3] and homogeneous settings [13, 25, 16].
We use momentum $\beta=0.9$ for MNIST and Fashion-MNIST and $0.99$ for CIFAR10 and Purchase100.
7.7 Models & Hyperparameters
On MNIST and Fashion-MNIST, we train a lightweight CNN for $T=800$ steps. On CIFAR10, we use a deeper CNN; on Purchase100, a fully connected neural network. Both models are trained for $T=2000$ steps. The model architectures are provided in Appendix B.2. Unless otherwise specified, the loss is cross-entropy. The batch size is $b=128$, learning rates decay over time, gradient clipping is fixed at $5$, and $\ell_{2}$ regularization is set to $10^{-4}$.
7.8 Reproducibility
All experiments are run using the three seeds $\{1,2,3\}$ to ensure reproducibility. The code will be made publicly available for reuse at https://anonymous.4open.science/r/WoLA-E4CB.
8 Empirical Results
In this section, we present experiments to answer the following Research Questions (RQ). We use WoLA methods to refer collectively to WoLA, WoLA, and their combinations with NNM.
  • RQ1
    Does WoLA guide training toward a better model?
  • RQ2
    Are WoLA and WoLA more robust overall than state-of-the-art strategies?
  • RQ3
    Are WoLA and WoLA more robust in worst-case scenarios than state-of-the-art strategies?
  • RQ4
    Does combining WoLA or WoLA with NNM improve robustness?
  • RQ5
    What explains the robustness of WoLA methods?
  • RQ6
    How do batch size and momentum affect WoLA performance?
[…]
8.1 Targeted Model with WoLA (RQ1)
As expected from Proposition 3 (Subsection 6.2), WoLA does not target a better model during training.
Figure 5 shows Test Accuracy over a training without Byzantine workers, using simple mean aggregation. The model trained with WoLA performs similarly to the unweighted loss, though slightly worse on Purchase100 under high heterogeneity.
8.2 Robustness of WoLA and WoLA (RQ2)
WoLA and WoLA significantly improve robustness. They perform slightly below the best state-of-the-art methods when $f=2$, offer marginal gains at $f=4$, and outperform all non-WoLA methods as the number of Byzantine workers increases.
When $f=2$: WoLA and WoLA show similar performance, slightly below the best non-WoLA strategies. Typical negative gains between $-1.1$ and 00 points (Figure 4(a)). Notable exceptions exist, such as Purchase100 with high heterogeneity (Table 2, $\alpha=0.3$), where NNM outperforms WoLA and WoLA by more than $6$ points.
When $f=4$: WoLA and WoLA still perform similarly and yield slightly better results than non-WoLA methods, with typical gains of a few points (Figure 4(a)). According to Table 2, gains range from $+5.4$ (CIFAR10, $\alpha=0.3$) to $-2.9$ (Purchase100, $\alpha=0.3$).
When $f=6$: WoLA slightly surpasses WoLA, and both outperform other methods. WoLA shows typical gains between $+3.3$ and $+12.1$ points, while WoLA reaches between $+1.8$ and $+8.1$ (Figure 4(a)). Table 2 reports maximum gains of $+21.4$ (WoLA) and $+20.5$ (WoLA) on MNIST with $\alpha=0.3$.
When $f=8$: WoLA is better than WoLA, and both methods clearly outperform all others (Figure 4(a)). WoLA typically achieves gains between $+5.6$ and $+14.7$ points (median around $+10.1$, Figure 4(a)), while WoLA ranges from $-3.7$ to $+7$ (median around $+5.2$, Figure 4(a)). The best results appear on MNIST with medium heterogeneity ($\alpha=1$): $+17.3$ for WoLA and $+13.6$ for WoLA (Table 2).
Importantly, when $f>2$, Table 2 shows that WoLA almost always ranks in the top three regardless of the scenario (exceptions: Purchase100 $\alpha=0.3$, $f=4$ and CIFAR10 $\alpha=0.3$, $f=8$). While WoLA does not frequently appear among the top values, it outperforms non-WoLA strategies most of the time on MNIST and Fashion MNIST, often on Purchase100, and occasionally on CIFAR10.
8.3 Worst-Case Robustness of WoLA and WoLA (RQ3)
Considering the strongest attack for each approach, WoLA and WoLA consistently enhance robustness, almost always outperforming other state-of-the-art methods, with gains reaching up to $+41.9$ points (Table 3, MNIST, $\alpha=1$, $f=8$, WoLA).
For $f=2$: WoLA and WoLA perform similarly, slightly surpass non-WoLA methods with typical gains of a few points (Figure 4(b)), and up to $+11.3$ points on CIFAR10, $\alpha=0.3$ (Table 3).
For $f=4$: WoLA and WoLA continue to perform similarly, outperforming all baselines with median gains around $+6.3$ and $4.6$ points respectively (Figure 4(b)). According to Table 3, WoLA’s gains range from $+0.1$ (MNIST, $\alpha=3$) to $+28.8$ (CIFAR10, $\alpha=0.3$), while WoLA ranges from 00 (MNIST, $\alpha=3$) to $+27$ (CIFAR10, $\alpha=0.3$).
For $f=6$ or $8$: WoLA outperforms WoLA, and both clearly surpass all other methods (Figure 4(b)). WoLA achieves median gains around $+13.4$ points, WoLA around $+12.4$ (Figure 4(b)). Table 3 shows that WoLA’s gains range from $-0.8$ (MNIST, $\alpha=3$, $f=6$) to $+41.9$ (MNIST, $\alpha=1$, $f=8$); WoLA ranges from $-3.4$ (CIFAR10, $\alpha=0.3$, $f=8$) to $+41.3$ (MNIST, $\alpha=1$, $f=8$).
8.4 Robustness Gains from Combining WoLA and WoLA with NNM (RQ4)
Combining WoLA or WoLA with NNM significantly improves robustness, consistently outperforming WoLA, WoLA or NNM alone.
WoLA+NNM consistently achieves the top or top-three performance in Table 2 as soon as $f>2$, reflecting strong robustness across settings. Gains over the best non-WoLA methods can reach around $20$ points (Table 2). For example, on MNIST with $\alpha=0.3$ and $f=8$, the gain is $+26.6$; on Fashion MNIST with $\alpha=0.3$ and $f=6$, $+19.8$; on Purchase100 with $\alpha=1$ and $f=8$, $+25.2$; and on CIFAR10 with $\alpha=1$ and $f=6$, $+18.5$. The median gain is $+10.4$ when $f=6$ and $+17.8$ when $f=8$ (Figure 4(a)). Considering worst attacks, improvements are even larger, with median gains around $+20$ for both $f=6$ and $f=8$ (Figure 4(b)).
WoLA+NNM also regularly reaches top-three values in Tables 2 and 3, confirming its robustness. It performs similarly to WoLA+NNM when $f\leq 4$, and slightly worse when $f>4$, but still outperforms WoLA alone and all non-WoLA strategies (Figure 4(a), 4(b)). Importantly, it turns nearly all negative gains of WoLA alone into positive ones (Table 2). For instance, on CIFAR10 with $\alpha=3$ and $f=8$, the score improves from $-4.4$ to $+6.7$; with $\alpha=1$ and $f=8$, from $-3.6$ to $+8.3$; and with $\alpha=0.3$ and $f=2$, from $-0.5$ to $+2.2$. Overall, median gains over non-WoLA methods are around $+8$ when $f=6$ or $8$ (Figure 4(a)). Considering worst attacks, the median gain is around $+13$ for both $f=6$ and $f=8$ (Figure 4(b)).
8.5 Sources of Robustness in WoLA Methods (RQ5)
Gradient dissimilarity is a key indicator of system robustness [24, 3]. As expected from Proposition 2 (Subsection 6.2), WoLA methods significantly reduce gradient dissimilarity, thereby enhancing robustness against Byzantine attacks.
Table 4 shows that WoLA and its variants consistently achieve much lower gradient dissimilarity across all scenarios. Interestingly, combining NNM with WoLA further reduces gradient dissimilarity, which may explain its greater robustness over WoLA alone.
Figure 3 shows the evolution of Test Accuracy and Gradient Dissimilarity during training on Fashion-MNIST with high heterogeneity ($\alpha=0.3$) and six Byzantine workers ($f=6$) under two attacks (ALIE and LF). We also observe that WoLA’s methods maintain lower gradient dissimilarity than other state-of-the-art approaches.
(a) Fashion MNIST
(b) CIFAR10
Figure 6: Heatmaps of mean Test Accuracy over training (averaged over three seeds). Each row corresponds to a different level of data heterogeneity: low ($\alpha=3$), medium ($\alpha=1$), and high ($\alpha=0.3$). Each column shows: WoLA accuracy, NNM accuracy, and WoLA’s accuracy gain over NNM. In each heatmap, the vertical axis represents batch size, and the horizontal axis represents momentum. Both scenarios involve $17$ workers ($n=17$), including $5$ Byzantine workers ($f=5$), under the ALIE attack mitigated by the CwTM aggregator. Accuracy is reported in %, with a shared color scale for WoLA and NNM accuracy within each row.
8.6 Impact of Batch Size & Momentum on WoLA (RQ6)
Figure 6 reports the mean Test Accuracy on Fashion MNIST and CIFAR-10 achieved by WoLA under three heterogeneity levels ($\alpha=3$, $1$, and $0.3$), with $17$ workers including $5$ Byzantine nodes ($f=5$), using the ALIE attack and the CwTM aggregator. The figure also shows the performance of NNM and the accuracy gain of WoLA over NNM.
For Fashion MNIST, regardless of the heterogeneity level or momentum value, increasing the batch size consistently improves WoLA’s performance. Similarly, increasing the momentum coefficient improves performance up to $0.9$.
For CIFAR-10, under high heterogeneity ($\alpha=0.3$), increasing the batch size (up to $256$) or momentum coefficient improves performance. Under low and medium heterogeneity ($\alpha=3$ and $1$), the effects are less consistent. However, with high momentum ($0.9$ or $0.99$), increasing the batch size still tends to help. The best performance is observed with a high momentum coefficient ($0.99$).
Comparing WoLA and NNM, we find that the best results are obtained with moderately large batch sizes and high momentum. These conclusions are specific to the given setting and are not a substitute for a full-scale study, but they provide useful insights into parameter sensitivity.
9 Conclusion and Future Work
We show that label skew heterogeneity can be partly addressed with WoLA, a weighted loss where weights are set according to each worker’s local label distribution and a shared training objective. WoLA is a simple, model-agnostic method that requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine workers, and adds no computational overhead. It is easy to implement and can be combined with other robustness strategies, such as pre-aggregators and robust aggregators.
We provide theoretical insights into WoLA’s behavior, supported by empirical evaluation. Extensive experiments demonstrate WoLA’s efficiency and its superiority over state-of-the-art strategies—both on average and in worst-case scenarios (i.e., under the strongest attack for each method)—especially in adversarial settings involving both Byzantine behavior and heterogeneity. Even better results are achieved when WoLA is combined with NNM. We also show that WoLA significantly reduces gradient dissimilarity among honest workers, which is key to its robustness.
For future work, WoLA’s robustness can be improved and its scope extended—for instance, by refining the weighting scheme, such as introducing thresholds or adapting it dynamically at the mini-batch scale. The impact of training objective attacks can also be explored more thoroughly, along with the design of stronger defenses against them. Finally, developing a decentralized version of WoLA also seems to be a relevant direction.
Appendix A Theoretical Results
In this section, we define the random counterparts of the local updates and provide all proofs of the theoretical results given in Section 6.2.
A.1 Gradients Modeling
In this subsection, we define the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$. Let the global dataset be $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ with $(x_{i},y_{i})\in\mathcal{X}\times[C]$, and let $z_{1},z_{2},\ldots,z_{N}$ be integers in $\mathcal{H}$ such that $z_{k}=i$ if $(x_{k},y_{k})\in\mathcal{D}_{i}$.
Since $N_{i}=\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)$, the following holds for $\nabla\mathcal{L}_{i}$:
$\nabla\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\ell\left(y,\Phi\left(x\right)\right)$
$=$ $\frac{N}{N_{i}}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\ell\left(y_{k},\Phi\left(x_{k}\right)\right)}{N}$
$=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}$
Let $L_{i}$ be the random counterpart of $\nabla\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $L_{i}$ is given by:
$L_{i}=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (5)
$\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
We now turn to $\nabla\mathcal{WL}_{i}$:
$\nabla\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\frac{q^{y_{k}}}{p^{y_{k}}_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\sum_{c=1}^{C}\mathbbm{1}(y_{k}=c)\frac{q^{c}}{N_{i}^{c}/N_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\sum_{c=1}^{C}q^{c}\frac{N}{N_{i}^{c}}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}}{N}$
Since $N^{c}_{i}=\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)$, we have:
$\nabla\mathcal{WL}_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}\Bigg{)}$
Let $W_{i}$ be the random counterpart of $\nabla\mathcal{WL}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $W_{i}$ is given by:
$W_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (6)
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
A.2 Lemmas
In this subsection, we give two lemmas that will be used in the subsequent theoretical proofs.
Lemma 2
Let $X$, $Y$, and $Z$ be random variables such that $X\mid Y$ is independent of $Z\mid Y$. Let $f$ be a function defined on the space of realizations of $X$. Then, for all $y$ and $z$ such that $\mathbb{P}(Y=y)>0$ and $\mathbb{P}(Z=z)>0$, we have:
$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
Proof.
$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=y)\mathbbm{1}(Z=z)f(X)\Big{]}}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\Big{|}Y=y\Big{]}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\mathbbm{1}(Y=y)\Big{]}}{\mathbb{P}(Y=y)}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
Lemma 3
The following expression converges almost surely (a.s.) as $N\to\infty$:
$\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
[…]
Proof.
By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
$\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
By the Law of Large Numbers, as $N\to\infty$, the expressions converge almost surely to:
$\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\Big{]}^{-1}$
$\quad\quad\quad\quad\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}}{\mathbb{P}(Y=c,Z=i)}$
$=$ $\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}$
According to Lemma 2:
$\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
A.3 Proof of Proposition 1
We begin by calculating the limit of $L_{i}$. By definition, we have:
$L_{i}=\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
By the Law of Large Numbers, it holds:
[…]
$\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
[…]
By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
$\lim_{N\to\infty}L_{i}=\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
$=$ $\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
[…]
Similarly, for $L$:
$\lim_{N\to\infty}L=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}L_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}L_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
[…]
By the Law of total expectation, it holds:
[…]
A.4 Proof of Proposition 2
We begin by calculating the limit of $W_{i}$. By the continuity of the involved functions, it holds:
$\lim_{N\to\infty}W_{i}$
$=$ $\lim_{N\to\infty}\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=$ $\sum_{c=1}^{C}q^{c}\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
Using Lemma 3, as $N\to\infty$, the expression converge almost surely to:
[…]
Similarly, for $W$:
$\lim_{N\to\infty}W=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}W_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}W_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
By the Law of Large Numbers, it holds:
[…]
Using above results, it concludes the lemma:
[…]
$=$ $\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
A.5 Proof of Gradient Dissimilarity Vanishing
According to Proposition 2, by the continuity of the gradient dissimilarity formula, we obtain:
$\lim_{N\to\infty}\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\left\|W_{i}-W\right\|^{2}$
[…]
A.6 Proof of Proposition 3
According to Proposition 2, it holds:
[…]
By assumption, it holds $\tilde{p}_{c}=\mathbb{P}(Y=c)$ for all $c\in[C]$. Then, using the Law of total expectation, it follows:
[…]
$=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
A.7 Worst Training Objective Attack
We show that the training objective $q$ is maximally deviated by a Byzantine attack that concentrates all its mass on the class with minimal weight among honest workers, under the $L_{1}$ norm denoted by $\|\cdot\|_{1}$.
Let $u$ and $v$ be two probability distributions over $C$ classes, with $u=(u^{1},u^{2},\ldots,u^{C})$ such that $u^{i}\in[0,1]$ and $\sum_{i=1}^{C}u^{i}=1$, and similarly for $v$. Let $u^{\min}$ and $u^{\max}$ denote the minimum and maximum components of $u$, and $v^{\max}$ the maximum of $v$.
We first recall the identity:
$\|u-v\|_{1}=\sum_{i=1}^{C}|u^{i}-v^{i}|$ $=\sum_{i=1}^{C}u^{i}+v^{i}-2\min(u^{i},v^{i})$
$=2-2\sum_{i=1}^{C}\min(u^{i},v^{i}).$
Observe that $v^{\max}\geq u^{\min}$. Otherwise, we would have $v^{i}<u^{i}$ for all $i$, which implies $\sum_{i}v^{i}<\sum_{i}u^{i}=1$, contradicting the assumption that $v$ is a probability distribution. It follows that:
$u^{\min}\leq\min(u^{\min},v^{\max})$ $\leq\min(u^{\max},v^{\max})$
$\leq\sum_{i=1}^{C}\min(u^{i},v^{i}),$
which gives the upper bound:
$\|u-v\|_{1}\leq 2-2u^{\min}.$
Now, consider the training objective $q=\sum_{i=1}^{n}q_{i}$, where each $q_{i}$ is a probability distribution sent by worker $i$. Let $u=\frac{1}{n-f}\sum_{i\in\mathcal{H}}q_{i}$ be the average contribution from honest workers, and $v=\frac{1}{f}\sum_{i\notin\mathcal{H}}q_{i}$ the average contribution from Byzantine workers. Then:
$\|q-u\|_{1}=\frac{f}{n}\|v-u\|_{1}\leq\frac{f}{n}(2-2u^{\min}).$
This bound is tight: it is achieved by setting each $q_{i}$ for $i\notin\mathcal{H}$ to put all mass on the class where $u$ is minimal, completing the proof.
Appendix B Experiment
B.1 Robust Distributed Stochastic Heavy Ball Algorithm
We present here the robust distributed stochastic Heavy Ball algorithm. More details can be found in [3]. We use superscript in brackets to denote the iteration index, i.e., (t), and subscripts to indicate the worker index.
Input: Initial model $\theta^{(0)}$, initial momentum $m_{i}^{(0)}=0$ for honest workers, robust aggregation $F$, learning rate $\gamma$, momentum coefficient $\beta$, and number of steps $T$.
for $t=1,\ldots,T$ do
       Server broadcasts $\theta^{(t-1)}$ to all workers;
       for every honest worker $i$ do
             Compute a stochastic gradient $g^{(t)}_{i}$;
             Update local momentum $m_{i}^{(t)}\leftarrow\beta m_{i}^{(t-1)}+(1-\beta)g^{(t)}_{i}$;
             Send $m_{i}^{(t)}$ to the server;
       end for
      Server aggregates the momentums:
           $R^{(t)}=F(m_{1}^{(t)},\ldots,m_{n}^{(t)})$;
       Server updates the model: $\theta^{(t)}=\theta^{(t-1)}-\gamma R^{(t)}$;
end for
Algorithm 1 Robust Distributed Stochastic Heavy Ball
B.2 Model Architectures
We adopt the following abreviations : L(#outputs) denotes a fully connected linear layer; R is a ReLU activation; T is a Tanh activation; C(#channels) represents a 2D convolutional layer with kernel size $5$, padding 00, and stride $1$; M denotes 2D max-pooling with kernel size $2$; B stands for batch normalization; and D represents dropout with a fixed probability of $0.25$.
The model architectures are as follows:
  • MNIST and Fashion MNIST (CNN from [3]): C(20)-R-M–C(20)-R-M–L(500)-R–L(10)
  • CIFAR10 (CNN from [3]):(3,32×32)-C(64)-R-B-C(64)-R-B-M-D-C(128)-R-B-C(128)-R-B-M-D-L(128)-R-D-L(10)
  • Purchase100 (fully connected neural networks from [31]): L(1024)-L(512)-T-L(256)-T-L(100)
The learning rates are defined as:
  • MNIST and Fashion MNIST (from [3]): $\gamma_{t}=0.75/(1+\left\lfloor\frac{t}{50}\right\rfloor)$
  • CIFAR10 (from [3]) and Purchase100: $\gamma_{t}=0.25$ if $t\leq 1500$, $0.025$ otherwise.
B.3 Mimic with Most Surrounded Outlier Selection
In the Mimic attack, we choose the honest outlier that is most surrounded by other honest updates. The selection procedure is detailed below. We first introduce the operators $\operatorname{dist}$ and $\operatorname{rank}$ used in this process.
Let $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}$ be the honest update vectors. For each vector $\mathbf{v}_{i}$, define $\sigma(i,\cdot)$ as the permutation that orders the vectors in decreasing order of their Euclidean distance from $\mathbf{v}_{i}$:
$\operatorname{dist}(\mathbf{v}_{i},\mathbf{v}_{\sigma(i,1)})\geq\cdots\geq\operatorname{dist}(\mathbf{v}_{i},\mathbf{v}_{\sigma(i,h)}),$
where $\operatorname{dist}$ denotes the Euclidean distance. For any vector $\mathbf{v}_{j}$, there exists an index $r$ such that $\sigma(i,r)=j$, meaning that $\mathbf{v}_{j}$ is the $r$‑th farthest vector from $\mathbf{v}_{i}$. Note that $\sigma(i,h)=i$. We define the operator $\operatorname{rank}$ as:
$\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})=r,$
where $j=\sigma(i,r)$.
The operators $\operatorname{dist}$ and $\operatorname{rank}$ allow us to compute a score for each update based on how surrounded a vector is and how far it is from the mean of the honest updates. The update to mimic is the one with the highest score, see Algorithm 2.
Input: Honest updates $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}\in\mathbb{R}^{d}$, number of Byzantine workers $f$
Output: Update $\mathbf{v}_{\text{argmax}}$ with highest score
for $i=1,\ldots,h$ do
       $s_{i}\leftarrow 0$ ;
end for
$\textit{center}\leftarrow\frac{1}{h}\sum_{i=1}^{h}\mathbf{v}_{i}$ ;
for $i=1,\ldots,h$ do
       for $j=1,\ldots,h$ do
             $r\leftarrow\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})$ ;
             $s_{j}\leftarrow s_{j}+\min(f,r)\cdot\operatorname{dist}(\mathbf{v}_{j},\textit{center})$ ;
       end for
end for
$k\leftarrow\arg\max_{j}s_{j}$ return $\mathbf{v}_{k}$
Algorithm 2 Mimic with most surrounded outlier selection
B.4 Comprehensive Results
We report the full experimental results. For MNIST, see Tables 5, 6, and 7; for Fashion MNIST, see Tables 8, 9, and 10; for Purchase100, see Tables 11, 12, and 13; and for CIFAR10, see Tables 14, 15, and 16. Each set of three tables corresponds to low, medium, and high heterogeneity levels ($\alpha=3$, $1$, and $0.3$).
Table 5: Accuracy (s.d.) % on MNIST with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 93.7
93.2
92.4
84.8
93.3
92.5
91.4
85.5
93.6
92.8
88.9
52.7
93.6
93.1
92.6
91.2
93.2
92.5
91.6
80.5
93.5
92.8
91.4
78.9
93.2
92.5
88.9
52.7
BKT 93.7
93.6
92.4
84.8
93.6
93.3
91.4
85.5
93.3
92.1
88.9
52.7
93.7
93.4
92.6
91.2
93.5
92.8
91.6
80.5
93.6
93.0
91.4
78.9
93.3
92.1
88.9
52.7
FoundFL 93.0
93.3
93.1
92.5
91.0
85.8
12.1
12.3
90.1
73.2
50.7
42.9
92.8
92.5
93.3
93.3
90.7
82.5
10.2
11.0
91.5
85.5
51.9
50.4
90.1
73.2
10.2
11.0
NNM 93.8
93.5
93.2
93.7
93.7
93.4
92.7
77.4
93.8
93.5
93.6
93.3
93.6
93.3
92.3
90.6
93.6
92.7
91.7
61.0
93.7
93.3
92.7
83.2
93.6
92.7
91.7
61.0
WoLA 93.7
93.5
93.1
91.5
93.5
93.0
92.1
90.8
93.8
93.6
93.3
91.0
93.7
93.6
93.3
93.5
93.5
92.9
92.0
90.8
93.6
93.3
92.8
91.5
93.5
92.9
92.0
90.8
WoLA+NNM 93.8
93.6
93.4
93.8
93.7
93.4
93.0
90.9
93.8
93.7
93.8
93.8
93.8
93.6
93.4
93.4
93.6
93.2
92.6
92.2
93.7
93.5
93.2
92.8
93.6
93.2
92.6
90.9
WoLA 93.7
93.4
92.8
90.9
93.5
93.0
91.9
90.5
93.7
93.5
92.8
90.5
93.7
93.5
93.3
92.7
93.4
92.8
91.8
90.7
93.6
93.2
92.5
91.1
93.4
92.8
91.8
90.5
WoLA+NNM 93.8
93.4
93.1
93.2
93.7
93.3
92.9
90.5
93.8
93.6
93.4
92.9
93.7
93.5
93.3
92.7
93.6
93.0
92.3
91.0
93.7
93.4
93.0
92.1
93.6
93.0
92.3
90.5
CwTM - 93.7
93.5
92.7
84.8
93.6
93.2
91.3
85.5
93.6
92.6
88.6
54.1
93.5
93.2
92.5
90.2
93.6
92.6
91.4
80.7
93.6
93.0
91.3
79.0
93.5
92.6
88.6
54.1
BKT 93.7
93.6
92.7
84.8
93.6
93.3
91.3
85.5
93.3
92.1
88.6
54.1
93.6
93.3
92.5
90.2
93.5
92.8
91.4
80.7
93.6
93.0
91.3
79.0
93.3
92.1
88.6
54.1
FoundFL 92.6
92.9
92.7
92.3
92.7
91.6
90.4
19.6
72.7
48.9
26.0
16.6
93.0
92.8
93.1
92.9
90.9
24.5
10.8
11.2
88.4
70.1
62.6
46.5
72.7
24.5
10.8
11.2
NNM 93.8
93.6
93.1
93.7
93.7
93.4
92.7
75.8
93.8
93.5
93.7
93.3
93.6
93.3
92.8
91.0
93.6
92.8
91.7
59.6
93.7
93.3
92.8
82.7
93.6
92.8
91.7
59.6
WoLA 93.8
93.5
93.2
91.4
93.7
93.3
92.5
90.7
93.7
93.5
93.2
91.0
93.7
93.6
93.5
93.4
93.5
92.9
91.9
90.9
93.7
93.4
92.9
91.5
93.5
92.9
91.9
90.7
WoLA+NNM 93.8
93.6
93.3
93.8
93.7
93.4
93.0
91.0
93.8
93.7
93.8
93.8
93.8
93.6
93.5
93.4
93.6
93.2
92.6
92.2
93.7
93.5
93.2
92.8
93.6
93.2
92.6
91.0
WoLA 93.7
93.5
92.9
90.9
93.6
93.2
92.3
90.5
93.7
93.4
92.7
90.3
93.7
93.5
93.4
92.7
93.5
92.9
91.7
90.4
93.7
93.3
92.6
90.9
93.5
92.9
91.7
90.3
WoLA+NNM 93.8
93.4
93.1
93.3
93.7
93.3
92.9
90.4
93.8
93.6
93.4
92.9
93.7
93.5
93.4
92.7
93.6
93.1
92.3
91.4
93.7
93.4
93.0
92.1
93.6
93.1
92.3
90.4
MKrum - 93.8
93.6
93.1
78.3
93.7
93.4
92.6
57.0
93.8
93.4
43.3
1.7
93.6
93.1
92.6
90.6
93.5
92.3
30.6
10.1
93.7
93.1
70.4
47.5
93.5
92.3
30.6
1.7
BKT 93.7
93.3
93.1
78.3
93.4
92.9
92.6
57.0
92.4
89.0
43.3
1.7
93.5
92.9
92.6
90.6
93.4
91.5
30.6
10.1
93.3
91.9
70.4
47.5
92.4
89.0
30.6
1.7
FoundFL 92.5
92.7
91.1
91.0
92.4
92.2
85.5
27.1
69.9
52.2
33.9
14.5
92.9
92.3
91.8
92.0
86.3
10.7
10.1
10.6
86.8
68.0
62.5
47.0
69.9
10.7
10.1
10.6
NNM 93.7
93.6
93.1
83.1
93.7
93.5
92.7
71.6
93.8
93.5
93.6
8.0
93.7
93.2
92.2
90.7
93.6
92.7
90.6
10.1
93.7
93.3
92.5
52.7
93.6
92.7
90.6
8.0
WoLA 93.8
93.6
93.3
91.0
93.7
93.3
92.7
76.8
93.8
93.6
93.8
1.7
93.7
93.5
93.4
93.4
93.6
92.9
87.5
10.1
93.7
93.4
92.2
54.6
93.6
92.9
87.5
1.7
WoLA+NNM 93.8
93.6
93.3
91.8
93.7
93.4
93.0
89.1
93.8
93.7
93.8
9.3
93.8
93.6
93.4
93.4
93.6
93.2
92.6
10.4
93.7
93.5
93.2
58.8
93.6
93.2
92.6
9.3
WoLA 93.8
93.4
93.1
90.3
93.7
93.3
92.6
70.0
93.8
93.5
93.4
4.6
93.7
93.4
93.4
92.7
93.6
92.8
87.8
10.1
93.7
93.3
92.1
53.6
93.6
92.8
87.8
4.6
WoLA+NNM 93.7
93.4
93.1
91.0
93.7
93.3
92.9
88.2
93.8
93.6
93.4
13.7
93.7
93.5
93.4
92.8
93.6
93.1
92.2
10.1
93.7
93.4
93.0
59.1
93.6
93.1
92.2
10.1
GM - 93.8
93.6
93.0
87.1
93.5
90.3
86.1
36.4
93.8
93.5
93.6
1.4
93.8
93.2
92.3
91.5
93.3
92.2
90.5
81.5
93.6
92.5
91.1
59.6
93.3
90.3
86.1
1.4
BKT 93.8
93.6
93.0
87.1
93.6
93.4
86.1
36.4
93.6
93.2
93.6
1.4
93.6
93.4
92.3
91.5
93.6
92.9
90.5
81.5
93.6
93.3
91.1
59.6
93.6
92.9
86.1
1.4
FoundFL 92.9
92.9
92.6
92.8
91.0
72.7
12.7
11.0
92.3
37.8
42.8
22.8
92.6
92.7
92.7
92.9
88.8
73.7
10.7
11.7
91.5
74.0
50.3
46.3
88.8
37.8
10.7
11.0
NNM 93.8
93.6
93.2
93.3
93.7
93.4
92.9
79.6
93.8
93.5
93.7
93.2
93.7
93.4
92.6
91.0
93.6
92.9
91.4
59.5
93.7
93.4
92.7
83.3
93.6
92.9
91.4
59.5
WoLA 93.8
93.6
93.5
92.2
93.7
93.3
92.9
87.6
93.8
93.6
93.6
93.5
93.8
93.6
93.5
93.6
93.5
92.7
90.9
83.6
93.7
93.4
92.9
90.1
93.5
92.7
90.9
83.6
WoLA+NNM 93.8
93.6
93.4
93.4
93.7
93.4
93.1
91.0
93.8
93.7
93.7
93.6
93.8
93.5
93.5
93.5
93.7
93.1
92.6
91.2
93.7
93.5
93.2
92.5
93.7
93.1
92.6
91.0
WoLA 93.8
93.5
93.1
91.7
93.6
93.2
92.5
87.9
93.7
93.6
93.2
92.7
93.7
93.5
93.2
92.8
93.5
92.7
90.9
83.9
93.6
93.3
92.6
89.8
93.5
92.7
90.9
83.9
WoLA+NNM 93.8
93.5
93.1
92.9
93.7
93.3
92.8
89.9
93.8
93.5
93.3
92.9
93.7
93.5
93.4
92.7
93.6
93.0
92.2
90.6
93.7
93.4
93.0
91.8
93.6
93.0
92.2
89.9
Table 6: Accuracy (s.d.) % on MNIST with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 93.9
93.1
90.6
55.4
93.2
91.7
86.7
62.4
93.2
91.0
68.4
21.5
93.2
92.6
89.8
79.9
92.6
90.9
86.1
35.1
93.2
91.8
84.3
50.9
92.6
90.9
68.4
21.5
BKT 93.9
93.6
90.6
55.4
93.8
93.3
86.7
62.4
93.2
91.5
68.4
21.5
93.8
92.8
89.8
79.9
93.6
92.6
86.1
35.1
93.7
92.8
84.3
50.9
93.2
91.5
68.4
21.5
FoundFL 86.7
91.4
91.4
89.5
84.2
64.8
23.9
10.0
80.7
78.5
78.9
50.7
82.7
87.8
92.5
92.4
81.4
30.8
11.3
13.6
83.1
70.6
59.6
51.2
80.7
30.8
11.3
10.0
NNM 94.0
93.6
92.6
93.1
93.9
93.5
92.2
58.3
93.9
93.7
93.7
92.4
93.9
93.0
87.0
73.7
93.8
90.8
87.2
32.1
93.9
92.9
90.5
69.9
93.8
90.8
87.0
32.1
WoLA 94.0
93.6
93.1
89.8
93.7
93.1
91.8
88.0
93.8
93.7
93.0
88.8
94.0
93.7
93.5
93.3
93.7
92.9
92.0
87.3
93.8
93.4
92.7
89.4
93.7
92.9
91.8
87.3
WoLA+NNM 94.1
93.7
93.2
93.9
93.9
93.4
92.9
84.7
94.1
93.9
93.9
93.9
94.1
93.7
93.5
93.1
93.9
93.2
92.4
91.3
94.0
93.6
93.2
91.4
93.9
93.2
92.4
84.7
WoLA 93.9
93.5
92.8
85.9
93.6
93.0
91.8
89.2
93.8
93.4
92.1
86.7
93.8
93.6
92.8
85.9
93.6
92.9
92.2
88.7
93.7
93.3
92.3
87.3
93.6
92.9
91.8
85.9
WoLA+NNM 93.9
93.6
93.2
91.7
93.8
93.5
92.7
72.3
94.0
93.8
93.4
92.4
93.9
93.5
92.9
86.6
93.7
93.1
91.8
65.1
93.9
93.5
92.8
81.6
93.7
93.1
91.8
65.1
CwTM - 94.1
93.5
91.4
56.4
93.9
93.1
88.1
63.5
93.6
91.9
74.9
20.5
93.7
92.3
90.4
78.3
93.8
91.5
86.5
39.1
93.8
92.5
86.3
51.6
93.6
91.5
74.9
20.5
BKT 93.9
93.6
91.4
56.4
93.8
93.3
88.1
63.5
93.2
91.5
74.9
20.5
93.8
92.9
90.4
78.3
93.6
92.6
86.5
39.1
93.7
92.8
86.3
51.6
93.2
91.5
74.9
20.5
FoundFL 91.9
91.1
90.0
89.7
91.6
87.8
80.0
20.6
79.6
48.5
55.4
21.5
90.1
90.7
91.3
91.5
88.1
19.3
10.6
14.3
88.3
67.5
65.5
47.5
79.6
19.3
10.6
14.3
NNM 94.0
93.7
92.5
93.1
94.0
93.5
92.4
59.0
94.0
93.7
93.8
92.3
93.9
92.5
90.9
76.5
93.9
92.4
87.0
25.0
93.9
93.2
91.3
69.2
93.9
92.4
87.0
25.0
WoLA 94.1
93.7
93.1
89.6
93.9
93.4
92.4
88.2
93.8
93.6
92.8
89.2
94.0
93.6
93.4
93.1
93.8
93.0
91.6
86.6
93.9
93.5
92.7
89.3
93.8
93.0
91.6
86.6
WoLA+NNM 94.1
93.8
93.3
93.9
93.9
93.4
92.9
85.5
94.1
93.9
93.9
93.9
94.0
93.6
93.5
93.2
93.9
93.2
92.2
91.8
94.0
93.6
93.2
91.7
93.9
93.2
92.2
85.5
WoLA 94.0
93.6
92.9
85.5
93.8
93.3
92.2
89.2
93.8
93.3
92.3
86.2
93.9
93.6
92.7
84.6
93.7
92.9
91.7
88.5
93.8
93.4
92.3
86.8
93.7
92.9
91.7
84.6
WoLA+NNM 93.9
93.7
93.2
91.7
93.8
93.5
92.8
71.9
94.0
93.8
93.4
92.4
93.9
93.5
92.9
83.3
93.8
93.1
92.0
65.0
93.9
93.5
92.9
80.8
93.8
93.1
92.0
65.0
MKrum - 93.9
93.6
92.4
32.4
93.8
93.5
92.1
56.4
93.8
93.5
49.4
11.4
93.8
92.6
90.7
73.7
93.7
90.6
30.6
6.5
93.8
92.7
71.0
36.1
93.7
90.6
30.6
6.5
BKT 93.7
93.3
92.4
32.4
93.4
92.8
92.1
56.4
92.7
87.4
49.4
11.4
93.4
92.1
90.7
73.7
93.1
90.1
30.6
6.5
93.3
91.1
71.0
36.1
92.7
87.4
30.6
6.5
FoundFL 91.6
91.2
88.7
77.3
91.7
90.0
55.2
29.9
85.8
72.8
53.8
29.3
91.8
90.9
90.2
86.8
82.3
12.8
11.0
14.2
88.6
71.6
59.8
47.5
82.3
12.8
11.0
14.2
NNM 94.1
93.6
92.4
39.9
94.0
93.5
92.2
70.7
93.9
93.7
93.5
26.1
93.9
92.7
74.6
59.9
93.9
92.0
81.4
6.5
93.9
93.1
86.8
40.6
93.9
92.0
74.6
6.5
WoLA 94.1
93.8
93.3
89.0
93.9
93.4
92.7
76.2
94.1
93.9
93.7
12.5
94.0
93.7
93.4
93.1
93.8
92.9
85.6
6.5
94.0
93.5
91.7
55.5
93.8
92.9
85.6
6.5
WoLA+NNM 94.1
93.8
93.2
90.5
93.9
93.5
92.9
71.4
94.1
93.9
93.9
23.9
94.1
93.7
93.6
93.1
93.9
93.2
92.2
6.7
94.0
93.6
93.2
57.1
93.9
93.2
92.2
6.7
WoLA 93.9
93.6
93.1
83.2
93.7
93.4
92.6
49.2
94.0
93.8
93.3
8.6
93.9
93.5
92.8
86.6
93.7
92.7
84.5
12.3
93.8
93.4
91.3
48.0
93.7
92.7
84.5
8.6
WoLA+NNM 93.9
93.7
93.1
85.9
93.8
93.4
92.8
58.1
94.0
93.8
93.4
17.7
93.8
93.5
92.8
86.3
93.8
93.1
91.7
12.3
93.9
93.5
92.8
52.0
93.8
93.1
91.7
12.3
GM - 94.0
93.6
92.4
51.0
93.1
87.4
77.2
24.6
93.9
93.7
93.6
2.1
93.9
92.9
84.6
31.2
93.2
91.9
89.5
68.4
93.6
91.9
87.5
35.5
93.1
87.4
77.2
2.1
BKT 94.0
93.7
92.4
51.0
93.9
93.5
77.2
24.6
93.7
93.1
93.6
2.1
93.9
93.3
84.6
31.2
93.7
92.8
89.5
68.4
93.8
93.3
87.5
35.5
93.7
92.8
77.2
2.1
FoundFL 91.0
92.7
91.7
90.6
62.7
49.6
10.9
17.2
86.1
69.7
24.0
61.7
87.3
85.5
91.2
91.9
81.2
45.7
11.2
15.5
81.7
68.6
45.8
55.4
62.7
45.7
10.9
15.5
NNM 94.1
93.7
92.7
92.7
94.0
93.6
92.8
72.7
94.0
93.7
93.8
90.9
93.8
92.7
91.5
82.9
93.9
91.1
87.2
26.2
93.9
93.0
91.6
73.1
93.8
91.1
87.2
26.2
WoLA 94.1
93.8
93.4
91.0
93.9
93.4
92.3
77.0
93.9
93.8
93.5
93.2
94.0
93.6
93.5
93.4
93.7
92.8
90.5
83.5
93.9
93.5
92.7
87.6
93.7
92.8
90.5
77.0
WoLA+NNM 94.1
93.8
93.3
93.6
93.9
93.5
93.1
90.1
93.9
93.9
93.6
93.5
94.0
93.6
93.5
93.3
93.8
93.2
92.3
90.7
94.0
93.6
93.2
92.2
93.8
93.2
92.3
90.1
WoLA 94.0
93.6
93.0
87.0
93.8
93.3
92.0
75.7
93.8
93.5
93.4
92.2
93.9
93.5
92.7
91.5
93.7
92.7
90.6
79.8
93.8
93.3
92.3
85.3
93.7
92.7
90.6
75.7
WoLA+NNM 93.9
93.7
93.2
91.0
93.8
93.4
92.9
64.5
93.8
93.5
93.5
92.7
93.9
93.6
92.9
87.0
93.8
93.1
91.9
63.2
93.9
93.5
92.9
79.7
93.8
93.1
91.9
63.2
Table 7: Accuracy (s.d.) % on MNIST with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 93.1
91.1
65.2
15.0
90.0
85.8
70.4
16.0
82.9
68.9
40.1
8.4
89.3
76.8
55.5
12.6
80.5
60.5
50.3
16.2
87.2
76.6
56.3
13.6
80.5
60.5
40.1
8.4
BKT 93.8
93.0
65.2
15.0
93.6
92.4
70.4
16.0
92.8
87.6
40.1
8.4
93.6
90.7
55.5
12.6
93.4
86.9
50.3
16.2
93.4
90.1
56.3
13.6
92.8
86.9
40.1
8.4
FoundFL 66.5
78.7
66.3
69.5
30.1
23.7
12.0
18.3
55.0
43.5
43.8
21.4
50.4
39.5
50.8
67.4
36.1
31.5
16.5
15.5
47.6
43.4
37.9
38.4
30.1
23.7
12.0
15.5
NNM 94.0
93.7
79.4
92.8
93.9
93.4
84.7
29.9
94.1
93.8
65.8
25.4
93.7
90.4
79.2
35.7
93.7
80.0
31.0
26.9
93.9
90.3
68.0
42.1
93.7
80.0
31.0
25.4
WoLA 93.7
93.4
92.0
68.9
93.2
92.4
87.0
49.1
93.4
92.9
88.9
41.3
93.5
93.3
92.8
91.4
93.0
92.0
88.1
42.5
93.4
92.8
89.8
58.6
93.0
92.0
87.0
41.3
WoLA+NNM 93.8
93.7
92.6
91.6
93.7
93.4
91.2
22.6
93.7
94.0
94.0
93.6
93.6
93.1
92.7
91.4
93.6
92.5
75.0
46.9
93.7
93.3
89.1
69.2
93.6
92.5
75.0
22.6
WoLA 93.7
93.2
91.7
57.1
93.2
92.2
86.9
65.5
93.4
92.7
87.7
36.4
93.5
92.9
91.3
56.2
93.1
92.1
85.7
61.7
93.4
92.6
88.7
55.4
93.1
92.1
85.7
36.4
WoLA+NNM 93.8
93.6
92.5
87.1
93.7
93.3
91.3
27.7
93.7
93.9
93.0
71.5
93.6
93.1
91.8
57.5
93.5
92.4
73.4
15.8
93.7
93.3
88.4
51.9
93.5
92.4
73.4
15.8
CwTM - 93.9
92.4
75.5
14.5
93.5
90.2
62.3
16.7
93.4
80.9
35.6
8.7
93.4
90.1
60.4
29.1
93.3
69.6
42.1
17.8
93.5
84.6
55.2
17.4
93.3
69.6
35.6
8.7
BKT 93.8
93.0
75.5
14.5
93.6
92.4
62.3
16.7
92.8
87.8
35.6
8.7
93.5
91.1
60.4
29.1
93.4
87.2
42.1
17.8
93.4
90.3
55.2
17.4
92.8
87.2
35.6
8.7
FoundFL 79.0
73.4
78.9
69.4
81.0
63.3
27.9
20.0
53.5
37.6
21.6
22.7
56.9
81.6
76.5
68.9
56.5
18.0
13.9
17.4
65.4
54.8
43.8
39.7
53.5
18.0
13.9
17.4
NNM 94.1
93.6
75.1
92.8
93.9
93.5
85.2
22.4
94.2
93.9
58.5
25.3
93.8
91.1
56.5
27.7
93.7
87.7
55.5
35.8
93.9
92.0
66.1
40.8
93.7
87.7
55.5
22.4
WoLA 93.8
93.7
92.4
69.3
93.6
93.3
91.3
49.2
93.6
93.2
91.1
44.9
93.6
93.3
92.8
89.7
93.5
92.5
90.4
43.8
93.6
93.2
91.6
59.4
93.5
92.5
90.4
43.8
WoLA+NNM 93.8
93.7
92.7
92.2
93.7
93.4
91.9
26.2
93.7
93.9
94.0
93.6
93.6
93.2
92.3
91.1
93.6
92.6
83.6
50.5
93.7
93.4
90.9
70.7
93.6
92.6
83.6
26.2
WoLA 93.8
93.6
92.1
57.3
93.7
93.2
90.8
65.2
93.6
92.7
91.0
36.4
93.7
93.3
91.1
60.7
93.5
92.4
89.8
62.0
93.6
93.0
91.0
56.3
93.5
92.4
89.8
36.4
WoLA+NNM 93.8
93.7
92.6
86.3
93.7
93.3
91.4
28.9
93.7
93.9
93.0
70.4
93.6
93.3
91.2
57.5
93.6
92.7
88.8
23.9
93.7
93.4
91.4
53.4
93.6
92.7
88.8
23.9
MKrum - 94.0
93.6
77.7
5.3
93.9
93.6
85.0
31.7
94.1
73.7
55.6
31.6
93.4
91.6
79.2
35.7
93.6
86.5
15.7
6.0
93.8
87.8
62.6
22.0
93.4
73.7
15.7
5.3
BKT 93.3
92.9
77.7
5.3
93.0
91.9
85.0
31.7
91.2
78.0
55.6
31.6
92.9
86.7
79.2
35.7
92.6
81.3
15.7
6.0
92.6
86.2
62.6
22.0
91.2
78.0
15.7
5.3
FoundFL 82.3
67.4
46.1
30.2
76.3
73.3
64.5
21.8
87.0
52.5
34.0
37.1
85.1
74.1
75.4
49.8
74.5
16.3
16.5
17.4
81.0
56.7
47.3
31.3
74.5
16.3
16.5
17.4
NNM 94.1
93.6
78.9
6.9
93.9
93.3
85.1
24.0
94.1
93.8
57.4
24.3
93.9
91.5
55.1
30.3
93.7
85.7
53.7
6.0
93.9
91.6
66.0
18.3
93.7
85.7
53.7
6.0
WoLA 93.9
93.6
92.6
67.9
93.7
93.3
91.4
50.8
93.7
94.0
93.6
18.3
93.6
93.2
92.5
91.4
93.5
92.2
48.4
5.1
93.7
93.2
83.7
46.7
93.5
92.2
48.4
5.1
WoLA+NNM 93.8
93.7
92.6
73.0
93.7
93.3
91.6
34.6
93.7
93.9
94.0
24.2
93.6
93.1
92.4
91.3
93.6
92.5
85.7
5.1
93.7
93.3
91.3
45.6
93.6
92.5
85.7
5.1
WoLA 93.8
93.6
92.3
39.3
93.7
93.3
91.3
33.5
93.7
93.9
93.1
15.4
93.5
93.0
91.3
57.5
93.5
92.1
42.4
7.5
93.7
93.2
82.1
30.7
93.5
92.1
42.4
7.5
WoLA+NNM 93.8
93.6
92.4
55.7
93.7
93.3
91.0
17.7
93.7
93.9
93.0
21.0
93.7
93.2
91.6
57.7
93.6
92.6
84.1
7.6
93.7
93.3
90.4
31.9
93.6
92.6
84.1
7.6
GM - 94.0
93.4
86.9
7.3
90.1
79.0
48.2
6.2
94.0
93.5
66.1
9.6
94.0
89.6
26.4
47.8
92.8
91.4
76.8
10.0
93.0
89.4
60.9
16.2
90.1
79.0
26.4
6.2
BKT 94.0
93.8
86.9
7.3
93.8
93.4
48.2
6.2
93.6
86.2
66.1
9.6
93.8
92.2
26.4
47.8
93.6
92.6
76.8
10.0
93.8
91.6
60.9
16.2
93.6
86.2
26.4
6.2
FoundFL 64.3
78.2
78.9
59.7
52.1
24.1
24.0
23.7
35.9
43.0
34.7
31.2
40.3
56.9
60.7
81.2
69.6
31.2
16.5
17.4
52.4
46.7
42.9
42.6
35.9
24.1
16.5
17.4
NNM 94.1
93.7
83.2
88.7
94.0
93.5
86.1
38.0
94.1
93.9
58.9
29.6
93.8
91.0
63.1
18.6
93.7
89.0
58.2
22.3
93.9
92.2
69.9
39.4
93.7
89.0
58.2
18.6
WoLA 93.8
93.8
92.9
76.7
93.5
92.1
84.8
37.1
93.7
93.8
93.3
24.9
93.6
93.3
92.2
92.0
93.3
91.6
89.4
36.6
93.6
92.9
90.5
53.4
93.3
91.6
84.8
24.9
WoLA+NNM 93.7
93.7
93.0
93.1
93.7
93.5
92.8
42.3
93.7
93.8
93.5
93.3
93.6
93.3
92.9
92.1
93.6
92.7
71.5
43.8
93.7
93.4
88.7
72.9
93.6
92.7
71.5
42.3
WoLA 93.7
93.6
92.4
71.2
93.5
92.1
85.8
27.2
93.7
93.6
92.9
17.8
93.7
93.4
91.6
86.4
93.3
91.4
89.9
31.1
93.6
92.8
90.5
46.7
93.3
91.4
85.8
17.8
WoLA+NNM 93.8
93.7
92.7
86.9
93.7
93.4
92.2
31.1
93.7
93.7
93.0
68.1
93.6
93.2
92.3
89.7
93.6
92.7
74.4
12.2
93.7
93.4
88.9
57.6
93.6
92.7
74.4
12.2
Table 8: Accuracy (s.d.) % on Fashion MNIST with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 78.0
75.1
70.5
61.4
74.2
70.6
65.3
55.3
75.5
71.3
61.3
32.2
75.9
73.8
71.2
70.7
74.1
71.2
65.2
45.2
75.5
72.4
66.7
53.0
74.1
70.6
61.3
32.2
BKT 77.7
76.6
70.5
61.4
77.0
75.5
65.3
55.3
76.0
72.4
61.3
32.2
76.9
75.9
71.2
70.7
76.5
72.8
65.2
45.2
76.8
74.7
66.7
53.0
76.0
72.4
61.3
32.2
FoundFL 72.1
73.2
73.1
69.6
70.0
58.2
11.4
11.7
70.0
60.3
58.4
44.5
71.5
72.4
73.4
74.8
66.1
52.9
11.7
14.1
69.9
63.4
45.6
42.9
66.1
52.9
11.4
11.7
NNM 79.0
77.4
74.1
76.4
78.3
76.8
73.1
30.7
79.3
79.2
79.1
79.3
77.6
75.2
72.0
70.8
77.8
72.6
62.1
38.7
78.4
76.2
72.1
59.2
77.6
72.6
62.1
30.7
WoLA 78.8
78.3
77.1
73.1
77.9
76.6
74.1
71.9
78.9
78.2
77.1
74.1
78.7
77.9
78.0
77.5
78.1
76.6
74.3
71.9
78.5
77.5
76.1
73.7
77.9
76.6
74.1
71.9
WoLA+NNM 79.0
78.6
78.0
78.7
78.3
77.4
75.8
71.5
78.9
79.2
79.2
79.1
78.7
77.9
77.7
77.1
78.0
76.6
75.1
74.8
78.6
77.9
77.2
76.2
78.0
76.6
75.1
71.5
WoLA 78.9
77.8
76.0
68.1
77.9
76.0
72.4
67.5
78.9
77.9
75.5
68.2
78.3
77.9
76.8
72.4
77.9
76.1
72.8
67.1
78.4
77.1
74.7
68.7
77.9
76.0
72.4
67.1
WoLA+NNM 78.7
78.6
76.5
74.9
78.6
77.4
74.6
63.7
78.9
78.6
77.9
75.6
78.5
77.4
77.0
73.1
77.7
75.9
73.1
67.8
78.5
77.6
75.8
71.0
77.7
75.9
73.1
63.7
CwTM - 78.8
77.0
72.4
58.8
78.0
75.0
66.4
55.2
77.6
72.1
63.7
29.8
77.4
75.4
72.7
70.9
77.2
70.9
67.3
46.2
77.8
74.1
68.5
52.2
77.2
70.9
63.7
29.8
BKT 77.6
76.6
72.4
58.8
77.0
75.5
66.4
55.2
75.8
72.4
63.7
29.8
76.6
75.9
72.7
70.9
76.4
72.9
67.3
46.2
76.7
74.7
68.5
52.2
75.8
72.4
63.7
29.8
FoundFL 75.0
73.7
73.2
69.7
74.0
72.2
66.6
18.8
68.6
59.9
54.1
36.3
75.0
74.2
74.3
74.9
68.9
21.5
13.2
13.2
72.3
60.3
56.3
42.6
68.6
21.5
13.2
13.2
NNM 79.1
77.6
73.9
76.3
78.6
77.1
73.1
30.7
79.3
79.4
79.1
79.3
77.7
75.6
72.2
71.8
78.1
73.7
63.8
38.7
78.6
76.7
72.4
59.4
77.7
73.7
63.8
30.7
WoLA 79.2
78.7
77.4
72.5
78.8
77.4
74.4
71.9
79.0
78.2
75.9
74.4
79.0
78.2
77.6
77.7
78.2
76.3
73.9
71.7
78.8
77.7
75.9
73.6
78.2
76.3
73.9
71.7
WoLA+NNM 79.1
78.8
78.1
78.7
78.5
77.3
75.7
71.7
78.8
79.3
79.2
79.0
78.5
77.9
77.9
77.5
78.1
76.7
74.7
74.3
78.6
78.0
77.1
76.3
78.1
76.7
74.7
71.7
WoLA 79.0
77.9
76.2
68.5
78.7
77.0
73.2
67.3
78.9
77.8
75.3
69.0
78.8
77.8
76.6
71.4
78.2
75.9
72.1
67.3
78.7
77.3
74.7
68.7
78.2
75.9
72.1
67.3
WoLA+NNM 79.1
78.6
76.4
74.7
78.9
77.5
74.6
64.8
78.9
78.6
77.9
75.6
78.2
77.5
76.9
72.6
77.8
75.8
73.5
68.0
78.6
77.6
75.9
71.1
77.8
75.8
73.5
64.8
MKrum - 78.5
77.0
73.5
52.8
78.3
76.4
72.6
21.4
79.4
78.8
37.8
1.7
77.1
75.3
73.1
70.8
77.7
72.0
21.4
8.5
78.2
75.9
55.7
31.0
77.1
72.0
21.4
1.7
BKT 77.4
75.5
73.5
52.8
76.2
73.9
72.6
21.4
75.7
73.0
37.8
1.7
76.9
73.7
73.1
70.8
75.2
67.5
21.4
8.5
76.3
72.7
55.7
31.0
75.2
67.5
21.4
1.7
FoundFL 74.4
74.2
71.6
67.3
74.4
73.4
66.8
25.6
68.7
61.0
75.6
10.6
74.3
74.7
72.5
73.6
64.6
22.4
11.0
41.9
71.3
61.1
59.5
43.8
64.6
22.4
11.0
10.6
NNM 79.0
77.7
73.7
59.0
78.5
76.9
72.9
44.2
79.4
79.3
78.7
4.8
77.3
75.7
72.1
69.8
77.9
73.5
61.8
8.5
78.4
76.6
71.8
37.3
77.3
73.5
61.8
4.8
WoLA 78.8
78.5
77.8
72.5
78.4
77.6
74.6
23.7
79.0
79.2
79.1
1.8
78.5
77.9
77.6
77.1
78.1
75.6
68.5
9.1
78.6
77.8
75.5
36.8
78.1
75.6
68.5
1.8
WoLA+NNM 79.0
79.0
77.9
74.2
78.3
77.7
75.5
65.0
79.0
79.3
79.2
5.1
79.0
77.8
77.6
77.5
78.1
76.7
74.9
8.5
78.7
78.1
77.0
46.0
78.1
76.7
74.9
5.1
WoLA 78.5
78.6
76.4
68.5
77.9
77.2
73.3
8.5
78.9
78.9
78.4
3.5
78.5
77.6
76.6
73.1
77.6
75.4
66.2
9.2
78.3
77.5
74.2
32.5
77.6
75.4
66.2
3.5
WoLA+NNM 78.8
78.3
76.4
69.8
78.3
77.4
74.9
57.0
78.9
78.6
77.9
9.8
78.6
77.6
76.5
72.6
77.9
75.7
73.1
9.2
78.5
77.5
75.8
43.7
77.9
75.7
73.1
9.2
GM - 78.6
77.2
74.3
64.5
73.9
64.9
59.0
34.1
79.0
78.9
78.1
9.4
77.8
74.7
72.0
72.7
76.0
71.4
56.1
45.3
77.1
73.4
67.9
45.2
73.9
64.9
56.1
9.4
BKT 78.0
77.5
74.3
64.5
77.5
75.9
59.0
34.1
77.2
77.2
78.1
9.4
77.0
76.0
72.0
72.7
76.5
73.2
56.1
45.3
77.3
75.9
67.9
45.2
76.5
73.2
56.1
9.4
FoundFL 75.5
76.8
74.3
71.4
72.1
49.1
13.8
22.9
72.3
71.7
63.8
65.1
74.9
74.6
75.2
75.0
62.8
46.9
11.1
13.0
71.5
63.8
47.6
49.5
62.8
46.9
11.1
13.0
NNM 79.0
77.9
74.2
76.0
78.4
77.1
73.4
54.0
79.4
79.3
78.9
79.2
77.4
75.1
72.7
72.1
77.7
73.5
63.9
35.2
78.4
76.6
72.6
63.3
77.4
73.5
63.9
35.2
WoLA 78.9
78.9
77.7
74.4
78.1
76.3
73.8
66.5
79.0
79.1
78.4
78.3
79.0
77.7
78.0
78.1
77.6
74.8
70.8
60.5
78.5
77.4
75.7
71.5
77.6
74.8
70.8
60.5
WoLA+NNM 78.9
78.9
78.0
78.7
78.5
77.8
76.1
71.3
79.1
78.9
78.7
78.8
78.9
77.9
78.0
77.8
77.9
76.4
74.5
70.8
78.7
78.0
77.1
75.5
77.9
76.4
74.5
70.8
WoLA 79.3
78.5
76.7
69.4
78.0
76.0
71.8
60.1
79.1
78.6
77.5
73.4
78.7
77.6
76.5
74.2
77.7
74.5
69.7
54.5
78.6
77.0
74.4
66.3
77.7
74.5
69.7
54.5
WoLA+NNM 78.9
78.5
77.0
74.3
78.3
77.6
74.9
65.9
79.1
78.5
78.0
75.4
78.4
77.9
76.6
73.5
78.0
76.1
73.3
65.4
78.6
77.7
75.9
70.9
78.0
76.1
73.3
65.4
Table 9: Accuracy (s.d.) % on Fashion MNIST with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 77.5
73.8
63.0
42.8
72.4
67.5
55.5
29.1
73.0
60.2
41.3
15.0
73.1
70.9
64.8
63.8
70.2
63.4
53.9
25.8
73.2
67.2
55.7
35.3
70.2
60.2
41.3
15.0
BKT 78.0
76.1
63.0
42.8
77.2
74.3
55.5
29.1
76.4
72.0
41.3
15.0
77.0
73.9
64.8
63.8
76.2
70.7
53.9
25.8
77.0
73.4
55.7
35.3
76.2
70.7
41.3
15.0
FoundFL 64.8
69.8
66.8
66.6
56.3
44.8
12.2
26.5
62.7
54.3
49.8
27.4
41.3
63.9
69.1
73.0
44.6
21.2
12.0
11.2
53.9
50.8
42.0
40.9
41.3
21.2
12.0
11.2
NNM 80.2
78.0
72.4
75.4
79.4
77.4
73.0
21.7
80.1
80.2
79.6
75.0
78.1
73.3
68.8
62.9
78.5
72.8
57.4
31.6
79.3
76.4
70.2
53.3
78.1
72.8
57.4
21.7
WoLA 80.0
79.0
76.9
71.6
79.0
76.8
74.2
70.8
79.9
78.5
77.5
72.3
79.7
78.8
78.4
78.0
78.8
76.7
74.6
69.8
79.5
78.0
76.3
72.5
78.8
76.7
74.2
69.8
WoLA+NNM 80.2
79.5
78.2
78.6
79.5
78.3
76.2
60.1
79.9
79.6
79.9
80.2
79.5
78.5
78.2
77.9
79.0
76.8
75.4
72.5
79.6
78.6
77.6
73.8
79.0
76.8
75.4
60.1
WoLA 79.8
77.3
73.5
60.7
78.8
75.8
71.2
65.4
79.3
77.7
74.0
62.9
79.5
77.5
76.1
69.7
78.7
76.0
71.8
62.5
79.2
76.9
73.3
64.2
78.7
75.8
71.2
60.7
WoLA+NNM 79.9
77.9
76.0
68.7
79.5
77.1
73.4
49.1
80.2
78.6
78.0
75.0
79.3
77.5
75.1
69.0
79.1
75.4
71.1
46.1
79.6
77.3
74.7
61.5
79.1
75.4
71.1
46.1
CwTM - 79.4
76.4
68.0
43.4
77.3
72.6
61.6
27.0
76.6
69.0
48.0
14.1
77.3
73.1
68.5
61.6
76.6
64.5
57.3
21.5
77.5
71.1
60.7
33.5
76.6
64.5
48.0
14.1
BKT 78.1
76.1
68.0
43.4
77.2
74.3
61.6
27.0
76.4
72.4
48.0
14.1
77.5
74.4
68.5
61.6
76.2
70.6
57.3
21.5
77.1
73.6
60.7
33.5
76.2
70.6
48.0
14.1
FoundFL 70.1
68.7
62.0
58.4
68.9
63.4
59.0
36.7
69.2
51.0
45.2
26.7
70.3
65.8
68.1
71.6
65.3
16.7
26.4
15.1
68.7
53.2
52.1
41.7
65.3
16.7
26.4
15.1
NNM 80.5
78.4
72.7
75.2
80.0
77.9
73.2
22.5
80.5
80.3
79.6
71.9
77.5
74.2
70.9
62.6
79.2
74.1
60.6
33.5
79.5
77.0
71.4
53.2
77.5
74.1
60.6
22.5
WoLA 80.0
79.6
77.7
71.6
79.7
78.0
74.7
70.9
79.5
78.3
76.7
72.3
79.8
78.8
78.7
78.3
79.1
76.6
74.3
69.2
79.6
78.3
76.4
72.5
79.1
76.6
74.3
69.2
WoLA+NNM 80.0
79.5
78.1
78.8
79.6
78.3
76.2
62.0
79.9
79.6
79.9
80.2
79.3
78.6
77.8
78.0
79.2
76.9
75.2
72.5
79.6
78.6
77.4
74.3
79.2
76.9
75.2
62.0
WoLA 79.9
78.1
74.1
59.9
79.4
77.0
71.5
65.6
79.2
77.4
70.7
64.4
79.1
77.5
76.1
69.1
79.1
75.3
71.0
62.1
79.3
77.1
72.7
64.2
79.1
75.3
70.7
59.9
WoLA+NNM 79.7
78.4
76.1
69.3
79.5
77.2
73.5
50.0
80.2
78.6
77.9
75.0
79.5
77.0
75.7
68.7
79.1
75.5
71.4
45.4
79.6
77.3
74.9
61.7
79.1
75.5
71.4
45.4
MKrum - 79.9
77.6
71.3
19.0
78.9
77.8
72.8
32.6
80.3
78.0
40.6
11.0
77.9
74.0
70.3
62.9
78.4
73.0
26.3
6.8
79.1
76.1
56.3
26.5
77.9
73.0
26.3
6.8
BKT 77.8
75.5
71.3
19.0
76.4
74.1
72.8
32.6
75.9
68.9
40.6
11.0
77.1
73.4
70.3
62.9
75.1
65.9
26.3
6.8
76.5
71.6
56.3
26.5
75.1
65.9
26.3
6.8
FoundFL 71.4
66.8
64.4
52.0
69.6
65.7
61.2
41.8
75.0
71.1
19.8
19.4
68.2
67.0
69.9
69.1
58.5
25.8
24.1
26.7
68.5
59.3
47.9
41.8
58.5
25.8
19.8
19.4
NNM 80.5
78.4
72.6
29.3
79.9
77.8
73.1
29.8
80.5
80.2
79.0
19.2
77.9
73.4
70.5
62.3
78.8
73.1
42.9
7.1
79.5
76.6
67.6
29.5
77.9
73.1
42.9
7.1
WoLA 80.1
79.2
77.8
72.1
79.3
77.9
75.2
24.1
79.9
79.6
79.9
8.1
79.5
78.4
78.2
77.9
79.0
75.9
67.4
5.4
79.6
78.2
75.7
37.5
79.0
75.9
67.4
5.4
WoLA+NNM 79.9
79.5
78.2
72.7
79.7
78.3
76.0
59.4
79.9
79.6
79.9
15.9
79.5
78.5
78.3
77.9
79.2
76.9
74.9
5.4
79.6
78.6
77.5
46.3
79.2
76.9
74.9
5.4
WoLA 79.8
77.6
75.6
61.0
79.2
76.9
71.9
20.5
80.1
78.6
77.8
5.8
79.2
77.2
75.7
69.0
78.9
74.8
59.7
8.4
79.4
77.0
72.1
32.9
78.9
74.8
59.7
5.8
WoLA+NNM 79.7
78.5
75.9
61.2
79.5
77.1
73.6
41.7
80.1
78.6
78.0
11.1
79.6
77.0
75.7
69.6
79.1
75.8
70.9
8.4
79.6
77.4
74.8
38.4
79.1
75.8
70.9
8.4
GM - 80.1
77.1
72.0
31.5
70.2
60.7
47.5
21.8
80.2
79.7
72.5
4.6
78.4
74.3
67.2
67.5
75.1
69.2
61.6
44.7
76.8
72.2
64.2
34.0
70.2
60.7
47.5
4.6
BKT 79.0
77.5
72.0
31.5
78.3
76.2
47.5
21.8
78.4
75.9
72.5
4.6
78.0
75.9
67.2
67.5
77.4
73.5
61.6
44.7
78.2
75.8
64.2
34.0
77.4
73.5
47.5
4.6
FoundFL 73.6
72.7
70.3
63.3
65.6
43.2
24.4
20.2
70.1
67.8
52.7
39.3
49.5
66.9
71.6
73.5
57.2
30.6
24.7
14.1
63.2
56.2
48.7
42.1
49.5
30.6
24.4
14.1
NNM 80.2
78.4
72.5
75.7
79.7
77.8
74.1
51.6
80.5
80.5
79.7
77.9
78.0
74.9
68.4
67.7
78.8
73.6
57.2
23.7
79.4
77.0
70.4
59.3
78.0
73.6
57.2
23.7
WoLA 79.9
79.4
77.9
73.0
78.6
76.4
70.5
61.1
80.1
79.6
79.7
78.9
79.6
78.7
77.7
78.6
78.5
75.8
71.5
41.5
79.4
77.9
75.5
66.6
78.5
75.8
70.5
41.5
WoLA+NNM 79.9
79.7
78.3
78.3
79.7
78.4
76.3
70.7
80.2
79.7
79.6
80.0
79.6
78.7
78.2
78.8
79.1
76.9
74.6
70.4
79.7
78.7
77.4
75.6
79.1
76.9
74.6
70.4
WoLA 79.8
78.0
75.5
62.5
78.6
74.7
67.3
51.6
79.9
78.7
77.8
56.7
79.6
77.1
75.5
70.3
78.3
74.6
68.8
52.6
79.3
76.6
73.0
58.7
78.3
74.6
67.3
51.6
WoLA+NNM 79.8
78.0
76.4
68.9
79.5
77.0
73.5
52.3
79.9
78.6
77.7
74.3
79.4
77.2
76.5
69.5
79.2
76.1
70.9
41.9
79.6
77.4
75.0
61.3
79.2
76.1
70.9
41.9
Table 10: Accuracy (s.d.) % on Fashion MNIST with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 72.6
65.3
39.6
19.5
65.1
54.5
39.4
16.0
58.2
41.6
23.6
6.0
65.3
47.2
50.0
33.1
50.3
38.2
29.0
16.3
62.3
49.4
36.3
18.2
50.3
38.2
23.6
6.0
BKT 76.1
69.8
39.6
19.5
75.0
68.8
39.4
16.0
74.6
63.7
23.6
6.0
76.2
68.3
50.0
33.1
73.4
55.2
29.0
16.3
75.1
65.1
36.3
18.2
73.4
55.2
23.6
6.0
FoundFL 30.6
47.7
41.2
39.6
43.8
23.7
11.6
13.0
34.5
42.6
27.2
15.5
42.9
40.5
43.9
37.2
24.1
22.5
16.5
9.7
35.2
35.4
28.1
23.0
24.1
22.5
11.6
9.7
NNM 80.4
76.0
52.2
68.5
78.9
76.0
65.6
27.1
79.9
77.4
58.4
15.8
77.3
69.4
58.9
35.1
78.2
69.1
29.4
16.6
79.0
73.6
52.9
32.7
77.3
69.1
29.4
15.8
WoLA 78.7
76.7
72.6
49.3
77.2
75.1
65.8
40.1
77.9
76.3
68.7
32.1
78.2
76.2
75.9
71.9
76.7
74.5
64.4
32.7
77.7
75.8
69.5
45.2
76.7
74.5
64.4
32.1
WoLA+NNM 78.9
78.2
76.2
69.4
78.4
77.5
75.0
19.6
79.0
79.7
80.5
79.3
78.5
76.3
77.0
71.4
77.6
75.0
60.2
25.6
78.5
77.4
73.8
53.1
77.6
75.0
60.2
19.6
WoLA 78.6
74.4
68.1
40.9
76.8
71.7
61.1
47.0
77.8
73.3
65.6
22.5
77.9
73.7
71.0
48.4
76.1
71.7
58.7
39.9
77.5
72.9
64.9
39.8
76.1
71.7
58.7
22.5
WoLA+NNM 78.6
76.3
71.1
53.7
78.3
75.0
69.0
17.9
78.7
77.7
76.8
70.7
77.7
73.8
70.0
63.3
77.8
71.9
55.7
12.2
78.2
74.9
68.5
43.6
77.7
71.9
55.7
12.2
CwTM - 77.0
70.0
43.3
19.1
74.6
65.2
41.1
16.0
74.2
54.8
23.0
5.9
73.1
62.7
54.9
26.0
72.1
38.1
24.3
17.2
74.2
58.2
37.3
16.8
72.1
38.1
23.0
5.9
BKT 76.1
69.6
43.3
19.1
75.0
68.8
41.1
16.0
74.6
63.7
23.0
5.9
75.1
68.6
54.9
26.0
73.4
55.3
24.3
17.2
74.9
65.2
37.3
16.8
73.4
55.3
23.0
5.9
FoundFL 59.3
52.0
65.4
49.3
65.8
41.5
31.6
23.4
57.3
44.7
35.9
14.7
65.5
29.6
55.0
62.3
28.1
14.1
18.8
17.4
55.2
36.4
41.3
33.4
28.1
14.1
18.8
14.7
NNM 80.9
76.5
51.9
69.6
80.1
77.3
66.6
26.7
80.6
76.0
61.2
15.6
77.6
71.1
56.5
35.4
79.7
72.8
40.9
16.3
79.8
74.7
55.4
32.7
77.6
71.1
40.9
15.6
WoLA 79.1
78.5
74.7
49.8
78.4
77.1
72.5
40.3
78.2
77.1
71.9
32.6
78.5
77.0
76.1
72.7
77.8
74.8
69.9
32.9
78.4
76.9
73.0
45.7
77.8
74.8
69.9
32.6
WoLA+NNM 79.1
78.4
76.1
66.8
78.2
77.4
75.5
21.0
79.0
79.7
80.4
79.2
78.4
76.7
76.7
70.6
77.9
74.9
65.3
23.6
78.5
77.4
74.8
52.2
77.9
74.9
65.3
21.0
WoLA 78.8
76.0
70.4
35.7
78.0
75.0
67.3
47.0
77.9
74.4
69.5
19.1
77.9
74.2
71.5
62.1
77.7
71.3
62.3
40.1
78.1
74.2
68.2
40.8
77.7
71.3
62.3
19.1
WoLA+NNM 78.8
76.2
72.0
50.8
78.2
75.2
69.8
19.3
78.8
77.7
76.8
70.7
77.9
73.9
71.3
48.8
78.0
71.6
60.7
15.5
78.4
74.9
70.1
41.0
77.9
71.6
60.7
15.5
MKrum - 79.4
75.3
47.5
5.1
79.5
75.9
62.6
22.0
80.1
60.8
45.2
18.0
77.1
64.5
47.7
35.1
78.4
66.0
13.8
6.1
78.9
68.5
43.4
17.3
77.1
60.8
13.8
5.1
BKT 76.9
72.2
47.5
5.1
75.5
71.5
62.6
22.0
74.1
60.4
45.2
18.0
75.6
68.0
47.7
35.1
74.0
56.2
13.8
6.1
75.2
65.7
43.4
17.3
74.0
56.2
13.8
5.1
FoundFL 70.5
57.5
53.6
28.5
68.9
59.4
48.8
26.4
36.0
45.1
29.4
17.8
70.1
61.9
52.0
54.4
54.3
41.3
16.4
17.4
60.0
53.0
40.0
28.9
36.0
41.3
16.4
17.4
NNM 80.6
75.9
55.6
6.2
80.2
77.1
66.6
23.7
80.4
75.7
50.1
16.7
76.9
66.6
63.2
42.1
79.5
69.7
33.6
6.1
79.5
73.0
53.8
19.0
76.9
66.6
33.6
6.1
WoLA 78.8
78.0
75.9
51.5
77.8
76.8
74.7
28.6
79.0
79.7
80.4
4.8
77.9
76.3
74.3
71.4
77.3
73.8
38.6
5.2
78.2
76.9
68.8
32.3
77.3
73.8
38.6
4.8
WoLA+NNM 79.1
78.3
75.8
52.4
78.3
77.5
75.0
26.9
79.0
79.7
80.5
10.3
78.4
76.3
76.0
71.9
78.0
74.5
64.1
5.1
78.5
77.3
74.3
33.3
78.0
74.5
64.1
5.1
WoLA 78.4
75.8
71.7
40.0
77.8
74.6
69.5
23.9
78.8
77.8
76.7
3.0
77.6
73.7
69.8
63.3
77.4
70.7
15.8
6.4
78.0
74.5
60.7
27.3
77.4
70.7
15.8
3.0
WoLA+NNM 78.8
76.2
71.0
40.7
78.2
74.9
69.2
15.9
78.8
77.7
76.8
7.6
77.9
74.1
70.0
61.7
78.0
71.8
58.5
6.4
78.3
74.9
69.1
26.5
77.9
71.8
58.5
6.4
GM - 79.4
75.2
58.5
6.5
67.2
43.6
29.5
12.6
80.4
76.3
50.7
1.4
77.2
68.2
48.1
44.8
71.1
62.3
52.7
8.1
75.1
65.1
47.9
14.7
67.2
43.6
29.5
1.4
BKT 78.6
76.3
58.5
6.5
78.0
75.9
29.5
12.6
77.0
70.1
50.7
1.4
77.1
74.5
48.1
44.8
77.3
71.3
52.7
8.1
77.6
73.6
47.9
14.7
77.0
70.1
29.5
1.4
FoundFL 42.5
62.5
61.6
52.3
42.0
35.9
26.7
41.2
37.3
48.5
48.2
29.1
42.7
37.1
62.8
56.4
40.1
25.0
16.4
17.4
40.9
41.8
43.1
39.3
37.3
25.0
16.4
17.4
NNM 80.5
76.6
52.9
68.0
79.8
77.5
65.9
29.8
80.6
78.5
51.1
19.1
77.1
70.2
60.4
51.1
79.3
71.0
42.4
19.3
79.5
74.8
54.5
37.5
77.1
70.2
42.4
19.1
WoLA 78.9
78.1
76.0
54.6
76.0
70.7
62.6
36.6
79.1
79.4
78.7
35.2
78.5
76.8
75.9
74.5
76.1
72.5
65.6
29.3
77.7
75.5
71.8
46.0
76.0
70.7
62.6
29.3
WoLA+NNM 79.0
78.5
76.5
69.7
78.2
77.5
75.4
29.8
79.0
79.5
79.7
78.5
78.4
76.9
75.8
73.1
78.1
75.2
56.9
31.2
78.6
77.5
72.8
56.4
78.1
75.2
56.9
29.8
WoLA 78.7
76.0
71.2
47.6
75.7
66.8
54.0
28.1
78.8
77.5
74.8
27.7
77.8
75.0
70.4
58.5
75.8
70.0
58.2
27.4
77.3
73.1
65.7
37.8
75.7
66.8
54.0
27.4
WoLA+NNM 78.8
76.4
72.1
54.6
78.2
75.4
70.1
24.8
78.9
77.7
76.6
71.1
77.7
74.3
71.7
61.0
78.1
71.8
56.9
11.2
78.3
75.1
69.5
44.5
77.7
71.8
56.9
11.2
Table 11: Accuracy (s.d.) % on Purchase100 with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 72.2
69.3
64.2
40.2
67.2
60.7
50.2
21.7
70.3
68.0
61.2
29.9
68.5
61.6
50.4
34.9
66.8
60.6
49.8
16.8
69.0
64.0
55.1
28.7
66.8
60.6
49.8
16.8
BKT 72.4
71.7
64.2
40.2
70.5
67.0
50.2
21.7
69.4
66.4
61.2
29.9
70.5
65.0
50.4
34.9
69.1
62.9
49.8
16.8
70.4
66.6
55.1
28.7
69.1
62.9
49.8
16.8
FoundFL 64.2
61.3
55.8
49.0
59.7
30.1
2.7
8.8
52.0
24.9
0.2
0.3
61.2
60.4
58.3
62.7
43.7
13.8
0.9
0.6
56.2
38.1
23.6
24.3
43.7
13.8
0.2
0.3
NNM 72.9
72.3
67.4
72.7
71.1
67.2
58.3
3.8
73.0
72.9
72.7
70.9
70.7
63.3
55.3
34.9
70.4
63.2
52.7
44.9
71.6
67.7
61.3
45.5
70.4
63.2
52.7
3.8
WoLA 71.5
71.2
69.9
66.5
69.9
67.8
64.9
58.5
71.2
70.7
69.4
64.8
71.0
69.4
67.9
63.7
69.8
67.3
64.2
55.8
70.7
69.3
67.3
61.8
69.8
67.3
64.2
55.8
WoLA+NNM 71.2
71.1
70.1
70.9
70.6
69.3
67.3
62.4
71.3
71.0
71.0
71.0
70.9
69.6
67.4
63.7
70.4
68.4
67.3
66.7
70.9
69.9
68.6
66.9
70.4
68.4
67.3
62.4
WoLA 71.3
70.7
68.5
61.3
69.8
67.3
63.3
53.8
71.1
70.3
68.3
55.5
70.7
69.3
65.9
61.8
69.6
66.9
62.8
53.7
70.5
68.9
65.8
57.2
69.6
66.9
62.8
53.7
WoLA+NNM 71.1
70.7
69.1
68.2
70.5
68.8
65.7
47.0
71.2
70.7
69.9
68.8
70.7
68.8
66.1
61.8
70.2
67.8
65.9
60.5
70.7
69.4
67.3
61.3
70.2
67.8
65.7
47.0
CwTM - 72.4
72.0
65.9
40.2
70.4
65.7
54.3
21.7
71.1
68.4
61.8
29.9
70.1
63.8
52.0
34.9
69.1
61.3
49.6
16.5
70.7
66.2
56.7
28.6
69.1
61.3
49.6
16.5
BKT 72.4
71.7
65.9
40.2
70.5
67.0
54.3
21.7
69.4
66.4
61.8
29.9
70.5
65.0
52.0
34.9
69.1
62.9
49.6
16.5
70.4
66.6
56.7
28.6
69.1
62.9
49.6
16.5
FoundFL 62.3
57.3
53.0
49.5
61.2
55.0
54.5
7.8
45.1
27.4
10.5
0.5
58.9
55.1
57.0
61.1
37.3
1.7
0.9
0.6
53.0
39.3
35.2
23.9
37.3
1.7
0.9
0.5
NNM 72.9
72.6
68.7
72.7
71.1
67.3
58.3
3.8
73.0
72.9
72.7
70.9
70.9
63.3
55.6
34.9
70.4
63.8
53.3
44.9
71.7
68.0
61.7
45.5
70.4
63.3
53.3
3.8
WoLA 71.4
71.4
70.4
66.5
70.2
68.1
65.0
58.5
71.1
70.4
69.1
64.8
71.0
69.5
67.9
63.7
69.8
67.1
63.5
55.8
70.7
69.3
67.2
61.9
69.8
67.1
63.5
55.8
WoLA+NNM 71.3
71.2
70.3
70.9
70.7
69.4
67.3
62.4
71.3
71.0
71.0
71.0
70.8
69.6
67.4
63.7
70.4
68.4
67.2
66.7
70.9
69.9
68.7
66.9
70.4
68.4
67.2
62.4
WoLA 71.2
70.9
69.2
61.3
70.0
67.7
63.4
53.8
71.0
70.0
68.0
55.2
70.8
69.2
66.4
61.8
69.6
66.6
62.1
53.7
70.5
68.9
65.8
57.2
69.6
66.6
62.1
53.7
WoLA+NNM 71.1
70.7
69.1
68.2
70.5
68.9
65.8
47.0
71.2
70.7
69.9
68.8
70.7
69.0
66.2
61.8
70.2
67.9
65.9
60.5
70.7
69.4
67.4
61.3
70.2
67.9
65.8
47.0
MKrum - 72.8
72.2
67.1
38.5
71.0
66.9
57.9
1.9
72.8
70.6
24.4
0.8
70.8
64.5
52.7
34.9
69.8
57.7
1.6
0.7
71.4
66.4
40.7
15.4
69.8
57.7
1.6
0.7
BKT 71.8
70.1
67.1
38.5
69.3
63.1
57.9
1.9
66.8
58.3
24.4
0.8
69.0
54.9
52.7
34.9
67.2
44.7
1.6
0.7
68.8
58.2
40.7
15.4
66.8
44.7
1.6
0.7
FoundFL 61.0
58.2
55.8
41.5
57.4
54.6
46.0
1.4
42.1
20.9
5.5
0.5
58.7
55.7
49.8
49.5
39.9
1.0
0.7
0.7
51.8
38.1
31.6
18.7
39.9
1.0
0.7
0.5
NNM 72.9
72.4
67.9
44.9
71.1
67.2
58.2
4.0
73.0
72.9
72.7
3.0
70.8
64.0
54.6
34.9
70.4
63.2
50.6
1.0
71.6
67.9
60.8
17.6
70.4
63.2
50.6
1.0
WoLA 71.1
71.0
70.1
66.8
70.5
68.8
61.6
3.4
71.3
71.0
71.0
0.8
70.8
69.2
67.0
63.7
70.2
66.8
58.3
0.7
70.8
69.4
65.6
27.1
70.2
66.8
58.3
0.7
WoLA+NNM 71.2
71.1
70.2
67.4
70.6
69.3
67.2
32.2
71.3
71.0
71.0
3.2
70.8
69.4
67.3
63.7
70.4
68.3
66.9
0.7
70.9
69.8
68.5
33.4
70.4
68.3
66.9
0.7
WoLA 70.9
70.5
68.6
61.8
70.3
68.3
59.1
1.9
71.2
70.7
69.8
0.9
70.6
68.7
65.3
61.8
70.0
66.3
56.3
0.5
70.6
68.9
63.8
25.4
70.0
66.3
56.3
0.5
WoLA+NNM 71.1
70.7
68.9
63.2
70.5
68.9
65.7
22.3
71.2
70.7
69.9
2.9
70.7
68.8
65.7
61.8
70.2
67.7
65.4
0.6
70.7
69.4
67.1
30.2
70.2
67.7
65.4
0.6
GM - 73.2
71.7
67.0
50.6
69.3
63.3
53.9
39.0
72.2
69.8
64.4
0.8
70.8
61.1
45.6
41.7
66.9
57.1
15.1
1.1
70.5
64.6
49.2
26.6
66.9
57.1
15.1
0.8
BKT 72.9
72.6
67.0
50.6
71.1
67.9
53.9
39.0
69.7
66.3
64.4
0.8
71.3
66.5
45.6
41.7
69.4
62.1
15.1
1.1
70.9
67.1
49.2
26.6
69.4
62.1
15.1
0.8
FoundFL 58.4
55.6
55.6
52.5
54.6
37.2
3.6
2.0
55.9
0.4
0.3
0.4
57.0
53.9
56.2
58.5
13.5
4.6
1.0
0.7
47.9
30.4
23.3
22.8
13.5
0.4
0.3
0.4
NNM 72.9
72.7
68.2
69.8
71.1
67.4
59.3
22.6
72.9
72.4
71.3
68.6
70.8
63.7
55.7
40.7
70.3
64.1
52.3
17.0
71.6
68.1
61.3
43.7
70.3
63.7
52.3
17.0
WoLA 71.3
71.3
70.9
68.3
69.8
67.6
65.0
58.4
71.0
70.2
69.0
65.7
70.9
70.0
69.2
68.4
69.0
65.5
59.7
38.5
70.4
68.9
66.8
59.9
69.0
65.5
59.7
38.5
WoLA+NNM 71.3
71.2
70.5
70.6
70.5
68.7
65.5
59.3
71.2
70.5
69.5
68.2
70.9
69.4
68.1
65.4
70.1
67.1
63.7
51.4
70.8
69.4
67.4
63.0
70.1
67.1
63.7
51.4
WoLA 71.2
70.9
69.6
65.3
69.6
67.1
63.6
53.2
70.9
69.9
68.1
61.7
70.8
69.3
68.0
66.3
68.8
64.9
58.2
19.7
70.3
68.4
65.5
53.2
68.8
64.9
58.2
19.7
WoLA+NNM 71.1
70.8
69.3
66.6
70.4
68.2
64.1
55.4
71.1
70.2
68.6
66.6
70.7
68.9
66.5
63.2
70.0
66.6
62.3
50.8
70.7
68.9
66.2
60.5
70.0
66.6
62.3
50.8
Table 12: Accuracy (s.d.) % on Purchase100 with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 69.9
63.2
50.0
5.9
61.6
49.0
29.3
12.4
66.6
61.3
41.4
11.1
61.6
50.3
29.6
12.4
61.1
49.3
27.9
1.5
64.2
54.6
35.6
8.7
61.1
49.0
27.9
1.5
BKT 71.5
68.9
50.0
5.9
69.2
63.9
29.3
12.4
68.0
62.7
41.4
11.1
65.8
53.7
29.6
12.4
67.1
57.3
27.9
1.5
68.3
61.3
35.6
8.7
65.8
53.7
27.9
1.5
FoundFL 58.5
50.6
17.5
4.7
50.8
22.5
1.9
2.2
24.3
5.9
5.2
10.0
54.6
55.7
56.2
56.8
23.7
10.3
1.5
1.5
42.4
29.0
16.5
15.0
23.7
5.9
1.5
1.5
NNM 72.9
71.2
59.6
73.1
71.1
66.6
57.0
1.2
72.6
71.6
71.7
69.4
67.3
53.5
29.5
12.6
69.0
55.7
41.0
24.1
70.6
63.7
51.7
36.1
67.3
53.5
29.5
1.2
WoLA 70.7
69.8
67.2
60.2
68.6
65.5
60.0
39.8
70.5
69.4
66.5
53.3
69.9
67.7
63.1
57.9
68.4
64.9
59.3
37.7
69.6
67.5
63.2
49.8
68.4
64.9
59.3
37.7
WoLA+NNM 70.0
69.7
68.0
69.8
68.8
67.0
63.6
51.1
70.1
69.9
69.7
69.8
69.0
66.7
62.4
57.9
68.7
66.0
63.2
62.3
69.3
67.9
65.4
62.1
68.7
66.0
62.4
51.1
WoLA 70.5
69.3
65.2
36.6
68.4
64.6
58.1
36.4
70.2
68.9
64.6
37.7
69.7
66.5
61.1
42.0
68.1
64.1
57.8
36.9
69.4
66.7
61.4
37.9
68.1
64.1
57.8
36.4
WoLA+NNM 69.7
69.2
67.1
59.0
68.7
66.3
58.3
24.6
69.9
69.4
68.8
67.2
69.0
66.1
60.4
41.8
68.5
65.3
60.9
34.3
69.1
67.2
63.1
45.4
68.5
65.3
58.3
24.6
CwTM - 71.6
67.5
52.1
6.0
69.0
61.4
37.7
12.4
69.7
64.0
47.2
10.5
64.4
50.2
31.5
11.7
66.9
53.1
31.9
1.5
68.3
59.2
40.1
8.4
64.4
50.2
31.5
1.5
BKT 71.5
69.1
52.1
6.0
69.2
63.9
37.7
12.4
68.0
62.7
47.2
10.5
65.8
53.7
31.5
11.7
67.1
57.3
31.9
1.5
68.3
61.3
40.1
8.4
65.8
53.7
31.5
1.5
FoundFL 45.0
47.5
17.0
3.3
48.9
46.9
52.8
4.1
35.9
23.2
9.1
11.4
44.9
40.2
40.3
52.6
24.1
1.6
2.0
1.9
39.7
31.9
24.2
14.6
24.1
1.6
2.0
1.9
NNM 73.2
71.9
59.9
73.1
71.2
66.6
56.5
1.2
72.6
71.7
71.6
69.4
67.0
52.4
34.5
14.4
69.1
56.2
40.7
24.1
70.6
63.7
52.6
36.4
67.0
52.4
34.5
1.2
WoLA 70.4
70.4
67.8
59.7
68.9
66.1
60.9
39.8
69.7
68.7
66.2
53.4
69.7
67.4
63.9
57.9
68.2
64.3
58.3
37.7
69.4
67.4
63.4
49.7
68.2
64.3
58.3
37.7
WoLA+NNM 70.0
69.8
68.4
69.8
68.9
67.0
63.6
51.1
70.1
69.9
69.7
69.8
69.1
67.0
62.7
57.9
68.7
66.1
63.6
62.3
69.4
68.0
65.6
62.1
68.7
66.1
62.7
51.1
WoLA 70.2
69.9
66.7
37.0
68.7
65.4
58.8
36.4
69.4
68.1
64.7
37.5
69.6
66.4
61.4
41.7
67.9
63.4
56.5
36.9
69.2
66.6
61.6
37.9
67.9
63.4
56.5
36.4
WoLA+NNM 69.8
69.1
67.1
59.1
68.7
66.3
58.9
24.6
69.9
69.4
68.8
67.2
69.1
66.2
61.0
41.9
68.5
65.3
60.9
34.3
69.2
67.2
63.3
45.4
68.5
65.3
58.9
24.6
MKrum - 72.9
69.8
58.4
6.3
71.0
66.7
56.3
1.3
72.4
63.4
29.0
3.1
67.3
55.2
36.0
12.6
69.0
55.4
1.3
0.3
70.5
62.1
36.2
4.7
67.3
55.2
1.3
0.3
BKT 70.6
67.1
58.4
6.3
67.7
60.9
56.3
1.3
65.1
51.3
29.0
3.1
63.4
39.0
36.0
12.6
64.9
35.0
1.3
0.3
66.3
50.7
36.2
4.7
63.4
35.0
1.3
0.3
FoundFL 49.7
41.5
40.5
11.1
44.3
37.2
43.0
1.9
25.0
26.0
12.3
2.1
43.9
37.9
20.5
30.9
42.8
2.0
1.7
1.9
41.2
28.9
23.6
9.6
25.0
2.0
1.7
1.9
NNM 73.1
71.3
58.5
13.1
71.1
66.7
56.2
1.2
72.6
71.6
71.7
7.1
67.5
54.1
33.9
14.8
69.1
55.5
33.7
0.3
70.7
63.8
50.8
7.3
67.5
54.1
33.7
0.3
WoLA 69.7
69.0
67.6
60.4
68.5
64.5
55.3
3.6
70.1
69.9
69.7
2.1
69.0
67.0
61.7
57.9
68.4
64.1
25.8
0.3
69.2
66.9
56.0
24.8
68.4
64.1
25.8
0.3
WoLA+NNM 70.0
69.8
68.2
62.3
68.9
67.0
63.4
26.3
70.1
69.9
69.7
6.5
69.1
66.7
62.1
57.9
68.7
66.0
63.2
0.3
69.4
67.9
65.3
30.6
68.7
66.0
62.1
0.3
WoLA 69.5
68.5
66.6
40.7
68.3
63.6
53.4
1.4
69.9
69.4
68.8
1.7
68.8
65.9
60.1
41.8
68.1
63.2
19.5
0.5
68.9
66.1
53.7
17.2
68.1
63.2
19.5
0.5
WoLA+NNM 69.8
69.2
67.2
44.4
68.7
66.2
58.0
12.0
69.9
69.4
68.8
5.4
69.0
65.8
59.3
41.6
68.5
65.2
60.8
0.5
69.2
67.1
62.8
20.8
68.5
65.2
58.0
0.5
GM - 73.4
68.6
56.2
15.7
66.6
56.1
40.2
18.8
71.7
68.3
53.9
0.4
67.4
48.2
19.8
14.9
63.4
39.8
13.3
1.9
68.5
56.2
36.7
10.3
63.4
39.8
13.3
0.4
BKT 72.9
71.2
56.2
15.7
70.9
67.1
40.2
18.8
68.9
63.2
53.9
0.4
68.3
58.8
19.8
14.9
68.9
59.6
13.3
1.9
70.0
64.0
36.7
10.3
68.3
58.8
13.3
0.4
FoundFL 44.9
50.3
45.7
19.2
43.9
22.9
2.0
1.8
13.2
1.1
1.0
1.1
42.3
42.3
42.6
46.5
16.4
12.3
1.7
1.9
32.1
25.8
18.6
14.1
13.2
1.1
1.0
1.1
NNM 73.2
71.1
59.9
60.3
71.2
67.0
57.8
14.1
72.5
71.6
71.0
67.5
66.8
55.0
36.9
10.2
69.2
57.6
38.8
14.7
70.6
64.5
52.9
33.4
66.8
55.0
36.9
10.2
WoLA 70.1
70.0
68.4
63.6
68.0
65.1
60.7
52.1
69.8
69.2
68.0
59.8
69.3
66.3
63.9
62.6
66.8
61.8
52.6
1.6
68.8
66.5
62.7
47.9
66.8
61.8
52.6
1.6
WoLA+NNM 70.0
69.9
68.6
66.5
68.8
66.7
62.6
53.9
70.0
69.5
68.9
68.0
69.2
67.1
63.3
59.3
68.5
64.9
60.9
48.0
69.3
67.6
64.9
59.2
68.5
64.9
60.9
48.0
WoLA 69.9
69.5
67.0
50.8
67.8
64.3
57.7
42.6
69.6
68.7
66.8
53.0
69.2
65.3
61.5
44.4
66.6
60.8
49.1
13.2
68.6
65.7
60.4
40.8
66.6
60.8
49.1
13.2
WoLA+NNM 69.8
69.4
68.0
56.4
68.6
65.9
59.3
42.9
69.8
69.1
67.9
65.2
69.0
66.8
62.0
44.9
68.3
64.1
58.6
30.1
69.1
67.1
63.1
47.9
68.3
64.1
58.6
30.1
Table 13: Accuracy (s.d.) % on Purchase100 with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 61.5
42.1
8.4
2.5
47.8
30.9
14.8
3.4
54.0
38.4
11.5
2.1
48.4
30.3
5.1
2.1
44.1
21.8
1.8
1.3
51.2
32.7
8.3
2.3
44.1
21.8
1.8
1.3
BKT 69.2
62.2
8.4
2.5
65.4
55.0
14.8
3.4
64.0
52.3
11.5
2.1
57.9
35.2
5.1
2.1
62.2
40.9
1.8
1.3
63.7
49.1
8.3
2.3
57.9
35.2
1.8
1.3
FoundFL 17.2
6.0
2.9
4.9
22.3
16.2
2.9
3.1
20.0
8.8
5.4
2.4
23.8
17.0
27.0
24.2
3.1
4.9
2.9
2.4
17.3
10.6
8.2
7.4
3.1
4.9
2.9
2.4
NNM 73.3
66.1
30.2
74.4
71.3
65.8
53.1
1.1
72.7
71.4
48.4
14.3
60.4
39.5
4.7
2.1
68.8
50.6
3.1
1.8
69.3
58.7
27.9
18.7
60.4
39.5
3.1
1.1
WoLA 66.1
62.9
54.0
28.5
59.4
47.5
29.8
13.8
63.3
57.1
43.3
18.8
62.8
55.0
45.6
36.6
58.5
45.8
29.1
11.0
62.0
53.6
40.4
21.7
58.5
45.8
29.1
11.0
WoLA+NNM 65.6
65.5
60.2
66.1
63.2
59.3
48.8
2.4
65.7
65.9
65.7
66.1
63.7
57.7
49.5
36.6
61.6
53.4
39.2
21.4
64.0
60.4
52.7
38.5
61.6
53.4
39.2
2.4
WoLA 65.8
61.7
41.1
7.4
59.1
46.2
28.7
15.1
63.0
55.4
38.0
10.9
62.8
55.2
41.5
15.8
58.3
44.9
29.5
11.9
61.8
52.7
35.7
12.2
58.3
44.9
28.7
7.4
WoLA+NNM 65.2
64.3
55.3
27.6
62.9
57.9
47.1
2.3
65.3
64.7
63.4
61.7
63.7
57.5
44.2
15.7
61.3
51.4
21.9
12.8
63.7
59.2
46.4
24.0
61.3
51.4
21.9
2.3
CwTM - 69.4
59.0
15.5
2.4
65.1
50.2
20.4
3.4
65.6
50.4
15.9
2.1
57.8
38.8
4.7
2.0
61.4
31.9
2.2
1.1
63.9
46.1
11.8
2.2
57.8
31.9
2.2
1.1
BKT 69.2
62.4
15.5
2.4
65.4
55.0
20.4
3.4
64.0
52.3
15.9
2.1
57.9
35.2
4.7
2.0
62.2
40.9
2.2
1.1
63.7
49.1
11.8
2.2
57.9
35.2
2.2
1.1
FoundFL 29.9
10.1
3.9
4.3
31.9
26.2
31.1
3.6
19.4
17.8
17.9
1.7
32.8
17.9
15.7
15.0
4.8
1.6
2.9
3.0
23.8
14.7
14.3
5.5
4.8
1.6
2.9
1.7
NNM 73.5
68.8
36.9
74.4
71.2
65.8
53.1
1.1
72.8
71.6
48.2
14.3
62.3
32.9
4.7
2.2
68.7
52.0
3.6
1.7
69.7
58.2
29.3
18.7
62.3
32.9
3.6
1.1
WoLA 66.1
66.0
57.6
28.8
63.6
58.9
41.5
13.8
64.2
60.7
49.3
18.7
64.1
58.6
47.8
36.6
62.2
53.0
32.6
10.9
64.0
59.4
45.8
21.8
62.2
53.0
32.6
10.9
WoLA+NNM 65.6
65.6
61.1
66.1
63.2
59.4
49.3
2.4
65.7
65.9
65.7
66.1
63.6
59.0
46.4
36.6
61.7
53.5
40.4
21.4
64.0
60.7
52.6
38.5
61.7
53.5
40.4
2.4
WoLA 65.7
64.5
51.5
7.4
63.3
57.4
38.4
15.1
63.8
59.1
44.1
11.1
63.8
56.6
41.0
15.5
61.9
51.4
33.0
11.9
63.7
57.8
41.6
12.2
61.9
51.4
33.0
7.4
WoLA+NNM 65.3
64.4
55.7
27.4
62.9
58.0
47.5
2.3
65.3
64.7
63.4
61.7
63.5
57.2
41.7
15.7
61.4
51.6
13.7
12.8
63.7
59.2
44.4
24.0
61.4
51.6
13.7
2.3
MKrum - 73.3
68.0
26.7
2.4
71.2
66.1
53.1
1.0
71.2
53.6
35.5
10.5
60.9
33.8
6.3
2.1
69.0
50.1
1.6
0.4
69.1
54.3
24.6
3.3
60.9
33.8
1.6
0.4
BKT 68.3
60.8
26.7
2.4
65.4
58.2
53.1
1.0
61.5
43.5
35.5
10.5
54.7
10.2
6.3
2.1
61.5
31.1
1.6
0.4
62.3
40.8
24.6
3.3
54.7
10.2
1.6
0.4
FoundFL 33.7
39.7
5.0
5.4
36.4
42.7
18.5
3.0
22.4
20.8
10.6
5.8
30.8
14.7
6.4
10.7
23.5
3.2
2.9
3.0
29.4
24.2
8.7
5.6
22.4
3.2
2.9
3.0
NNM 73.5
67.6
36.4
2.6
71.2
66.0
53.0
1.1
72.7
71.7
47.2
8.9
62.3
38.1
5.7
2.1
68.7
51.9
3.1
0.4
69.7
59.1
29.1
3.0
62.3
38.1
3.1
0.4
WoLA 65.4
64.6
58.8
29.0
62.9
58.6
48.2
4.1
65.7
65.9
65.7
2.6
63.5
57.8
47.8
36.6
61.1
48.2
2.6
0.3
63.7
59.0
44.6
14.5
61.1
48.2
2.6
0.3
WoLA+NNM 65.6
65.4
60.4
36.7
63.2
59.2
48.4
2.7
65.7
65.9
65.7
5.9
63.8
58.7
46.6
36.6
61.7
53.5
40.4
0.4
64.0
60.6
52.3
16.5
61.7
53.5
40.4
0.4
WoLA 65.0
63.3
54.7
8.5
62.6
57.8
46.9
1.7
65.3
64.7
63.4
1.9
63.1
57.7
40.5
15.7
60.8
46.4
2.1
0.3
63.4
58.0
41.5
5.6
60.8
46.4
2.1
0.3
WoLA+NNM 65.3
64.3
55.2
9.8
62.9
58.0
47.1
2.1
65.3
64.7
63.4
4.8
63.6
57.4
44.3
15.5
61.4
51.4
12.6
0.2
63.7
59.1
44.5
6.5
61.4
51.4
12.6
0.2
GM - 72.1
61.5
30.7
3.2
61.3
41.2
17.2
2.7
71.8
65.4
28.6
0.6
62.0
26.4
2.7
5.1
56.7
29.2
14.8
1.1
64.8
44.7
18.8
2.5
56.7
26.4
2.7
0.6
BKT 72.8
69.1
30.7
3.2
70.7
66.0
17.2
2.7
67.6
60.2
28.6
0.6
62.3
39.3
2.7
5.1
68.5
55.6
14.8
1.1
68.4
58.0
18.8
2.5
62.3
39.3
2.7
0.6
FoundFL 27.4
17.7
5.3
8.4
36.4
9.8
3.0
3.0
22.7
2.9
16.1
3.1
30.2
29.1
22.5
17.0
23.0
4.2
2.9
3.0
28.0
12.7
10.0
6.9
22.7
2.9
2.9
3.0
NNM 73.6
67.8
37.1
45.2
71.4
66.6
55.0
1.6
72.7
71.7
48.7
47.6
62.1
37.0
3.8
2.8
68.9
53.3
2.9
1.5
69.7
59.3
29.5
19.8
62.1
37.0
2.9
1.5
WoLA 65.9
64.8
59.5
42.5
61.2
55.0
46.0
35.3
65.1
64.4
61.3
24.4
63.8
55.9
42.4
42.0
59.2
42.4
8.7
2.2
63.0
56.5
43.6
29.3
59.2
42.4
8.7
2.2
WoLA+NNM 65.7
65.9
61.5
57.3
63.3
59.5
50.0
19.7
65.5
65.5
64.8
65.1
63.7
59.6
49.6
42.6
61.7
53.6
38.6
6.7
64.0
60.8
52.9
38.3
61.7
53.6
38.6
6.7
WoLA 65.5
64.1
54.8
15.4
60.9
53.2
40.2
20.3
64.7
63.2
58.2
8.6
63.5
55.8
40.2
27.4
58.8
40.5
29.9
1.8
62.7
55.4
44.7
14.7
58.8
40.5
29.9
1.8
WoLA+NNM 65.3
64.6
57.7
24.2
63.0
58.1
47.6
13.5
65.2
64.3
62.4
60.4
63.6
58.7
47.8
17.6
61.4
52.2
21.4
2.7
63.7
59.6
47.4
23.7
61.4
52.2
21.4
2.7
Table 14: Accuracy (s.d.) % on CIFAR10 with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 69.6
55.9
35.8
19.7
68.0
59.6
51.0
28.8
66.7
56.3
32.9
18.9
62.6
45.0
18.6
18.5
63.3
51.9
35.9
17.8
66.0
53.7
34.8
20.7
62.6
45.0
18.6
17.8
BKT 70.6
64.7
35.8
19.7
69.8
68.4
51.0
28.8
65.5
58.6
32.9
18.9
66.2
57.7
18.6
18.5
69.0
63.6
35.9
17.8
68.2
62.6
34.8
20.7
65.5
57.7
18.6
17.8
FoundFL 47.7
61.9
54.1
27.6
47.2
27.0
9.6
12.5
42.0
25.3
20.0
18.9
42.8
39.3
44.5
45.7
37.5
16.8
11.0
11.1
43.5
34.1
27.8
23.1
37.5
16.8
9.6
11.1
NNM 73.3
67.5
42.9
67.2
73.1
71.0
66.5
38.7
73.6
73.3
72.4
52.8
64.2
52.5
45.2
24.5
72.6
67.7
54.8
39.9
71.4
66.4
56.4
44.6
64.2
52.5
42.9
24.5
WoLA 71.7
63.4
47.2
28.4
70.6
65.2
59.9
51.7
71.0
68.1
60.7
42.4
69.4
64.1
58.7
52.3
68.6
62.2
56.5
42.4
70.2
64.6
56.6
43.5
68.6
62.2
47.2
28.4
WoLA+NNM 73.4
70.1
57.2
69.6
73.1
71.6
67.7
37.9
73.6
73.4
72.5
72.1
70.2
65.2
60.0
53.9
72.8
69.3
62.2
58.2
72.6
69.9
63.9
58.3
70.2
65.2
57.2
37.9
WoLA 71.3
62.3
44.1
19.9
70.5
64.5
58.5
48.0
70.9
67.3
58.0
33.1
69.5
63.1
56.4
45.6
68.4
61.7
54.3
37.4
70.1
63.8
54.3
36.8
68.4
61.7
44.1
19.9
WoLA+NNM 73.3
69.5
54.3
53.5
72.9
70.9
66.2
41.5
73.5
73.0
71.7
68.9
70.6
64.7
58.0
47.0
72.7
68.7
61.1
50.6
72.6
69.3
62.3
52.3
70.6
64.7
54.3
41.5
CwTM - 73.2
64.9
43.5
20.5
72.7
69.1
56.8
29.4
71.3
61.7
39.6
18.9
63.7
52.1
25.0
18.9
72.2
59.4
41.8
19.0
70.6
61.4
41.3
21.3
63.7
52.1
25.0
18.9
BKT 70.6
64.6
43.5
20.5
69.9
68.4
56.8
29.4
65.5
58.8
39.6
18.9
65.3
58.1
25.0
18.9
69.0
63.6
41.8
19.0
68.1
62.7
41.3
21.3
65.3
58.1
25.0
18.9
FoundFL 42.6
44.3
44.2
31.1
53.0
50.8
58.0
13.7
32.5
34.1
30.1
21.7
57.0
45.7
54.6
42.7
46.0
12.3
10.9
11.0
46.2
37.4
39.6
24.0
32.5
12.3
10.9
11.0
NNM 73.5
67.9
43.5
66.9
73.4
71.6
67.0
39.2
73.8
73.4
72.3
53.4
65.8
54.6
40.8
17.9
73.1
68.8
60.5
41.1
71.9
67.3
56.8
43.7
65.8
54.6
40.8
17.9
WoLA 73.2
68.7
51.6
27.9
72.8
70.7
62.1
51.5
72.3
68.3
60.5
42.7
70.5
64.3
59.6
51.9
72.4
65.2
56.6
41.2
72.3
67.4
58.1
43.0
70.5
64.3
51.6
27.9
WoLA+NNM 73.3
70.3
57.8
69.8
73.2
71.7
68.0
38.3
73.6
73.4
72.6
72.1
70.6
65.6
60.8
52.9
73.0
69.7
62.8
58.8
72.7
70.2
64.4
58.4
70.6
65.6
57.8
38.3
WoLA 73.3
68.0
48.7
19.8
72.8
70.1
60.0
47.8
72.1
67.7
58.0
33.1
70.3
63.3
58.2
44.8
72.4
64.4
55.2
36.5
72.2
66.7
56.0
36.4
70.3
63.3
48.7
19.8
WoLA+NNM 73.3
69.6
54.9
55.4
73.0
71.2
66.6
41.5
73.7
72.9
71.7
69.0
70.6
64.4
57.3
47.1
72.9
69.1
62.7
48.6
72.7
69.4
62.6
52.3
70.6
64.4
54.9
41.5
MKrum - 73.2
66.6
39.1
17.9
73.1
70.2
66.0
28.8
73.2
65.1
33.1
9.4
64.0
50.5
40.0
24.5
72.6
64.3
31.6
9.5
71.2
63.4
42.0
18.0
64.0
50.5
31.6
9.4
BKT 67.9
60.6
39.1
17.9
66.2
63.2
66.0
28.8
64.0
51.9
33.1
9.4
63.1
49.5
40.0
24.5
65.2
52.7
31.6
9.5
65.3
55.6
42.0
18.0
63.1
49.5
31.6
9.4
FoundFL 40.4
42.5
41.4
19.2
42.1
37.1
44.3
11.0
49.9
32.4
30.3
10.7
51.8
52.1
47.2
35.5
46.3
10.7
11.0
11.0
46.1
34.9
34.8
17.5
40.4
10.7
11.0
10.7
NNM 73.5
67.7
42.8
19.1
73.5
71.4
66.4
36.9
73.8
73.5
72.7
13.6
66.3
54.2
41.4
19.4
73.1
68.4
57.8
9.5
72.0
67.0
56.2
19.7
66.3
54.2
41.4
9.5
WoLA 73.0
69.9
56.5
29.0
72.5
70.8
66.7
34.0
73.6
72.9
67.7
10.8
70.3
66.2
60.4
53.9
72.3
67.1
41.7
9.1
72.4
69.4
58.6
27.4
70.3
66.2
41.7
9.1
WoLA+NNM 73.3
70.4
57.0
33.3
73.0
71.5
67.5
36.9
73.7
73.4
72.6
14.9
70.3
64.3
59.5
51.9
72.9
69.4
61.2
7.5
72.6
69.8
63.6
28.9
70.3
64.3
57.0
7.5
WoLA 72.9
69.2
53.5
21.0
72.6
70.2
65.4
30.2
73.5
72.5
66.1
11.5
70.4
64.2
58.3
47.0
72.2
66.4
30.9
10.5
72.3
68.5
54.8
24.0
70.4
64.2
30.9
10.5
WoLA+NNM 73.3
69.6
54.7
22.3
73.1
70.8
66.3
35.8
73.6
72.9
71.8
14.9
70.7
64.0
58.3
44.9
72.8
68.8
60.2
10.3
72.7
69.2
62.2
25.6
70.7
64.0
54.7
10.3
GM - 71.4
62.7
48.1
25.0
70.0
63.3
53.1
32.3
73.4
71.0
50.8
11.5
65.5
53.0
26.5
23.2
68.3
64.5
58.6
37.9
69.7
62.9
47.4
26.0
65.5
53.0
26.5
11.5
BKT 71.8
67.5
48.1
25.0
71.3
70.7
53.1
32.3
69.1
64.7
50.8
11.5
66.7
57.3
26.5
23.2
70.3
66.8
58.6
37.9
69.8
65.4
47.4
26.0
66.7
57.3
26.5
11.5
FoundFL 42.3
41.3
44.0
37.0
54.4
36.4
11.0
11.8
36.6
16.4
23.0
18.0
40.0
43.6
52.4
48.3
48.8
31.0
10.9
11.0
44.4
33.8
28.3
25.2
36.6
16.4
10.9
11.0
NNM 73.6
68.0
44.5
63.8
73.5
71.8
67.9
45.6
73.9
73.2
71.8
51.2
63.8
51.7
43.0
24.9
73.0
69.0
59.1
49.6
71.6
66.7
57.3
47.0
63.8
51.7
43.0
24.9
WoLA 72.5
67.3
55.3
35.7
71.6
68.6
64.9
59.4
73.3
72.6
70.2
49.7
70.8
63.5
58.5
57.3
69.2
64.6
61.2
28.5
71.5
67.3
62.0
46.1
69.2
63.5
55.3
28.5
WoLA+NNM 73.4
70.6
58.7
64.0
73.1
71.9
68.8
57.9
73.6
73.3
72.1
71.2
71.0
65.6
61.7
56.2
72.9
69.8
62.6
54.9
72.8
70.2
64.8
60.8
71.0
65.6
58.7
54.9
WoLA 72.4
66.6
52.6
25.3
71.7
67.9
63.0
48.9
73.2
72.1
68.9
34.4
70.7
62.5
56.8
52.7
69.0
63.7
59.4
40.3
71.4
66.5
60.1
40.3
69.0
62.5
52.6
25.3
WoLA+NNM 73.4
69.9
56.1
42.8
73.2
71.3
67.7
49.7
73.6
72.9
71.2
67.5
70.4
64.0
60.4
50.3
72.9
69.2
60.0
46.7
72.7
69.4
63.1
51.4
70.4
64.0
56.1
42.8
Table 15: Accuracy (s.d.) % on CIFAR10 with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 63.2
46.6
25.8
12.9
60.8
51.1
37.4
18.2
53.9
33.9
22.1
15.9
46.9
15.4
16.4
15.2
50.0
34.7
23.3
13.9
55.0
36.3
25.0
15.2
46.9
15.4
16.4
12.9
BKT 65.5
55.6
25.8
12.9
64.7
62.3
37.4
18.2
60.9
52.2
22.1
15.9
58.7
37.1
16.4
15.2
63.0
53.4
23.3
13.9
62.6
52.1
25.0
15.2
58.7
37.1
16.4
12.9
FoundFL 48.4
39.2
40.3
29.0
43.6
20.5
12.7
14.1
28.6
30.5
37.1
20.4
26.1
16.7
33.2
32.8
28.0
15.6
11.3
11.3
34.9
24.5
26.9
21.5
26.1
15.6
11.3
11.3
NNM 71.1
61.3
27.1
50.4
71.4
69.3
64.6
25.6
70.7
65.2
47.9
37.4
60.7
25.1
13.0
16.3
70.9
65.0
37.6
31.1
68.9
57.2
38.0
32.2
60.7
25.1
13.0
16.3
WoLA 68.3
55.6
38.1
25.1
67.1
61.2
54.6
41.5
67.6
62.9
53.1
30.7
66.0
59.6
50.9
41.8
63.8
55.8
47.0
26.9
66.5
59.0
48.8
33.2
63.8
55.6
38.1
25.1
WoLA+NNM 71.1
66.4
48.7
64.4
71.0
69.3
65.0
28.4
71.5
71.4
70.7
70.1
66.9
59.6
53.1
43.1
70.5
65.7
56.6
48.2
70.2
66.5
58.8
50.9
66.9
59.6
48.7
28.4
WoLA 68.1
54.4
31.2
12.8
66.8
60.1
51.6
41.5
67.3
61.9
48.1
24.4
65.3
57.5
45.9
31.4
63.1
54.9
45.3
24.9
66.1
57.7
44.4
27.0
63.1
54.4
31.2
12.8
WoLA+NNM 70.9
65.8
44.0
36.9
70.6
68.5
61.6
41.8
71.3
70.8
69.2
66.1
66.5
59.6
46.0
31.7
70.4
64.9
54.0
37.5
69.9
65.9
55.0
42.8
66.5
59.6
44.0
31.7
CwTM - 70.3
56.8
34.0
12.2
69.7
64.2
47.6
18.7
67.6
51.2
25.3
16.7
54.9
20.0
15.7
15.1
68.4
47.5
26.4
14.4
66.2
47.9
29.8
15.4
54.9
20.0
15.7
12.2
BKT 65.3
55.7
34.0
12.2
64.5
62.4
47.6
18.7
60.7
52.2
25.3
16.7
59.5
38.9
15.7
15.1
63.1
53.4
26.4
14.4
62.6
52.5
29.8
15.4
59.5
38.9
15.7
12.2
FoundFL 19.1
26.3
39.8
29.0
19.5
28.6
45.7
12.5
18.6
25.9
31.1
23.0
19.3
24.3
23.1
26.7
22.8
8.7
11.3
11.3
19.8
22.8
30.2
20.5
18.6
8.7
11.3
11.3
NNM 71.4
63.1
27.6
51.1
71.6
69.9
65.3
25.8
71.2
66.6
56.7
39.1
55.1
36.8
13.1
15.4
71.1
66.5
50.4
30.9
68.1
60.6
42.6
32.5
55.1
36.8
13.1
15.4
WoLA 70.8
64.7
43.9
24.7
70.5
67.9
58.2
41.2
69.7
64.3
54.0
31.5
66.4
60.6
53.2
43.7
69.9
60.6
49.0
26.6
69.5
63.6
51.6
33.5
66.4
60.6
43.9
24.7
WoLA+NNM 71.1
67.0
48.8
64.1
71.0
69.5
65.1
31.9
71.4
71.4
70.8
70.1
66.9
59.6
52.3
44.5
70.7
67.1
59.2
48.8
70.2
66.9
59.2
51.9
66.9
59.6
48.8
31.9
WoLA 70.8
64.1
36.6
13.6
70.4
67.2
54.1
42.0
69.6
63.5
49.8
24.5
66.4
59.7
45.7
29.6
69.7
59.0
46.4
27.9
69.4
62.7
46.5
27.5
66.4
59.0
36.6
13.6
WoLA+NNM 71.0
66.4
45.4
37.3
70.8
68.6
62.5
37.3
71.4
70.9
69.2
66.2
65.8
59.5
44.6
30.7
70.5
66.2
56.1
39.2
69.9
66.4
55.6
42.1
65.8
59.5
44.6
30.7
MKrum - 70.8
59.7
25.4
10.1
71.0
69.2
64.7
24.3
70.4
57.3
40.4
15.5
61.7
26.0
19.8
16.3
69.5
64.2
19.3
8.5
68.7
55.3
33.9
14.9
61.7
26.0
19.3
8.5
BKT 61.8
51.0
25.4
10.1
59.8
57.4
64.7
24.3
58.1
41.5
40.4
15.5
53.8
18.6
19.8
16.3
58.0
43.3
19.3
8.5
58.3
42.4
33.9
14.9
53.8
18.6
19.3
8.5
FoundFL 27.7
18.1
33.3
20.8
21.5
16.3
35.0
13.9
23.4
24.6
16.3
18.3
26.9
19.0
25.8
17.6
32.9
11.6
10.0
11.3
26.5
17.9
24.1
16.4
21.5
11.6
10.0
11.3
NNM 71.2
62.5
27.8
11.1
71.5
69.3
63.8
32.1
71.3
66.1
44.1
17.9
61.2
29.8
18.4
14.5
71.0
65.3
48.0
8.4
69.2
58.6
40.4
16.8
61.2
29.8
18.4
8.4
WoLA 70.8
66.0
48.8
25.2
70.4
68.4
63.0
30.2
71.6
71.1
61.8
14.8
67.0
60.4
52.6
43.1
69.9
63.8
29.8
8.4
69.9
66.0
51.2
24.3
67.0
60.4
29.8
8.4
WoLA+NNM 71.2
66.8
49.0
27.8
70.9
69.5
64.4
25.7
71.5
71.4
70.7
18.0
67.1
60.9
51.4
40.4
70.7
66.5
56.8
8.4
70.3
67.0
58.5
24.0
67.1
60.9
49.0
8.4
WoLA 70.5
65.6
42.8
12.4
70.3
67.8
60.9
25.7
71.3
70.4
50.2
12.8
66.5
58.3
46.2
31.7
69.6
63.4
25.4
10.1
69.6
65.1
45.1
18.5
66.5
58.3
25.4
10.1
WoLA+NNM 71.0
66.2
42.8
14.7
70.7
68.6
61.9
36.0
71.3
70.8
69.3
15.9
66.6
58.9
46.5
33.0
70.3
65.6
53.9
10.2
70.0
66.0
54.9
22.0
66.6
58.9
42.8
10.2
GM - 67.4
54.1
38.9
13.1
66.1
56.7
45.9
25.3
71.2
64.6
52.9
10.0
57.8
27.2
13.7
15.0
66.1
61.4
49.5
31.0
65.7
52.8
40.2
18.9
57.8
27.2
13.7
10.0
BKT 68.3
61.1
38.9
13.1
67.9
68.1
45.9
25.3
65.3
59.6
52.9
10.0
61.2
41.6
13.7
15.0
66.6
62.2
49.5
31.0
65.8
58.5
40.2
18.9
61.2
41.6
13.7
10.0
FoundFL 20.3
16.8
31.1
27.7
25.8
17.0
11.8
14.2
20.2
22.6
18.1
18.4
23.7
16.1
25.7
21.2
47.6
12.1
11.3
11.3
27.5
16.9
19.6
18.6
20.2
12.1
11.3
11.3
NNM 71.4
63.1
29.5
54.5
71.6
70.1
65.6
41.6
71.2
66.9
47.5
28.5
58.2
33.6
16.2
15.7
71.2
66.0
47.8
35.9
68.7
59.9
41.3
35.2
58.2
33.6
16.2
15.7
WoLA 69.5
62.3
47.0
28.5
68.9
65.5
61.2
44.9
71.4
70.1
66.4
38.0
67.2
60.2
48.2
48.3
66.8
63.1
57.0
30.8
68.7
64.2
56.0
38.1
66.8
60.2
47.0
28.5
WoLA+NNM 71.3
67.2
50.2
60.7
71.0
69.7
66.5
40.5
71.6
71.3
70.3
69.3
67.4
61.3
55.6
47.9
70.8
67.5
56.7
46.8
70.4
67.4
59.9
53.0
67.4
61.3
50.2
40.5
WoLA 69.5
61.4
42.1
16.7
68.8
64.2
57.4
36.9
71.2
69.1
63.8
27.9
66.9
58.1
40.8
33.2
66.3
61.0
53.5
31.7
68.5
62.8
51.5
29.3
66.3
58.1
40.8
16.7
WoLA+NNM 71.1
66.4
45.5
33.3
70.9
69.1
63.7
43.2
71.4
70.6
68.7
65.1
66.2
59.8
46.8
36.2
70.5
66.5
55.2
37.6
70.0
66.5
56.0
43.1
66.2
59.8
45.5
33.3
Table 16: Accuracy (s.d.) % on CIFAR10 with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 44.0
32.9
15.8
10.9
43.5
40.1
21.9
11.7
25.4
21.8
22.7
11.7
15.7
12.3
16.0
13.6
29.3
20.8
16.4
11.2
31.6
25.6
18.6
11.8
15.7
12.3
15.8
10.9
BKT 57.1
37.2
15.8
10.9
56.4
47.3
21.9
11.7
51.9
35.9
22.7
11.7
45.5
19.5
16.0
13.6
54.2
28.8
16.4
11.2
53.0
33.7
18.6
11.8
45.5
19.5
15.8
10.9
FoundFL 26.6
29.7
25.9
15.8
23.6
17.2
14.0
13.5
24.6
26.4
19.5
19.2
17.8
21.3
15.2
15.5
19.2
16.7
10.3
9.2
22.4
22.3
17.0
14.6
17.8
16.7
10.3
9.2
NNM 67.1
43.8
16.9
49.3
66.6
62.5
57.6
27.0
64.8
54.7
37.0
15.4
45.0
17.7
20.7
13.6
66.3
56.6
28.3
14.8
62.0
47.1
32.1
24.0
45.0
17.7
16.9
13.6
WoLA 61.3
43.1
31.3
18.8
61.3
53.9
42.7
23.3
60.0
51.2
31.6
19.1
54.0
43.3
33.2
28.3
52.9
45.1
30.4
16.6
57.9
47.3
33.8
21.2
52.9
43.1
30.4
16.6
WoLA+NNM 66.4
58.3
38.7
46.0
66.4
64.6
56.6
23.9
67.9
67.5
66.8
57.1
58.4
50.1
42.6
29.2
65.8
57.4
37.3
36.3
65.0
59.6
48.4
38.5
58.4
50.1
37.3
23.9
WoLA 61.4
41.3
19.8
9.0
60.9
51.7
44.3
29.2
60.0
49.1
27.2
15.6
54.0
39.6
29.6
19.8
53.5
42.2
35.3
19.6
58.0
44.8
31.2
18.6
53.5
39.6
19.8
9.0
WoLA+NNM 66.0
57.3
29.3
17.0
66.2
62.5
52.9
25.0
67.6
66.3
63.5
31.0
58.3
42.9
27.8
24.9
65.6
56.5
30.7
20.3
64.7
57.1
40.9
23.7
58.3
42.9
27.8
17.0
CwTM - 63.8
40.0
20.9
10.8
62.8
48.8
33.7
11.4
59.5
32.4
22.3
11.8
37.9
18.2
13.9
14.2
60.0
21.6
14.2
11.2
56.8
32.2
21.0
11.9
37.9
18.2
13.9
10.8
BKT 57.1
37.9
20.9
10.8
56.6
47.0
33.7
11.4
52.0
36.1
22.3
11.8
41.8
17.6
13.9
14.2
54.2
28.3
14.2
11.2
52.3
33.4
21.0
11.9
41.8
17.6
13.9
10.8
FoundFL 17.9
20.6
23.8
23.3
17.8
18.8
24.0
9.8
18.7
18.6
19.1
16.9
17.0
20.0
16.6
20.6
18.1
9.9
13.5
9.2
17.9
17.6
19.4
16.0
17.0
9.9
13.5
9.2
NNM 67.7
50.2
16.7
48.4
68.0
63.7
57.6
27.4
66.5
56.6
39.5
15.6
41.6
22.6
16.4
14.4
67.4
59.5
36.6
14.6
62.3
50.5
33.4
24.1
41.6
22.6
16.4
14.4
WoLA 66.4
55.1
34.8
18.9
66.3
62.8
50.8
23.6
64.8
56.5
37.2
19.2
58.6
46.5
36.1
30.7
65.6
48.1
36.5
16.2
64.3
53.8
39.1
21.7
58.6
46.5
34.8
16.2
WoLA+NNM 67.0
58.9
38.7
43.0
66.9
65.1
59.0
24.8
67.8
67.6
66.8
56.8
59.5
48.3
41.5
31.3
66.4
61.4
49.0
34.5
65.5
60.3
51.0
38.1
59.5
48.3
38.7
24.8
WoLA 66.3
53.7
24.4
8.1
66.1
61.1
45.5
30.0
64.2
54.6
31.0
16.1
57.7
45.8
29.4
18.2
65.3
47.8
39.2
18.5
63.9
52.6
33.9
18.2
57.7
45.8
24.4
8.1
WoLA+NNM 66.5
58.1
30.7
15.3
66.5
63.5
53.3
25.1
67.5
66.3
63.6
30.7
58.9
48.6
26.5
19.6
65.9
59.1
44.9
19.2
65.1
59.1
43.8
22.0
58.9
48.6
26.5
15.3
MKrum - 65.7
42.4
17.4
5.0
66.3
61.3
57.2
22.6
61.9
53.8
37.9
22.8
46.3
22.5
18.0
13.6
64.5
50.8
13.2
7.7
60.9
46.2
28.7
14.3
46.3
22.5
13.2
5.0
BKT 53.9
37.6
17.4
5.0
53.2
48.8
57.2
22.6
48.7
38.2
37.9
22.8
35.5
13.4
18.0
13.6
51.0
35.8
13.2
7.7
48.5
34.8
28.7
14.3
35.5
13.4
13.2
5.0
FoundFL 16.8
18.8
21.5
18.4
19.2
21.4
26.5
12.0
19.1
16.8
17.8
12.3
18.7
18.9
16.4
17.9
21.3
10.0
13.4
9.2
19.0
17.2
19.1
14.0
16.8
10.0
13.4
9.2
NNM 67.5
49.4
18.1
5.9
67.7
63.0
54.2
27.2
65.8
54.8
37.7
15.3
49.4
16.0
16.0
12.4
67.2
58.3
34.2
9.0
63.5
48.3
32.0
14.0
49.4
16.0
16.0
5.9
WoLA 65.9
57.5
38.7
19.6
65.5
63.8
56.5
23.9
67.8
67.4
38.5
12.2
57.1
49.1
37.9
29.2
65.0
56.3
13.8
7.5
64.3
58.9
37.1
18.5
57.1
49.1
13.8
7.5
WoLA+NNM 66.8
58.1
38.9
22.2
66.6
64.6
55.1
23.4
67.7
67.7
66.9
14.2
59.7
49.6
41.2
33.2
66.4
60.0
35.2
7.5
65.4
60.0
47.4
20.1
59.7
49.6
35.2
7.5
WoLA 65.7
56.8
28.6
7.3
65.9
62.2
52.9
25.0
67.6
65.9
26.1
11.2
57.4
47.7
29.4
24.9
64.7
54.4
17.2
9.5
64.3
57.4
30.9
15.6
57.4
47.7
17.2
7.3
WoLA+NNM 66.4
57.4
30.0
6.9
66.4
63.4
52.1
25.2
67.6
66.3
63.2
11.0
58.9
44.8
29.9
21.8
65.9
58.2
37.8
9.4
65.0
58.0
42.6
14.9
58.9
44.8
29.9
6.9
GM - 58.1
39.8
28.1
6.1
60.6
47.9
31.0
10.5
66.0
56.5
31.1
8.2
45.5
17.7
18.6
13.5
61.0
52.5
39.2
12.2
58.2
42.9
29.6
10.1
45.5
17.7
18.6
6.1
BKT 61.7
48.2
28.1
6.1
61.4
62.0
31.0
10.5
56.7
54.4
31.1
8.2
46.1
21.5
18.6
13.5
60.1
54.5
39.2
12.2
57.2
48.1
29.6
10.1
46.1
21.5
18.6
6.1
FoundFL 16.6
16.8
21.9
21.8
22.9
14.9
14.1
10.9
15.9
17.4
19.3
16.4
16.7
16.6
19.0
18.9
32.0
13.7
13.5
13.1
20.8
15.9
17.6
16.2
15.9
13.7
13.5
10.9
NNM 67.5
48.1
20.6
53.1
67.7
63.7
58.4
35.7
66.4
56.6
38.9
16.8
46.8
15.1
13.4
13.3
67.3
60.1
32.1
19.8
63.1
48.7
32.7
27.7
46.8
15.1
13.4
13.3
WoLA 63.5
51.7
38.5
23.1
64.6
60.3
53.7
18.0
66.9
64.1
51.0
17.9
59.5
48.6
34.7
32.6
62.3
56.3
50.2
7.4
63.3
56.2
45.6
19.8
59.5
48.6
34.7
7.4
WoLA+NNM 66.9
58.8
39.9
46.8
67.0
65.0
59.3
31.7
67.9
67.5
66.1
54.7
59.3
50.7
42.9
35.1
66.4
59.9
37.3
32.4
65.5
60.4
49.1
40.2
59.3
50.7
37.3
31.7
WoLA 63.1
49.9
29.2
9.3
63.9
58.4
50.3
10.8
66.7
62.4
46.1
11.2
58.6
47.0
24.5
23.0
61.4
55.4
47.1
9.3
62.7
54.6
39.4
12.7
58.6
47.0
24.5
9.3
WoLA+NNM 66.4
58.0
30.8
15.7
66.4
63.7
54.3
27.6
67.6
66.1
62.6
38.1
58.7
46.0
35.0
20.1
66.1
58.4
35.7
21.0
65.0
58.4
43.7
24.5
58.7
46.0
30.8
15.7

### v3

@@
 Abstract
-Federated learning (FL) is a machine learning paradigm that enables multiple data holders to collaboratively train a machine learning model without sharing their training data with external parties.
-In this paradigm, workers locally update a model and share with a central server their updated gradients (or model parameters).
-While FL seems appealing from a privacy perspective, it opens a number of threats from a security perspective as (Byzantine) participants can contribute poisonous gradients (or model parameters) harming model convergence.
-Byzantine-resilient FL addresses this issue by ensuring that the training proceeds as if Byzantine participants were absent.
-Towards this purpose, common strategies ignore outlier gradients during model aggregation, assuming that Byzantine gradients deviate more from honest gradients than honest gradients do from each other.
-However, in heterogeneous settings, honest gradients may differ significantly, making it difficult to distinguish honest outliers from Byzantine ones.
-In this paper, we introduce the Worker Label Alignement Loss (WoLA), a weighted loss that aligns honest worker gradients despite data heterogeneity, which facilitates the identification of Byzantines’ gradients.
-This approach significantly outperforms state-of-the-art methods in heterogeneous settings.
-In this paper, we provide both theoretical insights and empirical evidence of its effectiveness.
-1 Introduction
-Federated learning (FL) [19, 18, 41] enables a group of machines (workers) to collaboratively train a model without sharing their private data.
-In this paradigm, a central server coordinates the process by sending the current model to all workers.
-Each worker then performs a training step on its local data and sends an update back to the server.
-Typically, this update is the gradient computed from a mini-batch [19, 3].
-The server aggregates the gradients, updates the model, and sends it back to the workers, repeating the cycle.
-In an ideal setup, the server aggregates gradients by simply averaging them.
-However, some workers may fail or behave maliciously by sending corrupted updates (also referred to as Byzantine behaviour).
-Robust training seeks to achieve model performance comparable to that achieved in the absence of Byzantine workers.
-To this end, a common strategy is to replace simple averaging with robust aggregation methods on the server side [19, 18].
-These methods are designed to reduce the impact of outliers, which are expected to include Byzantine updates.
-However, their effectiveness strongly depends on the assumption that honest workers hold homogeneous data distributions [18, 3, 33, 4].
-In practice, this assumption may not hold [3, 24, 29, 15], as each worker typically has access to only a small, potentially unrepresentative portion of the dataset.
-With heterogeneous data, the divergence among honest gradients increases, making it easier for Byzantine machines to mask their updates [3, 19, 24].
-In this setting, a robust aggregator alone is not enough to resist Byzantine attacks.
-A common approach is to add a pre-aggregation step, such as Bucketing [24] or Nearest Neighbor Mean (NNM) [24, 3, 19, 18].
-While effective under weak adversarial conditions (i.e., when the proportion of Byzantine updates is low), we show that the performance of these solutions drops significantly in stronger adversarial scenarios.
-We present a novel solution for robust FL in heterogeneous and highly adversarial settings.
-Specifically, we address label distribution skew [29, 35, 46]—a form of data heterogeneity where label distributions vary across workers, while the feature distribution for each class remains consistent across them.
-This scenario is common in practice [29, 35, 46].
-For instance, in a mobile image recognition app, users of different ages may capture different subjects, resulting in variations in label distributions across devices.
-In another case, hospitals may specialize in certain types of diseases, leading to an overrepresentation of specific labels in their local patient records.
-Label skewness leads to inconsistent training across workers, as each worker optimizes the model based on its own label distribution [24, 46, 4].
-Local training tends to overemphasize majority classes while neglecting minority ones [46, 4].
-Since these class imbalances vary across workers, local objectives become misaligned, leading to highly inconsistent updates.
-Rather than attempting to reduce variance of updates through pre-aggregation, we address the root cause: misaligned worker objectives.
-We introduce the Worker Label Alignment (WoLA) loss, a weighted loss function that aligns local training objectives.
-WoLA reweights point-wise errors for each worker to generate gradients as if all local datasets were drawn from the same label distribution.
-By steering honest gradients in a common direction, WoLA reduces variance and improves Byzantine resilience.
-WoLA is model-agnostic, requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine machines, and adds no computational cost.
-It is also straightforward to implement and can be combined with other robustness strategies.
-The only requirement is that a common training objective—represented by a probability distribution—is fixed and shared by the server once before training begins.
-For instance, this objective could be the global label distribution (i.e., the distribution obtained by merging all local datasets), derived either from prior knowledge or from local label distributions shared by the workers with the server.
-Using this information, the server estimates and broadcasts the global distribution.
-Despite the fact that Byzantine nodes can lie in their data distributions ahead of training, we show in this paper that their influence remains limited and that this approach still brings a significant gain in robustness compared to state-of-the-art approaches.
-Our contributions can be summarized as follows:
-  • We introduce WoLA, a weighted loss function designed to address label skewness in highly adversarial settings (Section 5).
-  • We provide theoretical insights for the effectiveness of WoLA, with all proofs included in the appendix (Section 6)
-  • We empirically show that WoLA significantly outperforms state-of-the-art methods through extensive experiments involving various models, datasets, aggregators, and Byzantine attacks.
-Our code is publicly available [^1] https://anonymous.4open.science/r/WoLA-E4CB to ensure reproducibility (Section 8).
-The paper is organized as follows.
-Section 2 reviews the background and related work.
-Section 4 introduces the motivation behind the WoLA.
-Section 5 formally defines the loss and explores its properties.
-Section 6 presents theoretical justifications for its design.
-Section 7 describes the experimental setup, and Section 8 presents the experimental results.
-Finally, Section 9 concludes the paper and outlines directions for future work.
-2 Background
-This section outlines key FL concepts and defines our threat model. All notations are summarized in Table 1.
-Table 1: Notation—In (c) and (d), superscript indices denote classes, while subscripts denote workers.
-Notation Meaning
-$[k]$ Set of integers from $1$ to $k$, with $k\in\mathbb{N}$
-##\## Cardinality of a set
-$\mathbbm{1}$ Indicator function. For example, $\mathbbm{1}(x\leq 0)=1$ if $x\leq 0$, and 00 otherwise.
-$\nabla$ Gradient
-[…]
-$\mathbb{E}[X]$ Expected value of the random variable $X$
-$\mathbb{E}[X|A]$ Conditional expectation of the random variable $X$ given the event $A$
-i.i.d. Independent and identically distributed
-a.s. Almost surely is a probabilistic concept meaning that an event happens with probability $1$.
-[…]
-Notation Meaning
-$\Phi$ Model
-$n$ Number of workers
-$\mathcal{H}$ Set of honest workers, with $\mathcal{H}\subset[n]$
-$f$ Number of Byzantine workers, with $f=n-\#\mathcal{H}$
-Notation Meaning
-$C$ Number of classes
-$\mathcal{D}$ Global training set, composed of pairs $(x,y)\in\mathcal{X}\times[C]$, where $x$ is an observation, and $y$ its corresponding label.
-$\mathcal{D}_{i}$ Local dataset of worker $i$, with $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$
-$N$ Size of the global dataset: $N=\#\mathcal{D}$
-$N_{i}$ Size of worker $i$’s local dataset: $N_{i}=\#\mathcal{D}_{i}$
-$N^{c}$ Number of class-$c$ instances in the global dataset, with $N=\sum_{c=1}^{C}N^{c}$
-$N_{i}^{c}$ Number of class-$c$ instances in worker $i$’s dataset, with $N_{i}=\sum_{c=1}^{C}N_{i}^{c}$
-$p$ Label distribution of $\mathcal{D}$: $p=(p^{1},p^{2},\ldots,p^{C})$ with $p^{c}=N^{c}/N$
-$p_{i}$ Label distribution of $\mathcal{D}_{i}$: $p_{i}=(p_{i}^{1},p_{i}^{2},\ldots,p_{i}^{C})$ with $p^{c}_{i}=N_{i}^{c}/N_{i}$
-Notation Meaning
-$\mathcal{L}$ Loss
-$\ell(y,\Phi(x))$ Pointwise loss of prediction $\Phi(x)$ and label $y$
-$\mathcal{L}_{i}$ Loss of worker $i$: $\mathcal{L}_{i}=\sum_{(x,y)\in\mathcal{D}_{i}}\ell\Big{(}y,\Phi(x)\Big{)}/N_{i}$, with $\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\ N_{i}/N=\mathcal{L}$
-$\mu^{c}$ Class-wise gradient of class $c$ in $\mathcal{D}$, i.e., $\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}$
-$\mu^{c}_{i}$ Class-wise gradient of class $c$ in $\mathcal{D}_{i}$, i.e., $\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}_{i}$
-$\mathcal{WL}$ Worker Label Alignement Loss (WoLA)
-$\mathcal{WL}_{i}$ WoLA of worker $i$, with $\sum_{i\in\mathcal{H}}\mathcal{WL}_{i}\ N_{i}/N=\mathcal{WL}$
-$q$ Shared training objective across workers: a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$.
-2.1 Byzantine-free FL
-We consider a distributed architecture with $n$ workers and a central server, aimed at solving a classification task with $C$ classes.
-The workers hold local datasets $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$.
-Each worker $i$ holds $N_{i}$ data points $(x,y)\in\mathcal{X}\times[C]$, where $\mathcal{X}$ is the input space, $x$ is an observation, and $y$ is its label.
-Given a model $\Phi$, the standard local loss function (not WoLA) for worker $i$ is:
+Federated Learning (FL) allows multiple data holders (workers) to collaboratively train a machine learning model without sharing their private data.
+A central server coordinates the process by distributing the global model and receiving local gradients from workers.
+However, faulty or malicious (Byzantine) workers can compromise training.
+In heterogeneous environments, even honest gradients may differ substantially, making it challenging to distinguish them from Byzantine updates.
+We propose Worker Label Alignment Loss (WoLA), a weighted loss function that promotes alignment among honest gradients, thereby enhancing the detection of Byzantine behavior.
+Our method demonstrates superior performance compared to state-of-the-art techniques in empirical evaluations and is supported by theoretical analysis.
+Codehttps://github.com/anonymous14031992/WoLA
+Introduction
+Federated learning (FL) (Guerraoui, Gupta, and Pinot 2024b, a; Wen et al.
+2023) enables multiple machines (workers) to collaboratively train a model without sharing private data.
+A central server coordinates the training by distributing the current global model to all workers.
+Each worker computes an update based on its local data and returns it to the server.
+The server aggregates these updates, updates the global model, and repeats the process.
+Typically, these updates correspond to gradients computed on local minibatches (Allouah et al.
+2023).
+Under ideal conditions, the server aggregates gradients using simple averaging.
+However, some workers may fail or behave maliciously, sending corrupted (Byzantine) updates.
+Robust training aims to achieve performance comparable to that without Byzantine failures.
+A common approach replaces averaging with robust aggregation methods on the server (Guerraoui, Gupta, and Pinot 2024b, a).
+These methods mitigate the impact of outliers, which are expected to include Byzantine updates.
+Their effectiveness, however, depends heavily on honest workers having similar data distributions (Guerraoui, Gupta, and Pinot 2024a; Allouah et al.
+2023; Peng, Li, and Ling 2024; Bao, Wu, and He 2024).
+In practice, this assumption often fails (Allouah et al.
+2023; Karimireddy, He, and Jaggi 2020; Li et al.
+2022; Fang et al.
+2025), as each worker typically accesses only a small, potentially unrepresentative subset of the data.
+In heterogeneous settings, variance among honest gradients increases, enabling Byzantine workers to better conceal their malicious updates (Allouah et al.
+2023; Guerraoui, Gupta, and Pinot 2024b; Karimireddy, He, and Jaggi 2020).
+To enhance robustness, many methods introduce a pre-aggregation step (Guerraoui, Gupta, and Pinot 2024b, a), such as Bucketing (Karimireddy, He, and Jaggi 2020) or Nearest Neighbor Mean (NNM) (Allouah et al.
+2023).
+While effective when the proportion of Byzantine updates is low, we show that their performance deteriorates significantly in stronger adversarial scenarios.
+We propose a novel approach for robust FL in heterogeneous and highly adversarial environments.
+Specifically, we focus on label distribution skew (Li et al.
+2022; Ran, Ge, and Zhong 2021; Zhang et al.
+2022)—a common form of data heterogeneity where label distributions vary across workers, but feature distributions within each class remain consistent.
+For example, in a mobile image recognition app, users of different ages may capture different subjects, leading to variation in label distributions across devices (Li et al.
+2022; Ran, Ge, and Zhong 2021; Zhang et al.
+2022).
+Label skew causes inconsistent training, as each worker optimizes the model according to its local label distribution (Karimireddy, He, and Jaggi 2020; Zhang et al.
+2022; Bao, Wu, and He 2024), producing highly divergent updates.
+Instead of reducing update variance via pre-aggregation, we tackle the root cause: misaligned worker objectives.
+We introduce the Worker Label Alignment (WoLA) loss, a weighted loss function that aligns local training objectives across workers.
+WoLA reweights point-wise errors to generate gradients as if all local datasets shared the same label distribution.
+By steering honest gradients toward a common direction, WoLA reduces variance and improves resilience against Byzantine attacks.
+Our contributions are as follows:
+  • We introduce WoLA, a weighted loss function designed to address label skew in highly adversarial FL settings.
+  • We provide theoretical analysis supporting WoLA’s effectiveness, with proofs in the appendix.
+  • We empirically demonstrate that WoLA significantly outperforms state-of-the-art methods across multiple models, datasets, aggregators, and Byzantine attack scenarios.
+Background
+This section outlines key FL concepts and defines our threat model. All notations are introduced at first use and summarized in Appendix Table 2.
+Byzantine-free FL
+We consider a distributed FL setup with $n$ workers and a central server, tasked with a classification problem involving $C$ classes.
+Each worker $i$ holds a local dataset $\mathcal{D}_{i}$ of size $N_{i}$, consisting of samples $(x,y)\in\mathcal{X}\times[C]$, where $\mathcal{X}$ is the input space, $x$ an observation, and $y$ its label.
+Given a model $\Phi$, the local loss function for worker $i$ is:
 $\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\ell\big{(}y,\Phi(x)\big{)},$
-where $\ell$ is a point-wise loss function, assumed to be differentiable with respect to the model parameters.
-In each communication round, the server broadcasts the model parameters to all workers.
-Each worker $i$ is expected to compute the gradient using its own data $\mathcal{D}_{i}$ and send back the resulting gradient $\nabla\mathcal{L}_{i}$.
-2.2 Threat Model
-In this paper, we focus on handling Byzantine workers under the common assumption of a trusted FL server [3, 24, 4].
-We consider an adversarial setting where fewer than half the workers are Byzantine, i.e., $f<n/2$.
-These malicious workers can send arbitrary updates to the server and have full knowledge of all local updates, following the threat model in [3, 24, 4].
-The server’s objective is to identify a stationary point of the average loss over the honest workers.
-That is, to find a point where $\nabla\mathcal{L}=0$, with
-$\mathcal{L}:=\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\frac{N_{i}}{N},$
-and $N=\sum_{i\in\mathcal{H}}N_{i}$. Assuming such a point exists, it can be approximated via gradient descent and robustness strategies.
-2.3 Global Dataset, Local Datasets & Label Distribution Skew
-To model data heterogeneity, we consider the following label skew scenario: all data points are drawn from independent and identically distributed (i.i.d) random variables, forming the global dataset $\mathcal{D}$.
-These points are then randomly assigned to honest workers, resulting in local datasets $\mathcal{D}_{i}$ for $i\in\mathcal{H}$, such that $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$.
-Each worker receives data with a different label distribution, introducing heterogeneity.
+where $\ell$ is a point-wise loss, assumed differentiable with respect to model parameters.
+In each communication round, the server broadcasts the current model parameters to all workers.
+Each worker $i$ computes the gradient $\nabla\mathcal{L}_{i}$ on its local data $\mathcal{D}_{i}$ and sends this update back to the server.
+Threat Model
+We assume a Byzantine setting with a trusted server and fewer than half of the workers being malicious ($f<n/2$), following the threat model in (Allouah et al.
+2023; Karimireddy, He, and Jaggi 2020; Bao, Wu, and He 2024).
+Byzantine workers can send arbitrary updates and have full knowledge of all honest updates.
+The server’s goal is to find a stationary point of the average loss over honest workers, i.e.,
+$\nabla\mathcal{L}=0,\quad\text{with }\mathcal{L}:=\sum_{i\in\mathcal{H}}\frac{N_{i}}{N}\mathcal{L}_{i},\quad N=\sum_{i\in\mathcal{H}}N_{i},$
+where $\mathcal{H}$ denotes the set of honest workers. Assuming such a stationary point exists, it can be approximated via gradient descent combined with robustness mechanisms.
+Global Dataset, Local Datasets & Data Heterogeneity
+We focus on a label skew scenario: all data points are drawn from independent and identically distributed (i.i.d) random variables, forming the global dataset $\mathcal{D}$.
+These points are randomly assigned to honest workers, producing local datasets $\mathcal{D}_{i}$ for $i\in\mathcal{H}$, such that $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$.
+Each worker receives data with a different label distribution, introducing data heterogeneity.
 Although Byzantine workers may also have their own data, we exclude them from the global dataset as they play no role in our analysis.
-The global label distribution refers to the distribution of labels in the full dataset $\mathcal{D}$, while the local label distributions refer to label distributions in a worker’s datasets $\mathcal{D}_{i}$, for $i\in\mathcal{H}$.
-2.4 Gradient Dissimilarity
-Karimireddy et al.
-[24] introduce Gradient Dissimilarity as a measure of system robustness, lower gradient dissimilarity implies stronger robustness.
-It quantifies how much individual gradients deviate from the average of the honest gradients:
-Gradient $\text{Gradient Dissimilarity}=\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\|\nabla\mathcal{L}_{i}-\nabla\mathcal{L}\|^{2},$
-Data heterogeneity tends to increase gradient dissimilarity [19, 18, 24].
-3 Related Work
+We denote the global label distribution by $p$, corresponding to $\mathcal{D}$, and the local label distribution for worker $i$ by $p_{i}$, corresponding to $\mathcal{D}_{i}$.
+Gradient Dissimilarity
+Data heterogeneity increases divergence among honest gradients (Guerraoui, Gupta, and Pinot 2024b, a; Karimireddy, He, and Jaggi 2020).
+This divergence can be measured by Gradient Dissimilarity (Karimireddy, He, and Jaggi 2020):
+$\text{Gradient Dissimilarity}=\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\|\nabla\mathcal{L}_{i}-\nabla\mathcal{L}\|^{2},$
+where $\mathcal{H}$ is the number of honest workers.
+Gradient Dissimilarity is a proxy for system robustness. Lower values indicate higher robustness.
+Related Work
 This section reviews state-of-the-art approaches relevant to our context.
-3.1 Robustness Strategies with Homogeneous Data
-To mitigate Byzantine behavior during aggregation, a common approach is to replace simple updates averaging with robust aggregation methods [34, 19].
-Examples include the coordinate-wise median (CWMed) [45], geometric median (GM) [38, 34], coordinate-wise trimmed mean (CwTM) [45], Krum [7], Multi-Krum (MKrum) [20], Minimum Diameter Averaging [20], and Comparative Gradient Elimination [22].
-These methods heavily rely on the assumption that honest workers have homogeneous data.
-Their robustness under data heterogeneity remains unclear [18, 3, 33].
-Besides robust aggregators, additional strategies can enhance robustness in homogeneous settings.
-These approaches rely on a clean dataset held by the server to evaluate and filter worker updates [44, 9, 14, 36].
-3.2 Robustness Strategies Under Data Heterogeneity
-We begin by defining Gradient Dissimilarity, a robustness measure commonly used in studies addressing Byzantine resilience under data heterogeneity [19, 18, 24].
-Next, we review state-of-the-art approaches in Byzantine FL, followed by methods developed for the Byzantine-free setting.
-Finally, we highlight the advantages of our proposed method.
-3.2.1 Byzantine FL
-A common strategy to reduce gradient variance is to apply a pre-aggregation step, such as Bucketing [24], Nearest-Neighbor Mixing (NNM) [3], or FoundationFL [15].
-These methods act before the main aggregation rule, resulting in a two-step aggregation process—for example, applying NNM before Krum.
-Other methods like BOBA [4] and RAGE [12] use different strategies and do not rely on pre-aggregation.
-Bucketing [24] randomly partitions inputs into groups (buckets), averages the inputs in each bucket, and feeds these averages into the aggregation rule.
-This random partitioning reduces the empirical variance of honest updates in expectation (i.e., on average).
-However, in some iterations, it may fail to reduce heterogeneity, giving Byzantine workers a better chance to disrupt learning [18].
-In contrast, NNM [3] is a deterministic pre-aggregation method that replaces each gradient with the average of its $n-f$ nearest neighbors (including itself).
-NNM introduces additional computational cost, which prevents its use in large-scale distributed learning tasks [18].
-FoundationFL [15] synthesizes additional updates at each training step.
-The server assigns a score to each update and replicates the highest-scoring gradient $m$ times, where $m$ is a hyperparameter set by the authors to $m=n/2$.
-A critical limitation of FoundationFL is its underlying assumption [^2] The score is based on the distance to the coordinate-wise minimum and maximum.
-Given updates $g_{1},\dots,g_{n}$, define $g_{\text{max}}=\max(g_{1},\dots,g_{n})$ and $g_{\text{min}}=\min(g_{1},\dots,g_{n})$ (coordinate-wise).
+Robustness Strategies with Homogeneous Data
+Robust aggregation methods mitigate the impact of Byzantine failures during training (Pillutla, Kakade, and Harchaoui 2022; Guerraoui, Gupta, and Pinot 2024b).
+Examples include coordinate-wise median (CWMed)(Yin et al.
+2018), geometric median (GM)(Small 1990; Pillutla, Kakade, and Harchaoui 2022), coordinate-wise trimmed mean (CwTM)(Yin et al.
+2018), Krum(Blanchard et al.
+2017), Multi-Krum (MKrum)(Guerraoui, Rouault et al.
+2018), Minimum Diameter Averaging(Guerraoui, Rouault et al.
+2018), and Comparative Gradient Elimination (Gupta, Liu, and Vaidya 2021).
+These methods strongly rely on the assumption that honest workers have homogeneous data.
+Their robustness under data heterogeneity remains unclear (Guerraoui, Gupta, and Pinot 2024a; Allouah et al.
+2023; Peng, Li, and Ling 2024).
+Beyond robust aggregators, additional strategies enhance robustness in homogeneous settings by leveraging a clean dataset held by the server to evaluate and filter worker updates (Xie, Koyejo, and Gupta 2019; Cao et al.
+2020; Fang et al.
+2020; Regatti, Chen, and Gupta 2022).
+Robustness Strategies Under Data Heterogeneity
+To reduce gradient variance caused by data heterogeneity, a common approach applies a pre-aggregation step before the main aggregation rule, leading to a two-step process.
+Examples include Bucketing (Karimireddy, He, and Jaggi 2020), Nearest-Neighbor Mixing (NNM)(Allouah et al.
+2023), and FoundationFL(Fang et al.
+2025).
+Bucketing (Karimireddy, He, and Jaggi 2020) randomly partitions inputs into groups (buckets), averages the inputs within each bucket, and then feeds these averages into the aggregation rule.
+This random partitioning reduces the variance of honest updates in expectation.
+However, in some iterations, it may fail to reduce heterogeneity, giving Byzantine workers greater opportunity to disrupt training (Guerraoui, Gupta, and Pinot 2024a).
+In contrast, NNM (Allouah et al.
+2023) is a deterministic pre-aggregation method that replaces each gradient with the average of its nfn-f nearest neighbors.
+NNM achieves optimal probabilistic robustness guarantees but has a high computational cost, limiting its scalability in large distributed systems (Guerraoui, Gupta, and Pinot 2024a).
+FoundationFL (Fang et al.
+2025) synthesizes additional updates at each training step.
+The server scores each update based on its distance to the coordinate-wise extremes and replicates the highest-scoring update approximately $n/2$ times [^1] Given updates $g_{1},\dots,g_{n}$, define $g_{\text{max}}=\max(g_{1},\dots,g_{n})$ and $g_{\text{min}}=\min(g_{1},\dots,g_{n})$ (coordinate-wise).
 The score $s_{i}$ for worker $i$ is then $s_{i}=\min(\|g_{\text{max}}-g_{i}\|,\|g_{\text{min}}-g_{i}\|)$.
-A high score means the update is far from the extremes.: extreme values correspond to Byzantine updates.
-However, this assumption is not reliable under data heterogeneity, particularly against the A Little Is Enough (ALIE) attack [5].
-BOBA [4] differs from standard robustness strategies.
-It projects client gradients onto a $C-1$ dimensional subspace using truncated Singular Value Decomposition (SVD), estimates the $C-1$ simplex from a server-side dataset, reconstructs each client’s label distribution based on their position relative to the simplex, and discards those with abnormal values.
-In addition to requiring server-side data, at each round, BOBA incurs extra cost from computing the truncated SVD of the gradient matrix, which scales with both the number of clients and the gradient dimension [^3] The computational complexity of BOBA is $\mathcal{O}(kCnd)$, where $k$ is the number of truncated SVDs computed per round (typically between 3 and 5 in thei experiments), $C$ is the number of classes, $n$ the number of workers, and $d$ the gradient dimension..
-RAGE [12] performs a weighted average using confidence scores assigned to gradients.
-These scores are computed iteratively, downweighting gradients with large projections on the principal eigenvector of the empirical covariance matrix of gradients.
-The covariance is recomputed with updated scores until the largest eigenvalue is sufficiently small.
-The computational complexity of RAGE makes it impractical for large-scale optimization [24].
-3.2.2 Byzantine-free FL
-In a Byzantine-free setting, non-IID data remains a major obstacle to model convergence.
-Several techniques have been proposed to address this challenge, including knowledge distillation [^4] Knowledge distillation encourages a student model to approximate the teacher model’s output logits, enabling it to mimic the teacher’s behavior.
-and logit calibration.
-Some methods rely on knowledge distillation and require access to an additional test set [28, 11, 10, 30].
-In contrast, [47] and [39] avoid relying on proxy datasets.
-They use a generator trained on the server or client to enable data-free knowledge distillation.
-However, these approaches require significant communication and coordination, and introduce substantial computational overhead [40, 21].
-Other approaches introduce personalized loss functions based on local data distributions, including DMFL [35], FedLC [46], and FedDisk [32].
-DMFL and FedLC, inspired by [8], adjust logits [^5] Logits are the outputs of the final classification layer and inputs to the softmax.
-before applying softmax cross-entropy, increasing margins for minority classes.
-These methods require the tuning of a hyperparameter with no bounded range.
-FedDisk uses loss with sample-level weights, learned from a global and multiple local density estimators, independent of class labels.
-All these propositions either require an additional tes set, introduce significant overhead [40, 21, 32], or rely on hyperparameter tuning—a particularly difficult task in our setting: (i) without a test set, local tuning is possible, but data heterogeneity may cause inconsistent values; and (ii) since these methods assume a Byzantine-free setting, no robust protocol is provided to handle malicious behavior.
-3.2.3 Our proposition
-To the best of our knowledge, using a loss function to improve Byzantine resilience under data heterogeneity is a novel approach.
-WoLA is a model-agnostic method that requires no hyperparameters, no server-side dataset, and no knowledge of the number of Byzantine machines.
-It adds no computational cost, is easy to implement, and can be combined with other robustness strategies.
-Its only requirement is a shared training objective.
-While many approaches focus on handling update heterogeneity [3, 24, 4, 15, 12], WoLA tackles its root cause—label skewness—by aligning gradients from the start.
-This enables WoLA to significantly outperform state-of-the-art methods.
-In the next section, we explain the motivation behind its design.
-4 Rationale for Worker Label Alignement Loss
-This section first shows that local gradients approximate weighted averages of class-wise gradients.
-Next, it illustrates that class-wise gradients can differ significantly.
-Finally, it presents the main conclusion drawn from these observations.
-4.1 Gradients as Weighted Averages
-The key rationale behind WoLA is that the gradient over a sufficiently large local dataset can be approximated as a weighted average of class-wise gradients, with weights given by the local label distribution.
-More precisely, the gradient of worker $i$ for a standard loss is:
-$\nabla\mathcal{L}_{i}$ $=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
-$=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}$
-$=\sum_{c=1}^{C}\frac{N^{c}_{i}}{N_{i}}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}},$
-where:
-$\frac{N^{c}_{i}}{N_{i}}=\Bigg{(}\text{Proportion of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=p^{c}_{i},$
-and
-$\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}}$ (1)
-$\quad\quad=\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=\mu_{i}^{c}$
-Given a sufficient number of instances of class $c\in[C]$, the following approximation is reasonable (see Appendix Section A.2, Lemma 2 for a formal statement):
-(Mean gradient of class $\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}$ (2)
-$\quad\quad\approx\Bigg{(}\text{Mean gradient of class }c\text{ in}\mathcal{D}\Bigg{)}$
-$\quad\quad=\frac{\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}}\vcentcolon=\mu_{c}$
-We call $\mu_{i}^{c}$ the class-$c$ gradient of worker $i$, and $\mu^{c}$ the global class-$c$ gradient. When clear from context, we omit the term global.
-Using approximation equation (2), we can express the local gradient as a weighted average of global class gradients:
-$\nabla\mathcal{L}_{i}\approx\sum_{c=1}^{C}p_{i}^{c}\mu^{c}$ (3)
-Assuming that the vectors $\mu^{c}$ for $c\in[C]$ differ—either in norm or direction—the differences in local gradients arise from the local label distributions, through the weights $p_{i}^{c}$.
-In the next subsection, we show that the vectors $\mu^{c}$ for $c\in[C]$ are indeed different.
-4.2 Difference in Class Gradients
-Figure 1 shows standard (non-FL) training of a shallow neural network on the Iris dataset [17].
-The loss decreases and accuracy improves over epochs, indicating successful model convergence.
-Independently of the optimization, at each step we compute the class gradients for Setosa, Versicolour, and Virginica, and track the cosine similarity between these gradients over training steps.
-The figure reveals that the angles between class-wise gradients often exceed $\pi/2$, indicating that the gradients point in widely different directions.
-In some cases, such as between Versicolour and Virginica, the gradients are nearly opposite.
-This highlights that different classes can push the model in distinct—and sometimes conflicting—directions during training. Similar observations have been reported in previous studies [46, 4].
-[…]
-4.3 Key Takeaways
-Local gradients approximate weighted averages of global class gradients, with weights given by the local label distribution.
-Since class gradients may differ, aligning local gradients requires aligning their weightings—as if local datasets shared the same label distribution.
-WoLA is designed to achieve this goal.
-5 Worker Label Alignement Loss
+A high score means the update is far from the extremes..
+This scoring assumes that extreme values are Byzantine but becomes unreliable when data is heterogeneous (Baruch, Baruch, and Goldberg 2019).
+Other methods, such as BOBA (Bao, Wu, and He 2024) and RAGE (Data and Diggavi 2021), adopt different strategies without relying on pre-aggregation.
+However, they require access to a server-side dataset and are computationally expensive, limiting their practicality in large-scale optimization (Karimireddy, He, and Jaggi 2020) [^2] At each round, BOBA requires computing several times the truncated SVD of the gradient matrix, which scales with both the number of clients and the gradient dimension..
+Our proposition
+To the best of our knowledge, employing a loss function to improve robustness under data heterogeneity is novel.
+WoLA is model-agnostic, requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine workers, and adds no computational overhead.
+It is straightforward to implement and can be combined with other robustness strategies.
+WoLA only requires a common training objective shared by the server once before training begins.
+While many approaches address update heterogeneity (Allouah et al.
+2023; Karimireddy, He, and Jaggi 2020; Bao, Wu, and He 2024; Fang et al.
+2025; Data and Diggavi 2021), WoLA tackles its root cause—label skewness—by aligning gradients from the start.
+Worker Label Alignement Loss
 This section introduces WoLA.
-It first defines the training objective required for WoLA, then presents its formula.
-The final subsections provide an intuitive understanding of WoLA’s properties.
-These properties are rigorously stated in Section 6 and proved in the Appendix A.
-5.1 Shared Training Objective
-WoLA requires a shared training objective across workers.
-This objective is formalised by a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$.
-A larger $q^{i}$ means that point-wise errors for samples with label $i$ are given more weight during training, and vice versa.
-To define this distribution, we considered three situations:
+We first present the rationale behind its design, then provide its formal definition and training objective.
+The final subsections describe the theoretical framework and the properties of WoLA.
+Proofs and additional details are provided in the Appendix.
+Rationale for WoLA
+We begin by observing that local gradients are weighted averages of class-wise gradients:
+$\nabla\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}=\sum_{c=1}^{C}p_{i}^{c}\mu^{c}_{i},$
+where $p^{c}_{i}\vcentcolon=N^{c}_{i}/N_{i}$ is the proportion of class $c$ in $\mathcal{D}_{i}$, and
+$\mu_{i}^{c}\vcentcolon=\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbf{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}}$
+is the average gradient for class $c$ in worker $i$’s local dataset.
+With sufficient samples per class, local class-wise gradients $\mu_{i}^{c}$ approximate the global class-wise gradients $\mu^{c}$:
+$\nabla\mathcal{L}_{i}=\sum_{c=1}^{C}p_{i}^{c}\mu^{c}_{i}\approx\sum_{c=1}^{C}p_{i}^{c}\mu^{c},$ (1)
+where the global class-wise gradient is defined as:
+$\mu_{c}\vcentcolon=\frac{\sum_{(x,y)\in\mathcal{D}}\mathbf{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}}$
+In general, global class-wise gradients are not equal (Zhang et al. 2022; Bao, Wu, and He 2024); we illustrate this in the Appendix.
+Assuming the above approximation in (1) holds, differences in local gradients primarily stem from differences in local label distributions (via $p_{i}^{c}$).
+Hence, aligning local gradients requires aligning local label distributions or simulating such alignment.
+WoLA is specifically designed to achieve this goal.
+Training Objective & WoLA Definition
+WoLA requires a training objective $q=(q^{1},q^{2},\ldots,q^{C})$ with $\sum_{i\in[C]}q^{i}=1$ and $q^{i}\geq 0$, which is shared with all workers once before training begins.
+This vector $q$ represents the simulated local label distribution of each worker when using WoLA.
+In practice, a higher $q^{i}$ increases the weight given to errors on class $i$ during local training.
+The distribution $q$ may approximate the true global label distribution or be set arbitrarily. We present three settings for defining $q$:
+Local label distributions are shared
+Each worker communicates its local label distribution to the server, which computes and broadcasts the global label distribution.
+However, this introduces a potential vulnerability: Byzantine workers can manipulate the training objective by reporting false distributions.
+We analyze this attack in the Appendix (see Section Training Objective Attack).
 Prior knowledge
@@
 For example, in tumor detection on MRI scans or crop type recognition from satellite imagery, national statistics can serve as a reliable proxy for label distributions.
-Local label distributions are shared
-Each worker shares its local label distribution with the server before training.
-The server then computes the global distribution.
-In this case, it is important to note that (i) the process reduces data privacy, and (ii) it introduces a new opportunity for Byzantine workers to attack the system.
-We carefully evaluate this attack scenario in Section 7.2.
-Regarding privacy concerns, we assume in this paper that the server is trusted as we consider privacy as an orthogonal problem that is actively being investigated in the literature.
-Nevertheless, off-the-shelf SecureAggregation protocols such as [6] could be easily used to release the trusted server assumption.
 Consensus among workers on the training objective
-All workers agree that the global model should favor certain classes during training, regardless of the actual global label distribution.
-For example, in a task to detect hateful or illegal content on social networks—where hate content is rare—it may be agreed to assign higher importance to these cases.
-In medical diagnosis, rare diseases may also be given more weight.
-In contrast, for automated recruitment systems, all classes may be weighted equally to avoid algorithmic bias.
-The training objective $q$ is shared by the server with all workers once before training begins, allowing them to adjust their local loss functions accordingly.
-5.2 WoLA definition
-WoLA locally adjusts the weights of point-wise errors to simulate an identical label distribution across workers. More precisely, WoLA for worker $i$, denoted $\mathcal{WL}_{i}$, is defined as:
-$\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\ell\Big{(}y,\Phi(x)\Big{)},$
-While the formula is undefined for $p^{y}_{i}=0$, it is not problematic, as it indicates that label $y$ is absent from the local dataset $\mathcal{D}_{i}$, meaning the corresponding weight never appears in the sum.
-A sample from class $c$ is overweighted when class $c$ is underrepresented locally compared to the training objective, i.e., when $p_{i}^{c}<q^{c}\Leftrightarrow 1<q^{c}/p_{i}^{c}$; and underweighted otherwise.
-5.3 Local Updates Alignement
-WoLA makes local datasets appear as if they had identical label distributions matching the training objective $q$—that is, as if $p_{i}^{c}=q^{c}$ for all $c\in[C]$ and all $i\in\mathcal{H}$.
-Specifically, for worker $i$, the gradient under WoLA is:
-$\nabla\mathcal{WL}_{i}\vcentcolon=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
-$=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\frac{q^{c}}{p^{c}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
-$=$ $\sum_{c=1}^{C}q^{c}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}y,\Phi(x)\Big{)}}{N^{c}_{i}},$
-using approximation in equation (2), it follows:
-$\nabla\mathcal{WL}_{i}\approx\sum_{c=1}^{C}q^{c}\mu^{c}$ (4)
-With WoLA, the approximation of the worker $i$’s gradient no longer depends on local label proportions.
-It becomes a weighted average of class gradients, using the shared training objective $q$ as weights.
-Since equation (4) holds for all $i\in[C]$, it follows that $\nabla\mathcal{WL}_{i}\approx\nabla\mathcal{WL}_{j}$ for all $j\in[C]$.
-This suggests that local updates are more similar under WoLA, potentially reducing gradient dissimilarity and improving robustness.
-Figure 2 illustrates the impact of WoLA on updates using a fictitious example with three classes: Dog, Flower, and Frog, under the training objective $q=(1/3,1/3,1/3)$.
-It shows three local datasets with label distributions $(3/6,2/6,1/6)$, $(2/6,2/6,2/6)$, and $(1/6,1/6,4/6)$.
-Two of them differ from the training objective $q$.
-WoLA addresses this mismatch by upweighting samples from local minority classes and downweighting those from local majority classes, thereby aligning updates.
-(a) Point-wise standard loss gradients on $\mathcal{D}_{1}$
-(b) Point-wise standard loss gradients on $\mathcal{D}_{2}$
-(c) Point-wise standard loss gradients on $\mathcal{D}_{3}$
-(d) Point-wise WoLA gradients on $\mathcal{D}_{1}$
-(e) Point-wise WoLA gradients on $\mathcal{D}_{2}$
-(f) Point-wise WoLA gradients on $\mathcal{D}_{3}$
-Figure 2: Fictitious FL classification task with three classes (Dog, Flower, Frog) under the training objective $q=(1/3,1/3,1/3)$, comparing standard loss with WoLA.
-Each image represents a data point from a local dataset.
-Arrows indicate 2D point-wise loss gradients; in black squares, arrows show the resulting updates.
-Subfigures (a–c) display gradients under the standard loss, while (d–f) show gradients with WoLA.
-In (d) and (f), dataset sizes are scaled to reflect the effect of WoLA weights, although datasets are actually unchanged.
-For example, $\mathcal{D}_{3}$ has a local label distribution of $(1/6,1/6,4/6)$, while $q$ is uniform.
-This results in weights: $1/3\div 1/6=4/2$ for Dog and Flower, and $1/3\div 4/6=1/2$ for Frog.
-To avoid fractional images, the denominator 2 is omitted, giving 4 duplicates of Dog and Flower gradients, and 1 of Frog.
-The same logic applies to $\mathcal{D}_{1}$.
-Subfigures (b) and (e) are identical, as $\mathcal{D}_{2}$ label distribution matches the training objective.
-Local updates with WoLA are more aligned than with the standard loss.
-5.4 Targeted Model
-Assuming that the training objective $q$ matches the global label distribution $p$, i.e., $p=q$, the average gradients of the standard loss and the WoLA loss—denoted $\nabla\mathcal{L}$ and $\nabla\mathcal{WL}$ respectively—are similar.
-Using the approximations from equations (3) and (4), we have:
-$\nabla\mathcal{L}$ $=\sum_{i\in\mathcal{H}}\nabla\mathcal{L}i\frac{N_{i}}{N}$
-$\approx\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p_{i}^{c}\mu^{c}\frac{N_{i}}{N}=\sum_{c=1}^{C}p^{c}\mu^{c}=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p^{c}\mu^{c}\frac{N_{i}}{N}$
-$=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}q^{c}\mu^{c}\frac{N_{i}}{N}\approx\sum_{i\in\mathcal{H}}\nabla\mathcal{WL}_{i}\frac{N_{i}}{N}=\nabla\mathcal{WL}$
-Therefore, when $p=q$, the training targets the same model regardless of whether the standard or WoLA loss is used.
-5.5 Key Takeaways
-WoLA requires a shared training objective.
-Previous analyses suggest that: (i) using WoLA reduces gradient dissimilarity, and (ii) when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
-6 Theoretical Arguments
-This section uses probabilistic tools to rigorously state the properties of WoLA.
-We first describe the data modeling framework, then present asymptotic results that offer insights into WoLA’s behavior.
-6.1 Data Modeling
-Let $(X_{1},Y_{1},Z_{1}),(X_{2},Y_{2},Z_{2}),\ldots,(X_{N},Y_{N},Z_{N})$ be independent and identically distributed (i.i.d.) random triplets drawn from a distribution $\mathfrak{D}$ over the sample space $\mathcal{X}\times[C]\times\mathcal{H}$.
-Each pair $(X_{i},Y_{i})$ represents a datapoint in the dataset $\mathcal{D}$.
+All workers may agree on a predefined target distribution, regardless of the true label distribution. For example, in fair hiring systems, classes may be weighted equally to mitigate algorithmic bias.
+WoLA adjusts point-wise error contributions to simulate a common label distribution $q$ across all workers.
+Definition 1
+The local WoLA loss for worker $i$, denoted $\mathcal{W}\!\mathcal{L}_{i}$, is defined as:
+$\mathcal{W}\!\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\ell\Big{(}y,\Phi(x)\Big{)},$
+A sample from class $c$ is up-weighted when class $c$ is underrepresented in the local dataset compared to the training objective $q$, and down-weighted otherwise.
+This formula is directly derived from Importance Sampling (Tokdar and Kass 2010); further details are provided in the Appendix.
+Before establishing the theoretical properties that explain WoLA’s efficiency, we introduce the theoretical framework.
+Theoretical Framework
+Let $(X_{1},Y_{1},Z_{1}),(X_{2},Y_{2},Z_{2}),\ldots,(X_{N},Y_{N},Z_{N})$ be independent and identically distributed (i.i.d.) samples from a distribution $\mathfrak{D}$ over $\mathcal{X}\times[C]\times\mathcal{H}$.
+Each pair $(X_{i},Y_{i})$ represents a datapoint in $\mathcal{D}$.
 For each $k\in[N]$, the random variable $Z_{k}$ takes values in $\mathcal{H}$ and indicates the worker to which the $k$-th datapoint is assigned.
-We also denote by $(X,Y,Z)$ a generic triplet with the same distribution $\mathfrak{D}$.
-For any triplet $(X_{k},Y_{k},Z_{k})$, we assume that $X_{k}$ and $Z_{k}$ are conditionally independent given $Y_{k}$.
-That is, conditioned on the label, the features provide no information about the assignment, and vice versa.
-Formally, $P(X_{k},Z_{k}|Y_{k})=P(X_{k}|Y_{k})P(Z_{k}|Y_{k}).$
-6.2 Theoretical Results
-The results are asymptotic, considering the case where the size of the global dataset tends to infinity, i.e., $N\to\infty$, and are based on the Law of Large Numbers.
-Full proofs and details are given in Appendix A.
-Let $L_{i}$ and $W_{i}$ be the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$, respectively.
+Let $(X,Y,Z)$ be a generic triplet with the same distribution $\mathfrak{D}$.
+For any triplet $(X_{k},Y_{k},Z_{k})$, we assume $X_{k}$ and $Z_{k}$ are conditionally independent given $Y_{k}$, i.e., $P(X_{k},Z_{k}|Y_{k})=P(X_{k}|Y_{k})P(Z_{k}|Y_{k})$.
+Given the label, features provide no information about assignments, and vice versa.
+This implies $P(X_{k}|Y_{k},Z_{k})=P(X_{l}|Y_{l},Z_{l})$ for any $k,l$, so the feature distribution is identical across workers.
+WoLA Properties
+Let $L_{i}$ and $W_{i}$ be the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{W}\!\mathcal{L}_{i}$, respectively.
 Let $L$ and $W$ denote the average gradients of honest workers for the standard loss and WoLA, respectively.
-With the standard loss, local gradients remain different even with an infinite amount of data:
+Without WoLA, local gradients remain different even with an infinite amount of data.
 Proposition 1
 Under the standard loss, the gradient of worker $i$ converges almost surely as $N\to\infty$:
-[…]
+$L_{i}\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Z=i\Big{]}$
 Moreover, the average gradient satisfies:
-[…]
-[…]
-In contrast, with WoLA, local gradients become asymptotically identical, regardless of the training objective shared by the server:
+$L\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
+Under label skewness, and since class-wise gradients differ, there exists at least one worker $i$ for which the expected gradient differs from the global expectation:
+$\mathbb{E}[\nabla\ell(Y,\Phi(X))|Z=i]\neq\mathbb{E}[\nabla\ell(Y,\Phi(X))]$
+This implies $\lim L_{i}\overset{\text{a.s.}}{\neq}\lim L$ as $N\to\infty$, so the gradient dissimilarity remains strictly positive.
+With WoLA, local gradients become asymptotically identical, regardless of the training objective shared by the server.
 Proposition 2
 Let $q$ be the training objective distribution shared by the server. Under WoLA, the gradient of worker $i$ converges almost surely as $N\to\infty$:
-[…]
+$W_{i}\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
 Moreover, the average gradient also satisfies:
-[…]
-This directly implies that, under WoLA, the gradient dissimilarity vanishes asymptotically as $N\to\infty$ (see Appendix A.5).
-Finally, if the training objective matches the global label distribution, i.e., $q=p$, then the standard loss gradient and the WoLA gradient become asymptotically identical:
+$W\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
+As a result, the gradient dissimilarity under WoLA vanishes as $N\to\infty$.
+When the training objective matches the global label distribution (i.e., $q=p$), the WoLA and non-WoLA gradients become asymptotically identical.
 Proposition 3
 Let $q$ be the training objective and $p$ the global label distribution. If $q=p$, then:
-[…]
-This equality implies that, in the asymptotic regime, when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
-(a) Results under ALIE attack.
-(b) Results under LF attack.
-[…]
-(a) Boxplot of accuracy gains based on Table 2.
-(b) Boxplot of accuracy gains under the strongest attack, based on Table 3.
-[…]
-Table 2: Global Performance: Test Accuracy (s.d.) %, averaged over training.
-Results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF).
+$\lim_{N\to\infty}W\overset{\text{a.s.}}{=}\lim_{N\to\infty}L$
+In the asymptotic regime, when the training objective matches the global label distribution, WoLA and non-WoLA gradients target the same model.
+Table 1: Test accuracy (± standard deviation) in %, averaged over training, on Fashion MNIST and CIFAR10.
+Global results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF).
+Worst-case results correspond to the worst attack, averaged across aggregators.
 Colors rank values within columns: yellow (1st), blue (2nd), red (3rd).
 See full table in Appendix.
-(a) MNIST
-Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
-Accuracy
-- 93.6 (0.3) 92.9 (0.4) 86.1 (1.6) 66.3 (4.5) 93.6 (0.3) 92.2 (0.6) 82.3 (3.3) 43.5 (11.7) 91.9 (1.8) 84.6 (4.0) 58.8 (12.1) 17.3 (7.3)
-BKT 93.5 (0.3) 92.8 (0.5) 86.1 (1.6) 66.3 (4.5) 93.6 (0.4) 92.5 (0.7) 82.3 (3.3) 43.5 (11.7) 93.3 (0.7) 89.6 (1.9) 58.8 (12.1) 17.3 (7.3)
-FoundFL 89.6 (1.8) 74.4 (4.2) 56.8 (4.6) 47.5 (5.0) 85.4 (5.2) 69.6 (7.3) 57.7 (6.7) 50.4 (6.8) 61.6 (15.9) 50.4 (12.8) 43.0 (12.1) 38.0 (11.8)
-NNM 93.7 (0.2) 93.3 (0.3) 92.7 (0.4) 75.5 (6.9) 93.9 (0.3) 93.0 (0.6) 90.1 (2.5) 63.2 (9.0) 93.9 (0.6) 91.5 (1.4) 67.5 (14.5) 35.2 (8.4)
-WoLA 93.7 (0.3) 93.4 (0.3) 92.7 (0.3) 81.9 (0.6) 93.9 (0.3) 93.5 (0.4) 92.4 (0.4) 80.5 (2.2) 93.6 (0.5) 93.0 (0.7) 88.9 (2.6) 54.5 (8.0)
-WoLA+NNM 93.7 (0.3) 93.5 (0.3) 93.2 (0.3) 84.3 (0.6) 94.0 (0.3) 93.6 (0.4) 93.2 (0.3) 83.1 (2.8) 93.7 (0.6) 93.4 (0.7) 90.0 (3.0) 64.6 (8.7)
-WoLA 93.7 (0.3) 93.3 (0.3) 92.5 (0.3) 81.3 (0.8) 93.8 (0.3) 93.3 (0.3) 92.1 (0.6) 76.8 (5.4) 93.6 (0.6) 92.9 (0.7) 88.1 (2.8) 47.3 (17.0)
-WoLA+NNM 93.7 (0.3) 93.4 (0.3) 93.0 (0.3) 83.8 (0.7) 93.9 (0.4) 93.5 (0.3) 92.8 (0.4) 73.5 (13.0) 93.7 (0.6) 93.3 (0.7) 89.8 (2.8) 48.7 (15.8)
-Accuracy Gain Over Top Non-WoLA Methods
-WoLA 0.0 0.0 0.0 +6.4 0.0 +0.4 +2.4 +17.3 -0.4 +1.5 +21.4 +16.5
-WoLA+NNM 0.0 +0.2 +0.6 +8.8 +0.1 +0.6 +3.1 +19.9 -0.2 +1.8 +22.5 +26.6
-WoLA 0.0 0.0 -0.2 +5.9 -0.1 +0.3 +2.0 +13.6 -0.4 +1.4 +20.5 +9.3
-WoLA+NNM 0.0 +0.1 +0.3 +8.3 -0.1 +0.5 +2.8 +10.3 -0.2 +1.8 +22.3 +10.7
-(b) Fashion MNIST
+(a) Fashion MNIST (Global)
 Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
@@
 WoLA+NNM +0.1 +1.1 +3.6 +9.4 +0.2 +0.6 +5.0 +6.9 -1.1 +0.9 +15.1 +7.7
-(c) Purchase100
-Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
-Accuracy
-- 70.4 (0.7) 65.3 (1.1) 50.5 (1.9) 24.8 (2.8) 67.9 (1.0) 58.0 (2.0) 37.1 (3.1) 8.0 (2.1) 62.2 (1.8) 44.4 (3.3) 15.9 (2.3) 2.6 (1.1)
-BKT 70.1 (0.9) 64.6 (1.2) 50.5 (1.9) 24.8 (2.8) 68.2 (1.1) 59.3 (1.9) 37.1 (3.1) 8.0 (2.1) 64.5 (1.9) 49.3 (3.5) 15.9 (2.3) 2.6 (1.1)
-FoundFL 52.2 (3.6) 36.5 (3.1) 28.4 (2.6) 22.4 (2.4) 38.8 (5.9) 28.9 (4.1) 20.7 (4.5) 13.3 (4.9) 24.6 (8.2) 15.6 (6.6) 10.3 (4.3) 6.4 (2.8)
-NNM 71.6 (0.7) 67.9 (0.9) 61.3 (1.8) 38.1 (2.3) 70.6 (0.9) 63.9 (1.6) 52.0 (2.8) 28.3 (4.0) 69.6 (1.2) 58.8 (3.2) 28.9 (5.6) 15.1 (3.2)
-WoLA 70.6 (0.6) 69.2 (0.6) 66.7 (0.7) 52.7 (1.0) 69.2 (0.8) 67.0 (0.8) 61.3 (1.2) 43.1 (1.7) 63.2 (1.1) 57.1 (2.1) 43.6 (2.6) 21.8 (2.5)
-WoLA+NNM 70.9 (0.6) 69.8 (0.6) 68.3 (0.7) 57.6 (0.9) 69.3 (0.8) 67.8 (0.7) 65.3 (1.1) 53.5 (1.7) 64.0 (1.1) 60.6 (1.7) 52.6 (2.5) 33.0 (3.7)
-WoLA 70.5 (0.6) 68.8 (0.5) 65.2 (0.6) 48.2 (1.7) 69.0 (0.8) 66.3 (0.9) 59.3 (1.6) 33.5 (6.1) 62.9 (1.0) 55.9 (1.9) 40.9 (3.4) 11.2 (5.5)
-WoLA+NNM 70.7 (0.6) 69.3 (0.5) 67.0 (0.6) 53.3 (2.6) 69.2 (0.7) 67.2 (0.8) 63.1 (1.5) 39.9 (9.9) 63.7 (1.0) 59.3 (1.7) 45.7 (5.2) 19.5 (7.6)
-Accuracy Gain Over Top Non-WoLA Methods
-WoLA -1.0 +1.3 +5.4 +14.6 -1.4 +3.1 +9.3 +14.8 -6.4 -1.7 +14.6 +6.8
-WoLA+NNM -0.8 +1.8 +7.0 +19.5 -1.3 +3.9 +13.3 +25.2 -5.6 +1.8 +23.7 +17.9
-WoLA -1.2 +0.8 +4.0 +10.2 -1.6 +2.4 +7.2 +5.2 -6.7 -2.9 +11.9 -3.9
-WoLA+NNM -0.9 +1.4 +5.7 +15.3 -1.5 +3.2 +11.1 +11.6 -5.9 +0.5 +16.7 +4.5
-(d) CIFAR10
+(b) CIFAR10 (Global)
 Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
@@
 WoLA+NNM +0.9 +2.5 +5.9 +6.7 +1.2 +7.1 +14.7 +8.3 +2.2 +9.5 +10.2 -1.2
-Table 3: Worst-case Performance: Test Accuracy (s.d.) % under the worst attack, averaged over training.
-Results are averaged across aggregators (CWMed, CwTM, GM, MKrum).
-Colors rank values within columns: yellow (1st), blue (2nd), red (3rd).
-See full table in Appendix.
-(a) MNIST
-Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
-Accuracy
-- 93.4 (0.3) 91.9 (0.5) 73.6 (5.1) 27.5 (5.8) 93.3 (0.4) 90.1 (1.3) 62.8 (8.1) 12.6 (5.7) 89.3 (3.2) 70.7 (8.6) 29.5 (11.5) 7.1 (3.4)
-BKT 93.2 (0.5) 91.5 (0.9) 73.6 (5.1) 27.5 (5.8) 93.2 (0.5) 90.8 (1.5) 62.8 (8.1) 12.6 (5.7) 92.6 (1.1) 84.6 (4.0) 29.5 (11.5) 7.1 (3.4)
-FoundFL 80.4 (4.4) 36.5 (8.2) 10.5 (1.2) 10.9 (1.7) 76.4 (15.1) 27.1 (8.7) 10.9 (1.2) 13.5 (3.3) 48.5 (22.4) 20.5 (6.3) 14.7 (4.0) 16.9 (7.6)
-NNM 93.6 (0.2) 92.8 (0.3) 91.3 (0.6) 47.0 (27.5) 93.8 (0.3) 91.6 (1.7) 83.9 (8.9) 22.5 (11.7) 93.7 (0.6) 85.6 (3.7) 49.6 (24.0) 18.1 (5.9)
-WoLA 93.5 (0.3) 92.9 (0.3) 90.6 (0.4) 66.7 (0.7) 93.8 (0.3) 92.9 (0.3) 89.9 (0.7) 64.3 (3.0) 93.3 (0.5) 92.1 (0.8) 77.6 (8.1) 28.8 (6.2)
-WoLA+NNM 93.6 (0.3) 93.2 (0.3) 92.6 (0.2) 70.5 (1.6) 93.9 (0.3) 93.2 (0.3) 92.3 (0.4) 66.7 (4.7) 93.6 (0.5) 92.6 (0.7) 78.9 (11.2) 24.1 (11.4)
-WoLA 93.5 (0.3) 92.8 (0.3) 90.6 (0.4) 67.3 (1.0) 93.7 (0.3) 92.8 (0.3) 89.6 (0.8) 63.7 (8.9) 93.3 (0.5) 92.0 (0.8) 75.9 (8.1) 24.5 (7.3)
-WoLA+NNM 93.6 (0.3) 93.1 (0.3) 92.3 (0.4) 70.3 (0.7) 93.7 (0.3) 93.1 (0.3) 91.9 (0.5) 51.4 (27.9) 93.6 (0.6) 92.6 (0.8) 80.2 (9.5) 14.9 (6.6)
-Accuracy Gain Over Top Non-WoLA Methods
-WoLA -0.1 +0.1 -0.8 +19.7 -0.1 +1.3 +6.0 +41.9 -0.4 +6.5 +28.1 +10.7
-WoLA+NNM 0.0 +0.3 +1.3 +23.5 0.0 +1.6 +8.3 +44.3 -0.1 +7.0 +29.3 +6.0
-WoLA -0.1 0.0 -0.8 +20.3 -0.1 +1.3 +5.7 +41.3 -0.4 +6.4 +26.3 +6.5
-WoLA+NNM +0.1 +0.2 +0.9 +23.2 -0.1 +1.5 +7.9 +28.9 -0.1 +7.0 +30.6 -3.2
-(b) Fashion MNIST
+(c) Fashion MNIST (Worst-case)
 Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
@@
 WoLA+NNM +0.3 +2.6 +10.3 +25.4 +1.2 +2.3 +16.5 +16.7 +0.6 +2.5 +21.4 -3.5
-(c) Purchase100
-Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
-Accuracy
-- 68.2 (0.6) 59.2 (0.8) 29.0 (1.2) 8.7 (2.3) 64.0 (1.5) 48.5 (2.9) 18.5 (3.1) 0.9 (0.3) 54.9 (2.5) 28.5 (3.8) 2.1 (0.8) 0.9 (0.3)
-BKT 68.6 (0.8) 58.1 (0.9) 29.0 (1.2) 8.7 (2.3) 65.8 (1.6) 50.3 (3.1) 18.5 (3.1) 0.9 (0.3) 58.2 (3.4) 30.0 (7.1) 2.1 (0.8) 0.9 (0.3)
-FoundFL 33.6 (1.9) 4.2 (1.7) 0.6 (0.3) 0.4 (0.1) 21.5 (6.8) 2.6 (1.3) 1.6 (0.9) 1.6 (0.9) 13.2 (4.7) 3.2 (1.1) 2.9 (0.3) 2.5 (1.3)
-NNM 70.4 (0.5) 63.3 (1.5) 52.2 (1.4) 6.4 (1.0) 67.1 (1.7) 53.8 (4.3) 33.6 (4.4) 3.2 (1.8) 61.8 (3.1) 36.9 (8.1) 3.2 (2.3) 1.0 (0.4)
-WoLA 69.7 (0.6) 66.7 (0.6) 61.4 (0.7) 37.7 (1.0) 68.0 (0.8) 63.8 (0.8) 49.0 (1.4) 19.3 (1.2) 60.3 (1.0) 47.3 (2.3) 18.2 (2.6) 6.1 (2.8)
-WoLA+NNM 70.3 (0.6) 68.0 (0.6) 66.3 (0.6) 44.2 (0.8) 68.7 (0.8) 65.8 (0.6) 62.0 (1.5) 37.6 (1.9) 61.7 (1.0) 53.5 (1.5) 39.7 (3.3) 3.0 (1.8)
-WoLA 69.5 (0.6) 66.2 (0.5) 59.8 (0.6) 31.9 (1.9) 67.7 (0.8) 62.9 (0.8) 45.7 (2.5) 21.6 (4.6) 59.9 (0.9) 45.8 (2.0) 23.4 (4.5) 4.2 (1.0)
-WoLA+NNM 70.2 (0.6) 67.5 (0.5) 64.8 (0.4) 36.3 (6.4) 68.5 (0.7) 65.0 (0.7) 58.4 (1.6) 20.0 (13.5) 61.4 (0.9) 51.6 (1.6) 17.4 (15.2) 1.9 (0.7)
-Accuracy Gain Over Top Non-WoLA Methods
-WoLA -0.7 +3.4 +9.2 +29.0 +0.8 +10.0 +15.4 +16.1 -1.5 +10.4 +15.1 +3.6
-WoLA+NNM -0.1 +4.7 +14.1 +35.5 +1.5 +12.0 +28.4 +34.4 -0.1 +16.6 +36.5 +0.5
-WoLA -0.9 +2.8 +7.6 +23.2 +0.5 +9.1 +12.1 +18.4 -1.8 +8.9 +20.3 +1.7
-WoLA+NNM -0.2 +4.2 +12.6 +27.7 +1.3 +11.2 +24.8 +16.8 -0.4 +14.8 +14.2 -0.6
-(d) CIFAR10
+(d) CIFAR10 (Worst-case)
 Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
@@
 WoLA+NNM +5.4 +8.6 +13.0 +14.8 +7.5 +25.4 +28.0 +12.5 +13.0 +27.5 +13.1 +2.0
+(a) Boxplot of Global accuracy gains.
+(b) Boxplot of Worst-case accuracy gains.
+Figure 1: Accuracy gains over the best non-WoLA methods under low, medium, and high heterogeneity settings ($\alpha=3$, $1$, and $0.3$, respectively), based on the Global and Worst-case subtables reported in Tables 1 (Fashion MNIST & CIFAR10) and 4 (MNIST & Purchase100).
+Legend: $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM
+Experimental Setup
+This section outlines our experimental setup, including the training objectives of WoLA, baseline methods, and robust aggregators.
+We describe the datasets and heterogeneity regimes, the distributed system configuration, Byzantine attack scenarios, evaluation metrics, optimization algorithms, models, and hyperparameters.
+Training Objective, Baselines & Aggregators
+We assume the goal is to learn a model that performs well on a test set matching the global label distribution.
+Accordingly, the ideal WoLA training objective aligns with this distribution, i.e., $q=p$.
+We compare WoLA with four baselines: no pre-aggregation (-), Bucketing (BKT) (Karimireddy, He, and Jaggi 2020), Nearest-Neighbor Mixing (NNM) (Allouah et al.
+2023), and FoundationFL (FoundFL) (Fang et al.
+2025).
+We also evaluate WoLA combined with NNM.
+Methods requiring a server-side dataset or incurring substantial computational overhead(Karimireddy, He, and Jaggi 2020; Data and Diggavi 2021) are excluded.
+To evaluate robustness to training objective attacks, we introduce WoLA, a worst-case scenario where Byzantine workers submit highly misleading label distributions.
+A simple robust aggregator is used to counter the attack (see Appendix, Section Training Objective Attack).
+We also assess WoLA combined with NNM.
+All methods are tested with four robust aggregators: CWMed, CwTM, GM, and MKrum.
+Datasets & Heterogeneity
+We conduct experiments on four datasets: MNIST (LeCun et al.
+1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), CIFAR10 (Krizhevsky 2009), and Purchase100 (Shokri et al.
+2017).
+Due to space constraints, most results for MNIST and Purchase100 are provided in the Appendix.
+To simulate data heterogeneity across workers, we sample from the global dataset using a Dirichlet distribution with concentration parameter $\alpha$, following the method in (Allouah et al.
+2023; Hsu, Qi, and Brown 2019).
+We consider three heterogeneity levels—low, medium, and high—corresponding to $\alpha\in\{3,1,0.3\}$.
+Each worker holds a local dataset of fixed size, equal to the global dataset size divided by the number of honest workers.
+Distributed system & Byzantine attacks
+We consider a distributed system with $n=17$ workers, among which $f\in\{2,4,6,8\}$ are Byzantine.
+We evaluate five state-of-the-art gradient attacks: Fall of Empires (FOE) (Xie, Koyejo, and Gupta 2020), A Little is Enough (ALIE) (Baruch, Baruch, and Goldberg 2019), Sign Flipping (SF) (Allen-Zhu et al.
+2020), Label Flipping (LF) (Allen-Zhu et al.
+2020), and Mimic (Karimireddy, He, and Jaggi 2020).
+In the Mimic attack, all Byzantine workers replicate a carefully selected honest update.
+Inspired by ALIE (Baruch, Baruch, and Goldberg 2019), we select the honest outlier most surrounded by other honest updates (see Appendix).
+This makes Byzantine updates harder to detect while remaining effective in heterogeneous settings.
+Tables 1416 in Appendix in Appendix show that Mimic is often the most harmful attack for baseline methods such as –, BKT, and NNM.
+Metrics & Results
+We evaluate performance using two metrics averaged over training: Test Accuracy and Gradient Dissimilarity.
+Each experimental scenario—defined by dataset, heterogeneity level, defense method, and attack—is averaged over three runs with seeds $\{1,2,3\}$.
+Comprehensive results are provided in Appendix and summarized in the following tables: Table 1 presents Test Accuracy averaged over training for Fashion-MNIST and CIFAR10.
+The global accuracy corresponds to the mean performance across all aggregators (CWMed, CwTM, GM, MKrum) and all attacks (ALIE, FOE, LF, Mimic, SF).
+The worst-case accuracy refers to the lowest accuracy across attacks, averaged over aggregators.
+Table 4 provides analogous results for MNIST and Purchase100.
+Table 1 and 4 include Accuracy Gain Over Top Non-WoLA Methods, computed as the difference between each WoLA variant and the best-performing non-WoLA baseline.
+These gains are also visualized as boxplots in Figures 1(a) and 1(b).
+Table 3 shows Gradient Dissimilarity averaged over training, showing the mean across all aggregators and attacks.
+Due to space constraints, Table 3 and Table 4 are provided in the Appendix.
+Optimization & Models
+WoLA relies on approximating the global class-wise gradients with local class-wise gradients.
+The more data per class available at each optimization step, the better this approximation.
+As a result, the best performance is likely achieved using non-stochastic gradient descent, where honest workers compute gradients over the entire dataset.
+However, in practice, stochastic methods are more common.
+To align with this setting, we follow (Allouah et al.
+2023) and adopt the distributed stochastic heavy ball method (see Algorithm in Appendix).
+This method has shown strong robustness against Byzantine behavior in both heterogeneous (Allouah et al.
+2023) and homogeneous settings (El Mhamdi, Guerraoui, and Rouault 2021; Karimireddy, He, and Jaggi 2021; Farhadkhani et al.
+2022).
+We use momentum $\beta=0.9$ for MNIST and Fashion-MNIST, and $\beta=0.99$ for CIFAR10 and Purchase100.
+We use cross-entropy as point-wise loss, batch size $b=128$, learning rates that decay over time, gradient clipping at $5$, and $\ell_{2}$ regularization of $10^{-4}$.
+We train a lightweight CNN for $T=800$ steps on MNIST and Fashion-MNIST.
+For CIFAR10, we use a deeper CNN, and for Purchase100, a fully connected network.
+Both are trained for $T=2000$ steps (details in Appendix).
+Experimental Results
+In this section, we present experiments to answer the following Research Questions (RQ). We use WoLA methods to refer collectively to WoLA, WoLA, and their combinations with NNM.
+RQ 1 Are WoLA and WoLA more robust overall than state-of-the-art strategies?
+RQ 2 Are WoLA and WoLA more robust in worst-case scenarios than state-of-the-art strategies?
+RQ 3 Does combining WoLA or WoLA with NNM improve robustness?
+In Appendix, we also answer to questions:
+RQ A Does WoLA guide training toward a better model?
+RQ B What explains the robustness of WoLA methods?
+RQ C How do batch size and momentum affect WoLA performance?
+Global Robustness of WoLA & WoLA (RQ 1)
+According to Tables (a) and (b) in Table 1, WoLA and WoLA significantly improve robustness.
+The performance gains over the best state-of-the-art methods are particularly significant when the number of Byzantines is large.
+When $f=2$: WoLA and WoLA show similar performance, slightly below the best non-WoLA strategies. Typical negative gains between 1.1-1.1 and 0 points (Figure 1(a)).
+When $f=4$: WoLA and WoLA still perform similarly and yield slightly better results than non-WoLA methods, with typical gains of a few points (Figure 1(a)).
+According to Table 1, WoLA achieves gains of up to +5.4+5.4 (CIFAR10, $\alpha=0.3$).
+When $f=6$: WoLA slightly surpasses WoLA, and both outperform other methods.
+WoLA shows typical gains between +3.3+3.3 and +12.1+12.1 points, while WoLA reaches between +1.8+1.8 and +8.1+8.1 (Figure 1(a)).
+Table 1 reports maximum gains of +16.6+16.6 (WoLA) and +10.3+10.3 (WoLA) on Fashion MNIST with $\alpha=0.3$.
+When $f=8$: WoLA is better than WoLA, and both methods clearly outperform all others (Figure 1(a)).
+WoLA typically achieves gains between +5.6+5.6 and +14.7+14.7 points (median around +10.1+10.1, Figure 1(a)), while WoLA ranges from 3.7-3.7 to +7+7 (median around +5.2+5.2, Figure 1(a)).
+An exception occurs with CIFAR10, where both methods perform worse than NNM: 2.1-2.1 for WoLA and 6.2-6.2 for WoLA.
+Importantly, when $f>2$, Table 1 shows that WoLA almost always ranks in the top three regardless of the scenario (exceptions: Fashion MNIST $\alpha=0.3$, $f=2$; and CIFAR10 $\alpha=0.3$, $f=2,8$).
+While WoLA rarely ranks among the top values, it usually outperforms non-WoLA strategies.
+Worst-Case Robustness of WoLA & WoLA (RQ 2)
+According to Tables (c) and (d) in Table 1, Considering the strongest attack for each approach, WoLA and WoLA consistently enhance robustness, almost always outperforming other state-of-the-art methods, with gains reaching up to +28.8+28.8 points (CIFAR10, $\alpha=0.3$, $f=4$, WoLA).
+For $f=2$: WoLA and WoLA perform similarly, slightly surpass non-WoLA methods with typical gains of a few points (Figure 1(b)), and up to +11.3+11.3 points on CIFAR10, $\alpha=0.3$ (Table 1).
+For $f=4$, WoLA and WoLA continue to perform similarly, both outperforming all baselines with median gains of approximately +6.3+6.3 and +4.6+4.6 points, respectively (Figure 1(b)).
+WoLA achieves gains of up to +28.8+28.8 (CIFAR-10, $\alpha=0.3$), while WoLA reaches up to +27.0+27.0 under the same setting.
+For $f=6$ or 88: WoLA outperforms WoLA, and both clearly surpass all other methods (Figure 1(b)).
+WoLA achieves median gains around +13.4+13.4 points, WoLA around +12.4+12.4 (Figure 1(b)).
+Table 1 shows that WoLA achieves gains of up to +28.8+28.8, and WoLA up to +27.0+27.0 (CIFAR-10, $\alpha=0.3$, $f=4$).
+Robustness Gains from Combining WoLA & WoLA with NNM (RQ 3)
+Combining WoLA or WoLA with NNM significantly improves robustness, consistently outperforming WoLA, WoLA or NNM alone.
+WoLA+NNM consistently achieves the top or top-three performance in Table 1, reflecting strong robustness across settings.
+Gains over the best non-WoLA methods can reach around $20$ points.
+For example, on Fashion MNIST with $\alpha=0.3$ and $f=6$, +19.8+19.8; and on CIFAR10 with $\alpha=1$ and $f=6$, +18.5+18.5.
+The median gain is +10.4+10.4 when $f=6$ and +17.8+17.8 when $f=8$ (Figure 1(a)).
+Considering worst attacks, improvements are even larger, with median gains around +20+20 for both $f=6$ and $f=8$ (Figure 1(b)).
+WoLA+NNM also regularly reaches top-three values in Tables 1, confirming its robustness.
+It performs similarly to WoLA+NNM when $f\leq 4$, and slightly worse when $f>4$, but still outperforms WoLA alone and all non-WoLA strategies (Figure 1(a)1(b)).
+Importantly, it turns nearly all negative gains of WoLA alone into positive ones (Table 1).
+For instance, on CIFAR10 with $\alpha=3$ and $f=8$, the score improves from 4.4-4.4 to +6.7+6.7; with $\alpha=1$ and $f=8$, from 3.6-3.6 to +8.3+8.3; and with $\alpha=0.3$ and $f=2$, from 0.5-0.5 to +2.2+$2.$ Overall, median gains over non-WoLA methods are around +8+8 when $f=6$ or $8$ (Figure 1(a)).
+Considering worst attacks, the median gain is around +13+13 for both $f=6$ and $f=8$ (Figure 1(b)).
+Conclusion and Future Work
+We proposed WoLA, a simple and effective loss reweighting method to improve robustness in federated learning under label skew.
+Unlike existing approaches that act after gradient computation, WoLA addresses the root cause of gradient divergence by aligning worker objectives through a shared training distribution.
+It can be combined with existing defenses like NNM for additional gains.
+Empirically, WoLA significantly outperforms state-of-the-art methods: on CIFAR-10 with medium heterogeneity ($\alpha=1$, $f=6$), it improves accuracy by up to +11.3+11.3 points over the best non-WoLA baseline.
+On Fashion MNIST ($\alpha=0.3$, $f=6$), the gain is +16.6+16.6 points.
+Combined with NNM, WoLA+NNM achieves up to +18.5+18.5 on CIFAR-10 ($\alpha=1$, $f=6$), and +19.8+19.8 points more on Fashion MNIST ($\alpha=0.1$, $f=6$).
+Worst-case scenarios show even larger gains.
+Theoretically, we prove that WoLA aligns honest gradients asymptotically, reducing dissimilarity to zero even under label skew. This holds regardless of the training objective.
+Future work includes improving robustness further by refining the weighting scheme—e.g., at the mini-batch level—and designing stronger defenses against manipulations of the training objective.
+Appendix A Notations
+All notations used in the paper are summarized in Table 2.
+Table 2: Notation—In (c) and (d), superscript indices denote classes, while subscripts denote workers.
+Notation Meaning
+$[k]$ Set of integers from $1$ to $k$, with $k\in\mathbb{N}$
+#\# Cardinality of a set
+$\mathbf{1}$ Indicator function
+$\nabla$ Gradient
+$\|\cdot\|$ $\ell_{2}$-norm
+$\mathbb{E}[X]$ Expected value of a random variable $X$
+$\mathbb{E}[X|\cdot]$ Conditional expectation of $X$
+i.i.d. Independent and identically distributed
+a.s. Probabilistic property known as almost surely
+[…]
+Notation Meaning
+$\Phi$ Model
+$n$ Number of workers
+$\mathcal{H}$ Set of honest workers, with $\mathcal{H}\subset[n]$
+$f$ Number of Byzantine workers, $f=n-\#\mathcal{H}$
+Notation Meaning
+$C$ Number of classes
+$\mathcal{D}$ Global training set composed of observation-label pairs $(x,y)\in\mathcal{X}\times[C]$
+$\mathcal{D}_{i}$ Local dataset of worker $i$, with $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$
+NN Size of the global dataset: $N=\#\mathcal{D}$
+$N_{i}$ Size of worker $i$’s local dataset: $N_{i}=\#\mathcal{D}_{i}$
+$N^{c}$ Number of class-$c$ instances in the global dataset, with $N=\sum_{c=1}^{C}N^{c}$
+$N_{i}^{c}$ Number of class-$c$ instances in worker $i$’s dataset, with $N_{i}=\sum_{c=1}^{C}N_{i}^{c}$
+$p$ Label distribution of $\mathcal{D}$: $p=(p^{1},p^{2},\ldots,p^{C})$ with $p^{c}=N^{c}/N$
+$p_{i}$ Label distribution of $\mathcal{D}_{i}$: $p_{i}=(p_{i}^{1},p_{i}^{2},\ldots,p_{i}^{C})$ with $p^{c}_{i}=N_{i}^{c}/N_{i}$
+Notation Meaning
+$\mathcal{L}$ Loss
+$\ell(y,\Phi(x))$ Pointwise loss of prediction $\Phi(x)$ and label $y$
+$\mathcal{L}_{i}$ Worker $i$ loss: $\mathcal{L}_{i}=\sum_{(x,y)\in\mathcal{D}_{i}}\frac{\ell(y,\Phi(x))}{N_{i}}$, with $\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\ N_{i}/N=\mathcal{L}$
+$\mu^{c}$ Class-wise gradient of class $c$ in $\mathcal{D}$, i.e., $\sum_{(x,y)\in\mathcal{D}}\mathbf{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}$
+$\mu^{c}_{i}$ Class-wise gradient of class $c$ in $\mathcal{D}_{i}$, i.e., $\sum_{(x,y)\in\mathcal{D}_{i}}\mathbf{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}_{i}$
+$\mathcal{WL}$ Worker Label Alignement Loss (WoLA)
+$\mathcal{W}\!\mathcal{L}_{i}$ Worker $i$ WoLA, $\sum_{i\in\mathcal{H}}\mathcal{W}\!\mathcal{L}_{i}\ N_{i}/N=\mathcal{WL}$
+$q$ Shared training objective: a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, i.e., $\sum_{i\in[C]}q^{i}=1$ and $q^{i}\geq 0$.
+Appendix B Difference in Class-Wise Gradients
+This section illustrates that class-wise gradients differ. Similar observations have been reported in previous studies (Zhang et al. 2022; Bao, Wu, and He 2024).
+Figure 2 shows standard (non-FL) training of a shallow neural network on the Iris dataset (Fisher 1936).
+The loss decreases and accuracy improves over epochs, indicating successful model convergence.
+Independently of the optimization, at each step we compute the class-wise gradients Setosa, Versicolour, and Virginica, and track the cosine similarity between these gradients over training steps.
+The figure shows that angles between class-wise gradients often exceed $\pi/2$, meaning the gradients point in very different directions.
+For Versicolour and Virginica, the gradients are nearly opposite.
+This highlights that different classes can push the model in distinct—and sometimes conflicting—directions during training.
+Figure 2: Training a shallow neural network on the Iris dataset using batch gradient descent.
+The dataset is split into 70% training and 30% testing.
+While the model successfully learns Iris patterns, the class gradient directions vary significantly between classes.
+Legend: $\bullet\!$ Loss $\bullet\!$ Accuracy Gradient Cosine Similarity — $\bullet\!$ Setosa vs.
+Versicolour $\bullet\!$ Setosa vs.
+Virginica $\bullet\!$ Versicolour vs.
+Virginica
+Appendix C Importance Sampling
+The WoLA formula is derived from Importance Sampling, a Monte Carlo method that approximates an expectation under a target distribution by a weighted average of samples drawn from a different distribution (Tokdar and Kass 2010).
+In the following, we provide the detailed derivation of WoLA based on Importance Sampling.
+Let $(X,Y)$ be random variables with joint density $f$, where realizations represent points in the global dataset $\mathcal{D}$.
+Let $q$ be the marginal distribution of $Y$.
+Similarly, let $(X_{i},Y_{i})$ be random variables with joint density $f_{i}$, representing points in the local dataset $\mathcal{D}_{i}$, and $p_{i}$ the marginal distribution of $Y_{i}$.
+The label distribution skewness assumption states that the conditional distributions of features given labels are identical across workers, i.e., $X|Y$ and $X_{i}|Y_{i}$ follow the same law.
+From this, we have:
+$\frac{f(x,y)}{\int f(x,y)dx}=\frac{f_{i}(x,y)}{\int f_{i}(x,y)dx}$
+$\Leftrightarrow$ $\frac{f(x,y)}{q(y)}=\frac{f_{i}(x,y)}{p_{i}(y)}$
+$\Leftrightarrow$ $\frac{f(x,y)}{f_{i}(x,y)}=\frac{q(y)}{p_{i}(y)},$
+which leads to:
+$\mathbb{E}\Big{(}\nabla\ell(Y,\Phi(X))\Big{)}$ $=\int\nabla\ell(y,\Phi(x))f(x,y)dxdy$
+$=\int\nabla\ell(y,\Phi(x))\frac{f(x,y)}{f_{i}(x,y)}f_{i}(x,y)dxdy$
+$=\int\nabla\ell(y,\Phi(x))\frac{q(y)}{p_{i}(y)}f_{i}(x,y)dxdy$
+$=\mathbb{E}\Big{(}\nabla\ell(Y_{i},\Phi(X_{i}))\frac{q(Y_{i})}{p_{i}(Y_{i})}\Big{)}$
+As a result, to estimate $\mathbb{E}\big{(}\nabla\ell(Y,\Phi(X))\big{)}$ using data drawn from the distribution of $(X_{i},Y_{i})$, we have to estimate $\mathbb{E}\big{(}\nabla\ell(Y_{i},\Phi(X_{i})){q(Y_{i})}/{p_{i}(Y_{i})}\big{)}$.
+This corresponds to the probabilistic form of the WoLA formula.
+Appendix D Additional Research Questions
+This section addresses additional research questions.
 Table 3: Gradient Dissimilarity (s.d.), averaged over training.
@@
 WoLA+NNM 0.44 (0.02) 0.39 (0.02) 0.37 (0.04) 0.35 (0.05) 0.63 (0.04) 0.62 (0.05) 0.6 (0.08) 0.53 (0.2) 0.82 (0.07) 0.96 (0.15) 0.92 (0.22) 1.07 (0.47)
-7 Experimental Setup
-This section describes our experimental setup, including the baseline methods, the WoLA approach under training objective attack, and the datasets used.
-We detail the configuration of the heterogeneity regimes, workers, and the types of Byzantine attacks we consider.
-We also present the robust aggregation methods, evaluation metrics, optimization algorithms, models, and hyperparameters we used.
-Finally, we explain how we ensure the reproducibility of our experiments.
-7.1 Baseline Approaches
-We compare WoLA against three baselines: Bucketing (BKT) [24], NNM [3], and FoundationFL (FoundFL) [15].
-BKT requires setting a bucket size $s$.
-Following the authors’ recommendation, we use $s=\left\lfloor{n}/({2f})\right\rfloor$.
-When $f>n/4$, the bucket size drops to $1$, making BKT ineffective.
-For instance, with $17$ workers, BKT has no effect when $f=6$ or $8$.
-FoundFL requires a hyperparameter $m$.
-As suggested in the original paper, we set $m=n/2$, i.e., $m=8$ in all our experiments.
-We also include a baseline using only a robust aggregator without any additional defense, denoted as -.
-We evaluate different variants of WoLA.
-In particular, to assess robustness against training objective attacks, we implement the worst-case scenario described in Subsection 7.2, denoting the resulting loss as WoLA.
-Additionally, we examine the combination of WoLA and WoLA with NNM.
-We do not compare WoLA with methods that (i) require a server-side dataset [4, 28, 11, 10, 30], (ii) involve high computational costs [4, 12, 32, 47, 39], or (iii) rely on hyperparameter tuning [46, 35].
-7.2 WoLA under Training Objective Attack
-To define the training objective, we consider a scenario where workers share their local label distributions with the server.
-In this case, Byzantine machines may try to manipulate the objective by sending malicious distributions.
-We analyze the impact of such attacks.
-Just as Byzantine workers can send arbitrary updates with full knowledge of honest gradients, we assume they can also send arbitrary distributions with full knowledge of local label distributions.
-Rather than listing all possible attacks, we focus on one that causes the maximum distortion: Byzantine machines send a zero vector except for a single entry set to $1$, corresponding to the minority class in the global dataset (proof are provided in Appendix A.7).
-To defend against attacks, we propose using the geometric median as the aggregator.
-This approach offers two key advantages: (i) Unlike methods like Trimmed Mean or Krum, which exclude outliers, the geometric median treats all updates equally.
-This is important because outlier distributions are not necessarily malicious.
-Excluding them may discard honest distribution and degrade performance.
-(ii) When aggregating probability distributions, the geometric median produces a probability distribution—unlike the coordinate-wise median, for instance—ensuring the training objective retains its expected form.
-Our goal is to show that a strong attack, even when countered by a simple defense, has limited impact on model performance.
-As the chosen attack represents a worst-case scenario, it provides a meaningful benchmark for evaluating training objective attacks.
-We refer to our approach in this scenario as WoLA.
-7.3 Datasets, Heterogeneity & Workers
-We conduct the experiments on four datasets: MNIST [27], Fashion-MNIST [42], CIFAR10 [26], and Purchase100 [37].
-Purchase100 contains shopping records for thousands of customers.
-We use a simplified version [37], where each customer is represented by $600$ binary features indicating item purchases.
-The data is clustered into $100$ classes, and the task is to predict each customer’s class.
-To simulate data heterogeneity across workers, we sample from the global dataset using a Dirichlet distribution with concentration parameter $\alpha$, following the method in [3, 23].
-We consider three heterogeneity levels—low, medium, and high—corresponding to $\alpha\in\{3,1,0.3\}$.
-We use $n=17$ workers in all experiments.
-The number of Byzantine workers varies as $f\in\{2,4,6,8\}$.
-Each worker holds a local dataset of fixed size, equal to the total dataset size divided by the number of honest workers.
-7.4 Byzantine Attacks & Robust Aggregators
-Byzantine workers carry out one of five state-of-the-art gradient attacks: Fall of Empires (FOE) [43], A Little is Enough (ALIE) [5], Sign Flipping (SF) [2], Label Flipping (LF) [2], and Mimic [24].
-In the Mimic attack, all Byzantine workers replicate a selected honest update and send it to the server.
-Inspired by the ALIE attack—which places Byzantine values between the honest mean and outliers to overcome robustness defenses—we instead select the honest outlier that is most surrounded by other honest updates (see Appendix for details).
-With this strategy, the “extreme” honest outliers hide Byzantine updates, which makes the attack harder to detect while remaining effective in heterogeneous settings.
-For instance, based on the extensive results in Appendix B.4, Tables 14, 15, and 16, Mimic often appears as the worst attack for the strategies –, BKT, and NNM.
-All state-of-the-art strategies are evaluated using four robust aggregators: CWMed, CwTM, GM, and MKrum.
-7.5 Metrics, Tables & Boxplots
-We use two metrics: (i) Test Accuracy, averaged over training with a test set that matches the global training label distribution, and (ii) Gradient Dissimilarity, also averaged over training.
-We ran experiments on four datasets, three levels of heterogeneity, four aggregation methods, and five attacks, averaging results over three random seeds for each scenario.
-Full results are in Appendix B.4.
-For clarity, we summarize key findings in three tables.
-Table 2 shows Test Accuracy, averaged over all aggregation methods (CWMed, CwTM, GM, MKrum) and all attacks (ALIE, FOE, LF, Mimic, SF), to reflect overall performance.
-Table 4 presents the same summary for Gradient Dissimilarity.
-Table 3 shows Test Accuracy under the strongest attack (i.e., the one causing the lowest performance), averaged over all aggregators, to highlight worst-case performance.
-In Tables 2 and 3, we report the Accuracy Gain Over Top Non-WoLA Methods—the difference between each WoLA variant and the best-performing non-WoLA method (among –, BKT, FoundFL, and NNM).
-These values highlight performance gains over state-of-the-art strategies.
-We also include boxplots of the gains from Tables 2 and 3 in Figures 4(a) and 4(b), to provide an overall view.
-In the following, when describing a given situation, we refer to the typical performance as the interquartile range of the corresponding boxplot.
-The average gain is also taken from these boxplots.
-7.6 Optimization Algorithm
-WoLA relies on approximating the global class gradients using local datasets (see Subsection 5.3).
-Approximation accuracy improves with increased data per class at each optimization step.
-As a result, the best performance is likely achieved with non-stochastic gradient descent, where honest workers compute gradients over the entire dataset.
-In practice, however, stochastic methods are more common—workers typically compute gradients on random mini-batches.
-To match this setting, we follow [3] and adopt the distributed stochastic heavy ball method (see Algorithm in Appendix B.1).
-This method has demonstrated strong robustness against Byzantine behavior in both heterogeneous [3] and homogeneous settings [13, 25, 16].
-We use momentum $\beta=0.9$ for MNIST and Fashion-MNIST and $0.99$ for CIFAR10 and Purchase100.
-7.7 Models & Hyperparameters
-On MNIST and Fashion-MNIST, we train a lightweight CNN for $T=800$ steps.
-On CIFAR10, we use a deeper CNN; on Purchase100, a fully connected neural network.
-Both models are trained for $T=2000$ steps.
-The model architectures are provided in Appendix B.2.
-Unless otherwise specified, the loss is cross-entropy.
-The batch size is $b=128$, learning rates decay over time, gradient clipping is fixed at $5$, and $\ell_{2}$ regularization is set to $10^{-4}$.
-7.8 Reproducibility
-All experiments are run using the three seeds $\{1,2,3\}$ to ensure reproducibility. The code will be made publicly available for reuse at https://anonymous.4open.science/r/WoLA-E4CB.
-8 Empirical Results
-In this section, we present experiments to answer the following Research Questions (RQ). We use WoLA methods to refer collectively to WoLA, WoLA, and their combinations with NNM.
-  • RQ1
-    Does WoLA guide training toward a better model?
-  • RQ2
-    Are WoLA and WoLA more robust overall than state-of-the-art strategies?
-  • RQ3
-    Are WoLA and WoLA more robust in worst-case scenarios than state-of-the-art strategies?
-  • RQ4
-    Does combining WoLA or WoLA with NNM improve robustness?
-  • RQ5
-    What explains the robustness of WoLA methods?
-  • RQ6
-    How do batch size and momentum affect WoLA performance?
-[…]
-8.1 Targeted Model with WoLA (RQ1)
-As expected from Proposition 3 (Subsection 6.2), WoLA does not target a better model during training.
-Figure 5 shows Test Accuracy over a training without Byzantine workers, using simple mean aggregation.
-The model trained with WoLA performs similarly to the unweighted loss, though slightly worse on Purchase100 under high heterogeneity.
-8.2 Robustness of WoLA and WoLA (RQ2)
-WoLA and WoLA significantly improve robustness.
-They perform slightly below the best state-of-the-art methods when $f=2$, offer marginal gains at $f=4$, and outperform all non-WoLA methods as the number of Byzantine workers increases.
-When $f=2$: WoLA and WoLA show similar performance, slightly below the best non-WoLA strategies.
-Typical negative gains between $-1.1$ and 00 points (Figure 4(a)).
-Notable exceptions exist, such as Purchase100 with high heterogeneity (Table 2, $\alpha=0.3$), where NNM outperforms WoLA and WoLA by more than $6$ points.
-When $f=4$: WoLA and WoLA still perform similarly and yield slightly better results than non-WoLA methods, with typical gains of a few points (Figure 4(a)).
-According to Table 2, gains range from $+5.4$ (CIFAR10, $\alpha=0.3$) to $-2.9$ (Purchase100, $\alpha=0.3$).
-When $f=6$: WoLA slightly surpasses WoLA, and both outperform other methods.
-WoLA shows typical gains between $+3.3$ and $+12.1$ points, while WoLA reaches between $+1.8$ and $+8.1$ (Figure 4(a)).
-Table 2 reports maximum gains of $+21.4$ (WoLA) and $+20.5$ (WoLA) on MNIST with $\alpha=0.3$.
-When $f=8$: WoLA is better than WoLA, and both methods clearly outperform all others (Figure 4(a)).
-WoLA typically achieves gains between $+5.6$ and $+14.7$ points (median around $+10.1$, Figure 4(a)), while WoLA ranges from $-3.7$ to $+7$ (median around $+5.2$, Figure 4(a)).
-The best results appear on MNIST with medium heterogeneity ($\alpha=1$): $+17.3$ for WoLA and $+13.6$ for WoLA (Table 2).
-Importantly, when $f>2$, Table 2 shows that WoLA almost always ranks in the top three regardless of the scenario (exceptions: Purchase100 $\alpha=0.3$, $f=4$ and CIFAR10 $\alpha=0.3$, $f=8$).
-While WoLA does not frequently appear among the top values, it outperforms non-WoLA strategies most of the time on MNIST and Fashion MNIST, often on Purchase100, and occasionally on CIFAR10.
-8.3 Worst-Case Robustness of WoLA and WoLA (RQ3)
-Considering the strongest attack for each approach, WoLA and WoLA consistently enhance robustness, almost always outperforming other state-of-the-art methods, with gains reaching up to $+41.9$ points (Table 3, MNIST, $\alpha=1$, $f=8$, WoLA).
-For $f=2$: WoLA and WoLA perform similarly, slightly surpass non-WoLA methods with typical gains of a few points (Figure 4(b)), and up to $+11.3$ points on CIFAR10, $\alpha=0.3$ (Table 3).
-For $f=4$: WoLA and WoLA continue to perform similarly, outperforming all baselines with median gains around $+6.3$ and $4.6$ points respectively (Figure 4(b)).
-According to Table 3, WoLA’s gains range from $+0.1$ (MNIST, $\alpha=3$) to $+28.8$ (CIFAR10, $\alpha=0.3$), while WoLA ranges from 00 (MNIST, $\alpha=3$) to $+27$ (CIFAR10, $\alpha=0.3$).
-For $f=6$ or $8$: WoLA outperforms WoLA, and both clearly surpass all other methods (Figure 4(b)).
-WoLA achieves median gains around $+13.4$ points, WoLA around $+12.4$ (Figure 4(b)).
-Table 3 shows that WoLA’s gains range from $-0.8$ (MNIST, $\alpha=3$, $f=6$) to $+41.9$ (MNIST, $\alpha=1$, $f=8$); WoLA ranges from $-3.4$ (CIFAR10, $\alpha=0.3$, $f=8$) to $+41.3$ (MNIST, $\alpha=1$, $f=8$).
-8.4 Robustness Gains from Combining WoLA and WoLA with NNM (RQ4)
-Combining WoLA or WoLA with NNM significantly improves robustness, consistently outperforming WoLA, WoLA or NNM alone.
-WoLA+NNM consistently achieves the top or top-three performance in Table 2 as soon as $f>2$, reflecting strong robustness across settings.
-Gains over the best non-WoLA methods can reach around $20$ points (Table 2).
-For example, on MNIST with $\alpha=0.3$ and $f=8$, the gain is $+26.6$; on Fashion MNIST with $\alpha=0.3$ and $f=6$, $+19.8$; on Purchase100 with $\alpha=1$ and $f=8$, $+25.2$; and on CIFAR10 with $\alpha=1$ and $f=6$, $+18.5$.
-The median gain is $+10.4$ when $f=6$ and $+17.8$ when $f=8$ (Figure 4(a)).
-Considering worst attacks, improvements are even larger, with median gains around $+20$ for both $f=6$ and $f=8$ (Figure 4(b)).
-WoLA+NNM also regularly reaches top-three values in Tables 2 and 3, confirming its robustness.
-It performs similarly to WoLA+NNM when $f\leq 4$, and slightly worse when $f>4$, but still outperforms WoLA alone and all non-WoLA strategies (Figure 4(a), 4(b)).
-Importantly, it turns nearly all negative gains of WoLA alone into positive ones (Table 2).
-For instance, on CIFAR10 with $\alpha=3$ and $f=8$, the score improves from $-4.4$ to $+6.7$; with $\alpha=1$ and $f=8$, from $-3.6$ to $+8.3$; and with $\alpha=0.3$ and $f=2$, from $-0.5$ to $+2.2$.
-Overall, median gains over non-WoLA methods are around $+8$ when $f=6$ or $8$ (Figure 4(a)).
-Considering worst attacks, the median gain is around $+13$ for both $f=6$ and $f=8$ (Figure 4(b)).
-8.5 Sources of Robustness in WoLA Methods (RQ5)
-Gradient dissimilarity is a key indicator of system robustness [24, 3].
-As expected from Proposition 2 (Subsection 6.2), WoLA methods significantly reduce gradient dissimilarity, thereby enhancing robustness against Byzantine attacks.
-Table 4 shows that WoLA and its variants consistently achieve much lower gradient dissimilarity across all scenarios.
-Interestingly, combining NNM with WoLA further reduces gradient dissimilarity, which may explain its greater robustness over WoLA alone.
-Figure 3 shows the evolution of Test Accuracy and Gradient Dissimilarity during training on Fashion-MNIST with high heterogeneity ($\alpha=0.3$) and six Byzantine workers ($f=6$) under two attacks (ALIE and LF).
-We also observe that WoLA’s methods maintain lower gradient dissimilarity than other state-of-the-art approaches.
+Figure 3: Test Accuracy (averaged over three seeds) without Byzantine workers ($f=0$), using the mean aggregator on MNIST, Fashion MNIST, Purchase100, and CIFAR10, under low, medium, and high heterogeneity ($\alpha=3,1$, and $0.3$).
+Each column represents a dataset, and each row a heterogeneity level.
+Legend: $\bullet\!$ WoLA $\bullet\!
+-$
+(a) Results under ALIE attack.
+(b) Results under LF attack.
+Figure 4: Test Accuracy and Gradient Dissimilarity (averaged over three seeds) under ALIE and LF attacks on Fashion MNIST with high heterogeneity ($\alpha=0.3$) and $6$ Byzantine workers ($f=6$).
+Each column corresponds to a robust aggregator, and each row to an evaluation metric.
+Legend: $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM $\bullet\!$ WoLA $\bullet\!$ WoLA+NNM $\bullet\!$ NNM $\bullet\!$ FoundFL $\bullet\!
+-$ and BKT
 (a) Fashion MNIST
@@
 Accuracy is reported in %, with a shared color scale for WoLA and NNM accuracy within each row.
-8.6 Impact of Batch Size & Momentum on WoLA (RQ6)
+Targeted Model with WoLA (RQ A)
+As expected from Proposition 3 in Subsection WoLA Properties, WoLA does not target a better model during training.
+Figure 3 shows Test Accuracy over a training without Byzantine workers, using simple mean aggregation.
+The model trained with WoLA performs similarly to the unweighted loss, though slightly worse on Purchase100 under high heterogeneity.
+Sources of Robustness in WoLA Methods (RQ B)
+Gradient dissimilarity is a key indicator of system robustness (Karimireddy, He, and Jaggi 2020; Allouah et al.
+2023).
+As expected from Proposition 2 in Subsection WoLA Properties, WoLA methods significantly reduce gradient dissimilarity, thereby enhancing robustness against Byzantine attacks.
+Table 3 shows that WoLA and its variants consistently achieve much lower gradient dissimilarity across all scenarios.
+Interestingly, combining NNM with WoLA further reduces gradient dissimilarity, which may explain its greater robustness over WoLA alone.
+Figure 4 shows the evolution of Test Accuracy and Gradient Dissimilarity during training on Fashion-MNIST with high heterogeneity ($\alpha=0.3$) and six Byzantine workers ($f=6$) under two attacks (ALIE and LF).
+We also observe that WoLA’s methods maintain lower gradient dissimilarity than other state-of-the-art approaches.
+Impact of Batch Size & Momentum on WoLA (RQ C)
 Figure 5 reports the mean Test Accuracy on Fashion MNIST and CIFAR-10 achieved by WoLA under three heterogeneity levels ($\alpha=3$, $1$, and $0.3$), with $17$ workers including $5$ Byzantine nodes ($f=5$), using the ALIE attack and the CwTM aggregator.
@@
 These conclusions are specific to the given setting and are not a substitute for a full-scale study, but they provide useful insights into parameter sensitivity.
-9 Conclusion and Future Work
-We show that label skew heterogeneity can be partly addressed with WoLA, a weighted loss where weights are set according to each worker’s local label distribution and a shared training objective.
-WoLA is a simple, model-agnostic method that requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine workers, and adds no computational overhead.
-It is easy to implement and can be combined with other robustness strategies, such as pre-aggregators and robust aggregators.
-We provide theoretical insights into WoLA’s behavior, supported by empirical evaluation.
-Extensive experiments demonstrate WoLA’s efficiency and its superiority over state-of-the-art strategies—both on average and in worst-case scenarios (i.e., under the strongest attack for each method)—especially in adversarial settings involving both Byzantine behavior and heterogeneity.
-Even better results are achieved when WoLA is combined with NNM.
-We also show that WoLA significantly reduces gradient dissimilarity among honest workers, which is key to its robustness.
-For future work, WoLA’s robustness can be improved and its scope extended—for instance, by refining the weighting scheme, such as introducing thresholds or adapting it dynamically at the mini-batch scale.
-The impact of training objective attacks can also be explored more thoroughly, along with the design of stronger defenses against them.
-Finally, developing a decentralized version of WoLA also seems to be a relevant direction.
-Appendix A Theoretical Results
-In this section, we define the random counterparts of the local updates and provide all proofs of the theoretical results given in Section 6.2.
-A.1 Gradients Modeling
-In this subsection, we define the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$.
-Let the global dataset be $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ with $(x_{i},y_{i})\in\mathcal{X}\times[C]$, and let $z_{1},z_{2},\ldots,z_{N}$ be integers in $\mathcal{H}$ such that $z_{k}=i$ if $(x_{k},y_{k})\in\mathcal{D}_{i}$.
-Since $N_{i}=\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)$, the following holds for $\nabla\mathcal{L}_{i}$:
-$\nabla\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\ell\left(y,\Phi\left(x\right)\right)$
-$=$ $\frac{N}{N_{i}}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\ell\left(y_{k},\Phi\left(x_{k}\right)\right)}{N}$
-$=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}$
-Let $L_{i}$ be the random counterpart of $\nabla\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $L_{i}$ is given by:
-$L_{i}=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (5)
-$\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
-We now turn to $\nabla\mathcal{WL}_{i}$:
-$\nabla\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
-$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\frac{q^{y_{k}}}{p^{y_{k}}_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
-$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\sum_{c=1}^{C}\mathbbm{1}(y_{k}=c)\frac{q^{c}}{N_{i}^{c}/N_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
-$=$ $\sum_{c=1}^{C}q^{c}\frac{N}{N_{i}^{c}}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}}{N}$
-Since $N^{c}_{i}=\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)$, we have:
-$\nabla\mathcal{WL}_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}\Bigg{)}$
-Let $W_{i}$ be the random counterpart of $\nabla\mathcal{WL}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $W_{i}$ is given by:
-$W_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (6)
-$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
-A.2 Lemmas
-In this subsection, we give two lemmas that will be used in the subsequent theoretical proofs.
-Lemma 2
-Let $X$, $Y$, and $Z$ be random variables such that $X\mid Y$ is independent of $Z\mid Y$.
-Let $f$ be a function defined on the space of realizations of $X$.
-Then, for all $y$ and $z$ such that $\mathbb{P}(Y=y)>0$ and $\mathbb{P}(Z=z)>0$, we have:
-$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
-Proof.
-$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=y)\mathbbm{1}(Z=z)f(X)\Big{]}}{\mathbb{P}(Y=y,Z=z)}$
-$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
-$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\Big{|}Y=y\Big{]}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
-$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\mathbbm{1}(Y=y)\Big{]}}{\mathbb{P}(Y=y)}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
-$=$ $\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
-Lemma 3
-The following expression converges almost surely (a.s.) as $N\to\infty$:
-$\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
-[…]
-Proof.
-By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
-$\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
-$=\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
-By the Law of Large Numbers, as $N\to\infty$, the expressions converge almost surely to:
-$\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\Big{]}^{-1}$
-$\quad\quad\quad\quad\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}$
-$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}}{\mathbb{P}(Y=c,Z=i)}$
-$=$ $\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}$
-According to Lemma 2:
-$\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
-A.3 Proof of Proposition 1
-We begin by calculating the limit of $L_{i}$. By definition, we have:
-$L_{i}=\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
-By the Law of Large Numbers, it holds:
-[…]
-$\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
-[…]
-By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
-$\lim_{N\to\infty}L_{i}=\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\quad\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
-$=$ $\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
-[…]
-Similarly, for $L$:
-$\lim_{N\to\infty}L=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}L_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
-$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}L_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
-[…]
-By the Law of total expectation, it holds:
-[…]
-A.4 Proof of Proposition 2
-We begin by calculating the limit of $W_{i}$. By the continuity of the involved functions, it holds:
-$\lim_{N\to\infty}W_{i}$
-$=$ $\lim_{N\to\infty}\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
-$=$ $\sum_{c=1}^{C}q^{c}\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
-$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
-Using Lemma 3, as $N\to\infty$, the expression converge almost surely to:
-[…]
-Similarly, for $W$:
-$\lim_{N\to\infty}W=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}W_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
-$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}W_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
-By the Law of Large Numbers, it holds:
-[…]
-Using above results, it concludes the lemma:
-[…]
-$=$ $\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
-A.5 Proof of Gradient Dissimilarity Vanishing
-According to Proposition 2, by the continuity of the gradient dissimilarity formula, we obtain:
-$\lim_{N\to\infty}\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\left\|W_{i}-W\right\|^{2}$
-[…]
-A.6 Proof of Proposition 3
-According to Proposition 2, it holds:
-[…]
-By assumption, it holds $\tilde{p}_{c}=\mathbb{P}(Y=c)$ for all $c\in[C]$. Then, using the Law of total expectation, it follows:
-[…]
-$=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
-A.7 Worst Training Objective Attack
+Appendix E Training Objective Attack
+Our experiments include scenarios where workers share local label distributions with the server, while Byzantine machines manipulate the training objective by sending malicious distributions.
+We describe this attack setting below.
+WoLA Scenario
+Similar to how Byzantine workers can send arbitrary gradient updates with full knowledge of honest gradients, we assume they can also send arbitrary label distributions with full knowledge of the honest workers’ local label distributions.
+Rather than enumerating all possible attacks, we focus on a worst-case attack: Byzantine machines submit a zero vector except for a single entry set to $1$, corresponding to the minority class in the global dataset (proofs are provided in the next subsection).
+To defend against such attacks, we adopt the geometric median as the aggregation method.
+This choice offers two main advantages: (i) Unlike methods such as Trimmed Mean or Krum, which exclude outliers, the geometric median treats all updates equally.
+This is crucial since outlier distributions are not necessarily malicious; excluding them risks discarding honest data and harming performance.
+(ii) When aggregating probability distributions, the geometric median produces a valid probability distribution—unlike the coordinate-wise median—thus preserving the training objective’s expected form.
+Our aim is to show that even a worst-case attack has limited impact on model performance when countered by a simple defense. We denote our method under this attack scenario as WoLA.
+Proof that WoLA is a Worst-Case Attack
 We show that the training objective $q$ is maximally deviated by a Byzantine attack that concentrates all its mass on the class with minimal weight among honest workers, under the $L_{1}$ norm denoted by $\|\cdot\|_{1}$.
@@
 This bound is tight: it is achieved by setting each $q_{i}$ for $i\notin\mathcal{H}$ to put all mass on the class where $u$ is minimal, completing the proof.
-Appendix B Experiment
-B.1 Robust Distributed Stochastic Heavy Ball Algorithm
+Appendix F Theoretical Proofs
+In this section, we define the random counterparts of the local updates and provide all proofs of the theoretical results given in Section WoLA Properties.
+Gradients Modeling
+In this subsection, we define the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{W}\!\mathcal{L}_{i}$.
+Let the global dataset be $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ with $(x_{i},y_{i})\in\mathcal{X}\times[C]$, and let $z_{1},z_{2},\ldots,z_{N}$ be integers in $\mathcal{H}$ such that $z_{k}=i$ if $(x_{k},y_{k})\in\mathcal{D}_{i}$.
+Since $N_{i}=\sum_{l=1}^{N}\mathbf{1}(z_{l}=i)$, the following holds for $\nabla\mathcal{L}_{i}$:
+$\nabla\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\ell\left(y,\Phi\left(x\right)\right)$
+$=$ $\frac{N}{N_{i}}\frac{\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\ell\left(y_{k},\Phi\left(x_{k}\right)\right)}{N}$
+$=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}$
+Let $L_{i}$ be the random counterpart of $\nabla\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection Theoretical Framework), $L_{i}$ is given by:
+$L_{i}=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (2)
+$\quad\quad\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
+We now turn to $\nabla\mathcal{W}\!\mathcal{L}_{i}$:
+$\nabla\mathcal{W}\!\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
+$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\frac{q^{y_{k}}}{p^{y_{k}}_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
+$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\sum_{c=1}^{C}\mathbf{1}(y_{k}=c)\frac{q^{c}}{N_{i}^{c}/N_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
+$=$ $\sum_{c=1}^{C}q^{c}\frac{N}{N_{i}^{c}}\frac{\sum_{k=1}^{N}\mathbf{1}(y_{k}=c)\mathbf{1}(z_{k}=i)\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}}{N}$
+Since $N^{c}_{i}=\sum_{l=1}^{N}\mathbf{1}(y_{l}=c)\mathbf{1}(z_{l}=i)$, we have:
+$\nabla\mathcal{W}\!\mathcal{L}_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(y_{l}=c)\mathbf{1}(z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(y_{k}=c)\mathbf{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}\Bigg{)}$
+Let $W_{i}$ be the random counterpart of $\nabla\mathcal{W}\!\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection Theoretical Framework), $W_{i}$ is given by:
+$W_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (3)
+$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
+Lemmas
+In this subsection, we give two lemmas that will be used in the subsequent theoretical proofs.
+Lemma 1
+Let $X$, $Y$, and $Z$ be random variables such that $X\mid Y$ is independent of $Z\mid Y$.
+Let $f$ be a function defined on the space of realizations of $X$.
+Then, for all $y$ and $z$ such that $\mathbb{P}(Y=y)>0$ and $\mathbb{P}(Z=z)>0$, we have:
+$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
+Proof.
+$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\frac{\mathbb{E}\Big{[}\mathbf{1}(Y=y)\mathbf{1}(Z=z)f(X)\Big{]}}{\mathbb{P}(Y=y,Z=z)}$
+$=$ $\mathbb{E}\Big{[}\mathbf{1}(Z=z)f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
+$=$ $\mathbb{E}\Big{[}\mathbf{1}(Z=z)\Big{|}Y=y\Big{]}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
+$=$ $\frac{\mathbb{E}\Big{[}\mathbf{1}(Z=z)\mathbf{1}(Y=y)\Big{]}}{\mathbb{P}(Y=y)}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
+$=$ $\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
+Lemma 2
+The following expression converges almost surely (a.s.) as $N\to\infty$:
+$\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
+$\quad\quad\quad\quad\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
+Proof.
+By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
+$\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
+$=\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
+By the Law of Large Numbers, as $N\to\infty$, the expressions converge almost surely to:
+$\mathbb{E}\Big{[}\mathbf{1}(Y=c)\mathbf{1}(Z=i)\Big{]}^{-1}$
+$\quad\quad\quad\quad\mathbb{E}\Big{[}\mathbf{1}(Y=c)\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}$
+$=$ $\frac{\mathbb{E}\Big{[}\mathbf{1}(Y=c)\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}}{\mathbb{P}(Y=c,Z=i)}$
+$=$ $\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}$
+According to Lemma 1:
+$\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
+Proof of Proposition 1
+We begin by calculating the limit of $L_{i}$. By definition, we have:
+$L_{i}=\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
+By the Law of Large Numbers, it holds:
+$\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\mathbf{1}(Z=i)\Big{]}=\mathbb{P}(Z=i),$
+$\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
+$\quad\quad\quad\quad\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}[\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))].$
+By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
+$\lim_{N\to\infty}L_{i}=\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\quad\quad\quad\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
+$=$ $\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
+$\overset{\text{a.s.}}{=}$ $\frac{\mathbb{E}[\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))]}{\mathbb{P}(Z=i)}=\mathbb{E}[\nabla\ell(Y,\Phi(X))|Z=i]$
+Similarly, for $L$:
+$\lim_{N\to\infty}L=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}L_{i}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
+$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}L_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
+$\overset{\text{a.s.}}{=}$ $\sum_{i\in\mathcal{H}}\mathbb{E}[\nabla\ell(Y,\Phi(X))|Z=i]\mathbb{P}(Z=i)$
+By the Law of total expectation, it holds:
+$\lim_{N\to\infty}L\overset{\text{a.s.}}{=}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
+Proof of Proposition 2
+We begin by calculating the limit of $W_{i}$. By the continuity of the involved functions, it holds:
+$\lim_{N\to\infty}W_{i}$
+$=$ $\lim_{N\to\infty}\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
+$=$ $\sum_{c=1}^{C}q^{c}\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
+$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
+Using Lemma 2, as $N\to\infty$, the expression converge almost surely to:
+$\lim_{N\to\infty}W_{i}\overset{\text{a.s.}}{=}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
+Similarly, for $W$:
+$\lim_{N\to\infty}W=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}W_{i}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
+$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}W_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
+By the Law of Large Numbers, it holds:
+$\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\overset{\text{a.s.}}{=}\mathbb{E}\Big{[}\mathbf{1}(Z=i)\Big{]}=\mathbb{P}(Z=i),$
+Using above results, it concludes the lemma:
+$\lim_{N\to\infty}W\overset{\text{a.s.}}{=}\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}\mathbb{P}(Z=i)$
+$=$ $\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
+Proof of Gradient Dissimilarity Vanishing
+According to Proposition 2, by the continuity of the gradient dissimilarity formula, we obtain:
+$\lim_{N\to\infty}\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\left\|W_{i}-W\right\|^{2}$
+$=$ $\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\left\|\lim_{N\to\infty}W_{i}-\lim_{N\to\infty}W\right\|^{2}\overset{\text{a.s.}}{=}0$
+Proof of Proposition 3
+According to Proposition 2, it holds:
+$W\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
+By assumption, it holds $\tilde{p}_{c}=\mathbb{P}(Y=c)$ for all $c\in[C]$. Then, using the Law of total expectation, it follows:
+$W\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}$ $\sum_{c=1}^{C}\mathbb{P}(Y=c)\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
+$=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
+Appendix G Experimental Details
+Robust Distributed Stochastic Heavy Ball Algorithm
 We present here the robust distributed stochastic Heavy Ball algorithm.
-More details can be found in [3].
+More details can be found in (Allouah et al.
+2023).
 We use superscript in brackets to denote the iteration index, i.e., (t), and subscripts to indicate the worker index.
-Input: Initial model $\theta^{(0)}$, initial momentum $m_{i}^{(0)}=0$ for honest workers, robust aggregation $F$, learning rate $\gamma$, momentum coefficient $\beta$, and number of steps $T$.
-for $t=1,\ldots,T$ do
-       Server broadcasts $\theta^{(t-1)}$ to all workers;
-       for every honest worker $i$ do
-             Compute a stochastic gradient $g^{(t)}_{i}$;
-             Update local momentum $m_{i}^{(t)}\leftarrow\beta m_{i}^{(t-1)}+(1-\beta)g^{(t)}_{i}$;
-             Send $m_{i}^{(t)}$ to the server;
-       end for
-      Server aggregates the momentums:
-           $R^{(t)}=F(m_{1}^{(t)},\ldots,m_{n}^{(t)})$;
-       Server updates the model: $\theta^{(t)}=\theta^{(t-1)}-\gamma R^{(t)}$;
-end for
 Algorithm 1 Robust Distributed Stochastic Heavy Ball
-B.2 Model Architectures
-We adopt the following abreviations : L(#outputs) denotes a fully connected linear layer; R is a ReLU activation; T is a Tanh activation; C(#channels) represents a 2D convolutional layer with kernel size $5$, padding 00, and stride $1$; M denotes 2D max-pooling with kernel size $2$; B stands for batch normalization; and D represents dropout with a fixed probability of $0.25$.
+0: Initial model $\theta^{(0)}$, initial momentum $m_{i}^{(0)}=0$ for honest workers, robust aggregation $F$, learning rate $\gamma$, momentum coefficient $\beta$, and number of steps $T$
+1:for $t=1$ to $T$ do
+2:  Server broadcasts $\theta^{(t-1)}$ to all workers
+3:  for every honest worker $i$ do
+4:   Compute a stochastic gradient $g^{(t)}_{i}$
+5:   Update local momentum: $m_{i}^{(t)}\leftarrow\beta m_{i}^{(t-1)}+(1-\beta)g_{i}^{(t)}$
+6:   Send $m_{i}^{(t)}$ to the server
+7:  end for
+8:  Server aggregates the momenta:
+9:    $R^{(t)}=F(m_{1}^{(t)},\ldots,m_{n}^{(t)})$
+10:  Server updates the model:
+11:    $\theta^{(t)}=\theta^{(t-1)}-\gamma R^{(t)}$
+12:end for
+Model Architectures
+We adopt the following abreviations : L(#outputs) denotes a fully connected linear layer; R is a ReLU activation; T is a Tanh activation; C(#channels) represents a 2D convolutional layer with kernel size $5$, padding 0, and stride $1$; M denotes 2D max-pooling with kernel size $2$; B stands for batch normalization; and D represents dropout with a fixed probability of $0.25$.
 The model architectures are as follows:
-  • MNIST and Fashion MNIST (CNN from [3]): C(20)-R-M–C(20)-R-M–L(500)-R–L(10)
-  • CIFAR10 (CNN from [3]):(3,32×32)-C(64)-R-B-C(64)-R-B-M-D-C(128)-R-B-C(128)-R-B-M-D-L(128)-R-D-L(10)
-  • Purchase100 (fully connected neural networks from [31]): L(1024)-L(512)-T-L(256)-T-L(100)
+  • MNIST and Fashion MNIST (CNN from (Allouah et al. 2023)): C(20)-R-M–C(20)-R-M–L(500)-R–L(10)
+  • CIFAR10 (CNN from (Allouah et al. 2023)):(3,32×32)-C(64)-R-B-C(64)-R-B-M-D-C(128)-R-B-C(128)-R-B-M-D-L(128)-R-D-L(10)
+  • Purchase100 (fully connected neural networks from (Nasr, Shokri, and Houmansadr 2018)): L(1024)-L(512)-T-L(256)-T-L(100)
 The learning rates are defined as:
-  • MNIST and Fashion MNIST (from [3]): $\gamma_{t}=0.75/(1+\left\lfloor\frac{t}{50}\right\rfloor)$
-  • CIFAR10 (from [3]) and Purchase100: $\gamma_{t}=0.25$ if $t\leq 1500$, $0.025$ otherwise.
-B.3 Mimic with Most Surrounded Outlier Selection
+  • MNIST and Fashion MNIST (from (Allouah et al. 2023)): $\gamma_{t}=0.75/(1+\left\lfloor\frac{t}{50}\right\rfloor)$
+  • CIFAR10 (from (Allouah et al. 2023)) and Purchase100: $\gamma_{t}=0.25$ if $t\leq 1500$, $0.025$ otherwise.
+Mimic with Most Surrounded Outlier Selection
 In the Mimic attack, we choose the honest outlier that is most surrounded by other honest updates.
@@
 where $\operatorname{dist}$ denotes the Euclidean distance.
-For any vector $\mathbf{v}_{j}$, there exists an index $r$ such that $\sigma(i,r)=j$, meaning that $\mathbf{v}_{j}$ is the $r$‑th farthest vector from $\mathbf{v}_{i}$.
+For any vector $\mathbf{v}_{j}$, there exists an index $r$ such that $\sigma(i,r)=j$, meaning that $\mathbf{v}_{j}$ is the rr‑th farthest vector from $\mathbf{v}_{i}$.
 Note that $\sigma(i,h)=i$.
@@
 The update to mimic is the one with the highest score, see Algorithm 2.
-Input: Honest updates $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}\in\mathbb{R}^{d}$, number of Byzantine workers $f$
-Output: Update $\mathbf{v}_{\text{argmax}}$ with highest score
-for $i=1,\ldots,h$ do
-       $s_{i}\leftarrow 0$ ;
-end for
-$\textit{center}\leftarrow\frac{1}{h}\sum_{i=1}^{h}\mathbf{v}_{i}$ ;
-for $i=1,\ldots,h$ do
-       for $j=1,\ldots,h$ do
-             $r\leftarrow\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})$ ;
-             $s_{j}\leftarrow s_{j}+\min(f,r)\cdot\operatorname{dist}(\mathbf{v}_{j},\textit{center})$ ;
-       end for
-end for
-$k\leftarrow\arg\max_{j}s_{j}$ return $\mathbf{v}_{k}$
 Algorithm 2 Mimic with Most Surrounded Outlier Selection
-B.4 Comprehensive Results
-We report the full experimental results.
-For MNIST, see Tables 5, 6, and 7; for Fashion MNIST, see Tables 8, 9, and 10; for Purchase100, see Tables 11, 12, and 13; and for CIFAR10, see Tables 14, 15, and 16.
+0: Honest updates $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}\in\mathbb{R}^{d}$, number of Byzantine workers $f$
+0: Update $\mathbf{v}_{\text{argmax}}$ with highest score
+1:for $i=1$ to $h$ do
+2:  $s_{i}\leftarrow 0$
+3:end for
+[…]
+5:for $i=1$ to $h$ do
+6:  for $j=1$ to $h$ do
+7:   $r\leftarrow\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})$
+8:   $s_{j}\leftarrow s_{j}+\min(f,r)\cdot\operatorname{dist}(\mathbf{v}_{j},\textit{center})$
+9:  end for
+10:end for
+[…]
+12:return  $\mathbf{v}_{k}$
+Appendix H Full and Additional Results
+We report summarized performance in Table 4 for MNIST and Purchase100, which show similar method behaviors as Table 1.
+Additionally, full experimental results are provided.
+For MNIST, see Tables 56, and 7; for Fashion MNIST, see Tables 89, and 10; for Purchase100, see Tables 1112, and 13; and for CIFAR10, see Tables 1415, and 16.
 Each set of three tables corresponds to low, medium, and high heterogeneity levels ($\alpha=3$, $1$, and $0.3$).
+Table 4: Test accuracy (± standard deviation) in %, averaged over training, on MNIST and Purchase100.
+Global results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF).
+Worst-case results correspond to the worst attack, averaged across aggregators.
+Colors rank values within columns: yellow (1st), blue (2nd), red (3rd).
+See full table in Appendix.
+(a) MNIST (Global)
+Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
+Accuracy
+- 93.6 (0.3) 92.9 (0.4) 86.1 (1.6) 66.3 (4.5) 93.6 (0.3) 92.2 (0.6) 82.3 (3.3) 43.5 (11.7) 91.9 (1.8) 84.6 (4.0) 58.8 (12.1) 17.3 (7.3)
+BKT 93.5 (0.3) 92.8 (0.5) 86.1 (1.6) 66.3 (4.5) 93.6 (0.4) 92.5 (0.7) 82.3 (3.3) 43.5 (11.7) 93.3 (0.7) 89.6 (1.9) 58.8 (12.1) 17.3 (7.3)
+FoundFL 89.6 (1.8) 74.4 (4.2) 56.8 (4.6) 47.5 (5.0) 85.4 (5.2) 69.6 (7.3) 57.7 (6.7) 50.4 (6.8) 61.6 (15.9) 50.4 (12.8) 43.0 (12.1) 38.0 (11.8)
+NNM 93.7 (0.2) 93.3 (0.3) 92.7 (0.4) 75.5 (6.9) 93.9 (0.3) 93.0 (0.6) 90.1 (2.5) 63.2 (9.0) 93.9 (0.6) 91.5 (1.4) 67.5 (14.5) 35.2 (8.4)
+WoLA 93.7 (0.3) 93.4 (0.3) 92.7 (0.3) 81.9 (0.6) 93.9 (0.3) 93.5 (0.4) 92.4 (0.4) 80.5 (2.2) 93.6 (0.5) 93.0 (0.7) 88.9 (2.6) 54.5 (8.0)
+WoLA+NNM 93.7 (0.3) 93.5 (0.3) 93.2 (0.3) 84.3 (0.6) 94.0 (0.3) 93.6 (0.4) 93.2 (0.3) 83.1 (2.8) 93.7 (0.6) 93.4 (0.7) 90.0 (3.0) 64.6 (8.7)
+WoLA 93.7 (0.3) 93.3 (0.3) 92.5 (0.3) 81.3 (0.8) 93.8 (0.3) 93.3 (0.3) 92.1 (0.6) 76.8 (5.4) 93.6 (0.6) 92.9 (0.7) 88.1 (2.8) 47.3 (17.0)
+WoLA+NNM 93.7 (0.3) 93.4 (0.3) 93.0 (0.3) 83.8 (0.7) 93.9 (0.4) 93.5 (0.3) 92.8 (0.4) 73.5 (13.0) 93.7 (0.6) 93.3 (0.7) 89.8 (2.8) 48.7 (15.8)
+Accuracy Gain Over Top Non-WoLA Methods
+WoLA 0.0 0.0 0.0 +6.4 0.0 +0.4 +2.4 +17.3 -0.4 +1.5 +21.4 +16.5
+WoLA+NNM 0.0 +0.2 +0.6 +8.8 +0.1 +0.6 +3.1 +19.9 -0.2 +1.8 +22.5 +26.6
+WoLA 0.0 0.0 -0.2 +5.9 -0.1 +0.3 +2.0 +13.6 -0.4 +1.4 +20.5 +9.3
+WoLA+NNM 0.0 +0.1 +0.3 +8.3 -0.1 +0.5 +2.8 +10.3 -0.2 +1.8 +22.3 +10.7
+(b) Purchase100 (Global)
+Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
+Accuracy
+- 70.4 (0.7) 65.3 (1.1) 50.5 (1.9) 24.8 (2.8) 67.9 (1.0) 58.0 (2.0) 37.1 (3.1) 8.0 (2.1) 62.2 (1.8) 44.4 (3.3) 15.9 (2.3) 2.6 (1.1)
+BKT 70.1 (0.9) 64.6 (1.2) 50.5 (1.9) 24.8 (2.8) 68.2 (1.1) 59.3 (1.9) 37.1 (3.1) 8.0 (2.1) 64.5 (1.9) 49.3 (3.5) 15.9 (2.3) 2.6 (1.1)
+FoundFL 52.2 (3.6) 36.5 (3.1) 28.4 (2.6) 22.4 (2.4) 38.8 (5.9) 28.9 (4.1) 20.7 (4.5) 13.3 (4.9) 24.6 (8.2) 15.6 (6.6) 10.3 (4.3) 6.4 (2.8)
+NNM 71.6 (0.7) 67.9 (0.9) 61.3 (1.8) 38.1 (2.3) 70.6 (0.9) 63.9 (1.6) 52.0 (2.8) 28.3 (4.0) 69.6 (1.2) 58.8 (3.2) 28.9 (5.6) 15.1 (3.2)
+WoLA 70.6 (0.6) 69.2 (0.6) 66.7 (0.7) 52.7 (1.0) 69.2 (0.8) 67.0 (0.8) 61.3 (1.2) 43.1 (1.7) 63.2 (1.1) 57.1 (2.1) 43.6 (2.6) 21.8 (2.5)
+WoLA+NNM 70.9 (0.6) 69.8 (0.6) 68.3 (0.7) 57.6 (0.9) 69.3 (0.8) 67.8 (0.7) 65.3 (1.1) 53.5 (1.7) 64.0 (1.1) 60.6 (1.7) 52.6 (2.5) 33.0 (3.7)
+WoLA 70.5 (0.6) 68.8 (0.5) 65.2 (0.6) 48.2 (1.7) 69.0 (0.8) 66.3 (0.9) 59.3 (1.6) 33.5 (6.1) 62.9 (1.0) 55.9 (1.9) 40.9 (3.4) 11.2 (5.5)
+WoLA+NNM 70.7 (0.6) 69.3 (0.5) 67.0 (0.6) 53.3 (2.6) 69.2 (0.7) 67.2 (0.8) 63.1 (1.5) 39.9 (9.9) 63.7 (1.0) 59.3 (1.7) 45.7 (5.2) 19.5 (7.6)
+Accuracy Gain Over Top Non-WoLA Methods
+WoLA -1.0 +1.3 +5.4 +14.6 -1.4 +3.1 +9.3 +14.8 -6.4 -1.7 +14.6 +6.8
+WoLA+NNM -0.8 +1.8 +7.0 +19.5 -1.3 +3.9 +13.3 +25.2 -5.6 +1.8 +23.7 +17.9
+WoLA -1.2 +0.8 +4.0 +10.2 -1.6 +2.4 +7.2 +5.2 -6.7 -2.9 +11.9 -3.9
+WoLA+NNM -0.9 +1.4 +5.7 +15.3 -1.5 +3.2 +11.1 +11.6 -5.9 +0.5 +16.7 +4.5
+(c) MNIST (Worst-case)
+Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
+Accuracy
+- 93.4 (0.3) 91.9 (0.5) 73.6 (5.1) 27.5 (5.8) 93.3 (0.4) 90.1 (1.3) 62.8 (8.1) 12.6 (5.7) 89.3 (3.2) 70.7 (8.6) 29.5 (11.5) 7.1 (3.4)
+BKT 93.2 (0.5) 91.5 (0.9) 73.6 (5.1) 27.5 (5.8) 93.2 (0.5) 90.8 (1.5) 62.8 (8.1) 12.6 (5.7) 92.6 (1.1) 84.6 (4.0) 29.5 (11.5) 7.1 (3.4)
+FoundFL 80.4 (4.4) 36.5 (8.2) 10.5 (1.2) 10.9 (1.7) 76.4 (15.1) 27.1 (8.7) 10.9 (1.2) 13.5 (3.3) 48.5 (22.4) 20.5 (6.3) 14.7 (4.0) 16.9 (7.6)
+NNM 93.6 (0.2) 92.8 (0.3) 91.3 (0.6) 47.0 (27.5) 93.8 (0.3) 91.6 (1.7) 83.9 (8.9) 22.5 (11.7) 93.7 (0.6) 85.6 (3.7) 49.6 (24.0) 18.1 (5.9)
+WoLA 93.5 (0.3) 92.9 (0.3) 90.6 (0.4) 66.7 (0.7) 93.8 (0.3) 92.9 (0.3) 89.9 (0.7) 64.3 (3.0) 93.3 (0.5) 92.1 (0.8) 77.6 (8.1) 28.8 (6.2)
+WoLA+NNM 93.6 (0.3) 93.2 (0.3) 92.6 (0.2) 70.5 (1.6) 93.9 (0.3) 93.2 (0.3) 92.3 (0.4) 66.7 (4.7) 93.6 (0.5) 92.6 (0.7) 78.9 (11.2) 24.1 (11.4)
+WoLA 93.5 (0.3) 92.8 (0.3) 90.6 (0.4) 67.3 (1.0) 93.7 (0.3) 92.8 (0.3) 89.6 (0.8) 63.7 (8.9) 93.3 (0.5) 92.0 (0.8) 75.9 (8.1) 24.5 (7.3)
+WoLA+NNM 93.6 (0.3) 93.1 (0.3) 92.3 (0.4) 70.3 (0.7) 93.7 (0.3) 93.1 (0.3) 91.9 (0.5) 51.4 (27.9) 93.6 (0.6) 92.6 (0.8) 80.2 (9.5) 14.9 (6.6)
+Accuracy Gain Over Top Non-WoLA Methods
+WoLA -0.1 +0.1 -0.8 +19.7 -0.1 +1.3 +6.0 +41.9 -0.4 +6.5 +28.1 +10.7
+WoLA+NNM 0.0 +0.3 +1.3 +23.5 0.0 +1.6 +8.3 +44.3 -0.1 +7.0 +29.3 +6.0
+WoLA -0.1 0.0 -0.8 +20.3 -0.1 +1.3 +5.7 +41.3 -0.4 +6.4 +26.3 +6.5
+WoLA+NNM +0.1 +0.2 +0.9 +23.2 -0.1 +1.5 +7.9 +28.9 -0.1 +7.0 +30.6 -3.2
+(d) Purchase100 (Worst-case)
+Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
+Accuracy
+- 68.2 (0.6) 59.2 (0.8) 29.0 (1.2) 8.7 (2.3) 64.0 (1.5) 48.5 (2.9) 18.5 (3.1) 0.9 (0.3) 54.9 (2.5) 28.5 (3.8) 2.1 (0.8) 0.9 (0.3)
+BKT 68.6 (0.8) 58.1 (0.9) 29.0 (1.2) 8.7 (2.3) 65.8 (1.6) 50.3 (3.1) 18.5 (3.1) 0.9 (0.3) 58.2 (3.4) 30.0 (7.1) 2.1 (0.8) 0.9 (0.3)
+FoundFL 33.6 (1.9) 4.2 (1.7) 0.6 (0.3) 0.4 (0.1) 21.5 (6.8) 2.6 (1.3) 1.6 (0.9) 1.6 (0.9) 13.2 (4.7) 3.2 (1.1) 2.9 (0.3) 2.5 (1.3)
+NNM 70.4 (0.5) 63.3 (1.5) 52.2 (1.4) 6.4 (1.0) 67.1 (1.7) 53.8 (4.3) 33.6 (4.4) 3.2 (1.8) 61.8 (3.1) 36.9 (8.1) 3.2 (2.3) 1.0 (0.4)
+WoLA 69.7 (0.6) 66.7 (0.6) 61.4 (0.7) 37.7 (1.0) 68.0 (0.8) 63.8 (0.8) 49.0 (1.4) 19.3 (1.2) 60.3 (1.0) 47.3 (2.3) 18.2 (2.6) 6.1 (2.8)
+WoLA+NNM 70.3 (0.6) 68.0 (0.6) 66.3 (0.6) 44.2 (0.8) 68.7 (0.8) 65.8 (0.6) 62.0 (1.5) 37.6 (1.9) 61.7 (1.0) 53.5 (1.5) 39.7 (3.3) 3.0 (1.8)
+WoLA 69.5 (0.6) 66.2 (0.5) 59.8 (0.6) 31.9 (1.9) 67.7 (0.8) 62.9 (0.8) 45.7 (2.5) 21.6 (4.6) 59.9 (0.9) 45.8 (2.0) 23.4 (4.5) 4.2 (1.0)
+WoLA+NNM 70.2 (0.6) 67.5 (0.5) 64.8 (0.4) 36.3 (6.4) 68.5 (0.7) 65.0 (0.7) 58.4 (1.6) 20.0 (13.5) 61.4 (0.9) 51.6 (1.6) 17.4 (15.2) 1.9 (0.7)
+Accuracy Gain Over Top Non-WoLA Methods
+WoLA -0.7 +3.4 +9.2 +29.0 +0.8 +10.0 +15.4 +16.1 -1.5 +10.4 +15.1 +3.6
+WoLA+NNM -0.1 +4.7 +14.1 +35.5 +1.5 +12.0 +28.4 +34.4 -0.1 +16.6 +36.5 +0.5
+WoLA -0.9 +2.8 +7.6 +23.2 +0.5 +9.1 +12.1 +18.4 -1.8 +8.9 +20.3 +1.7
+WoLA+NNM -0.2 +4.2 +12.6 +27.7 +1.3 +11.2 +24.8 +16.8 -0.4 +14.8 +14.2 -0.6
 Table 5: Accuracy (s.d.) % on MNIST with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).